### v2

# A simplified proof of the CSP Dichotomy Conjecture and XY-symmetric operations

Abstract
We develop a new theory of strong subalgebras and linear congruences that are defined globally. Using this theory we provide a new proof of the correctness of Zhuk’s algorithm for all tractable CSPs on a finite domain, and therefore a new simplified proof of the CSP Dichotomy Conjecture. Additionally, using the new theory we prove that composing a weak near-unanimity operation of an odd arity $n$ we can derive an $n$-ary operation that is symmetric on all two-element sets. Thus, CSP over a constraint language $\Gamma$ on a finite domain is tractable if and only if there exist infinitely many polymorphisms of $\Gamma$ that are symmetric on all two-element sets.
1 Introduction
The Constraint Satisfaction Problem (CSP) is the problem of deciding whether a set of constraints has a satisfying assignment. In general, the problem is NP-hard (or even undecidable for infinite domains) and to obtain tractable cases we restrict the set of admissible constraints. Let $A$ be a finite set and $\Gamma$ be a set of relations on $A$, called the constraint language. Then $\operatorname{CSP}(\Gamma)$ is the problem of deciding whether a conjunctive formula
$R_{1}(\dots)\wedge R_{2}(\dots)\wedge\dots\wedge R_{s}(\dots),$ (*)
where $R_{1},\dots,R_{s}\in\Gamma$, is satisfiable. It was conjectured that $\operatorname{CSP}(\Gamma)$ is either in P, or NP-complete [18]. In 2017, two independent proofs of this conjecture appeared [33, 32, 15, 16], and the conjecture became a theorem. To formulate it properly, we need two definitions.
An operation $f$ on a set $A$ is called a weak near-unanimity (WNU) operation if it satisfies $f(y,x,\ldots,x)=f(x,y,x,\ldots,x)=\dots=f(x,x,\ldots,x,y)$ for all $x,y\in A$. We say that an operation $f\colon A^{n}\to A$ preserves a relation $R\subseteq A^{m}$ if
$(a_{1,1},\ldots,a_{1,m}),\dots,(a_{n,1},\ldots,a_{n,m})\in R\Rightarrow(f(a_{1,1},\ldots,a_{n,1}),\ldots,f(a_{1,m},\ldots,a_{n,m}))\in R.$
We say that an operation preserves a set of relations $\Gamma$ if it preserves every relation in $\Gamma$. If $f$ preserves $R$ or $\Gamma$, we also say that $f$ is a polymorphism of $R$ or $f$ is a polymorphism of $\Gamma$, and write $f\in\operatorname{Pol}(R)$ or $f\in\operatorname{Pol}(\Gamma)$, respectively.
Theorem 1 ([33, 32, 15, 16]).
Suppose $\Gamma$ is a finite set of relations on a finite set $A$. Then $\operatorname{CSP}(\Gamma)$ can be solved in polynomial time if there exists a WNU preserving $\Gamma$; $\operatorname{CSP}(\Gamma)$ is NP-complete otherwise.
The NP-hardness for constraint languages without a WNU follows from [13, 14] and [28]. The essential part of each proof of the CSP Dichotomy Conjecture is a polynomial algorithm that works for all tractable cases, and the tricky and cumbersome part is to show that the algorithm works correctly.
One of the two main ingredients of Zhuk’s proof is the idea of strong/linear subagebras that exist in every finite algebra with a WNU term operation. We may assume that the domain of each variable $x$ in (*) is a subset (subuniverse) $D_{x}$ of $A$, and each $D_{x}$ has a strong/linear subset. We prove the existence of a solution (or some properties) of the instance by gradually reducing the domains $D_{x}$ of the variables to such strong subsets until all the domains are singletons.
The crucial disadvantage of this approach is that the linear subalgebras we obtain only exist locally, whereas the properties we want to prove are global. For example, we could start with a domain $D_{x}=\{0,1,\dots,99\}$ for some variable $x$. We gradually reduce this domain $D_{x}\supsetneq D_{x}^{(1)}\supsetneq D_{x}^{(2)}\supsetneq\dots\supsetneq D_{x}^{(10)}=\{0,1\}$, and on $\{0,1\}$ our instance is just a system of linear equations modulo 2. Nevertheless, the strong properties we have on $\{0,1\}$ do not say much about the behaviour on the whole domain $\{0,1,\dots,99\}$. As a result, we are forced to go forward and backward from global $\{0,1,\dots,99\}$ to local $\{0,1\}$, and use a very complicated induction to prove most of the claims.
In this paper, we develop a new theory such that every reduction is either strong or global. Precisely, for every domain $D_{x}$ we can build a sequence $D_{x}\supsetneq D_{x}^{(1)}\supsetneq D_{x}^{(2)}\supsetneq\dots\supsetneq D_{x}^{(s)}=\{a\}$ such that for every $i\in\{1,\dots,s\}$
  1. 1.
    either $D_{x}^{(i+1)}$ is a strong subset of $D_{x}^{(i)}$,
  2. 2.
    or there exists an equivalence relation $\sigma$ on $D_{x}$ satisfying very strong properties such that $D_{x}^{(i+1)}$ is an intersection of $D_{x}^{(i)}$ with a block of this equivalence relation.
In the above example with $D_{x}^{(10)}=\{0,1\}$ we would have an equivalence relation on $\{0,1,\dots,99\}$ such that 00 and $1$ are in different blocks of this equivalence relation, and the linear behaviour on $\{0,1\}$ is due to the properties of the equivalence relation.
Another good feature of the new approach is that whenever we have such a sequence of “good” subsets $D_{s}\subsetneq D_{s-1}\subsetneq\dots\subsetneq D_{1}$, we do not care about the types of the subsets in the middle. We only need to know that such a sequence exists, which we denote by $D_{s}\lll D_{1}$.
Finally, the new theory connected the two main ideas of Zhuk’s proof: the idea of strong subalgebras and the idea of bridges and connectedness. Originally, they lived separate lifes. Using strong subalgebras and reductions we showed that all the relations have the parallelogram property, which gives us an irreducible congruence for every constraint and its variable. Then manipulating with the instance we tried to connect the congruences (variables) with bridges. In the new theory, bridges appear naturally from strong/linear subuniverses: whenever a restriction to strong/linear subuniverses gives an empty set, it immediately gives a bridge between congruences (equivalence relations) defining these subuniverses.
Using the new theory we obtain two results presented in the next two subsections.
1.1 A simplified proof of the CSP Dichotomy Conjecture
First, we provide a new proof of the correctness of Zhuk’s algorithm. Three main statements that imply the correctness are formulated in Section ”Correctness of the algorithm“ in [33]. Below we formulate informal analogues of these statements, and the formal statements can be found in Section 3.4.
Informal Claim 1.
Suppose $\Gamma$ is a constraint language preserved by a WNU operation $w$. Then each $D_{x}$ of size at least 2 has a strong subset (subuniverse) or an equivalence relation $\sigma$ such that $D_{x}/\sigma\cong\mathbb{Z}_{p}$ for some prime $p$.
Informal Claim 2.
Suppose
  1. 1.
    $\Gamma$ is a constraint language preserved by a WNU operation $w$;
  2. 2.
    $\mathcal{I}$ is a consistent enough (cycle-consistent + irreducible) instance of $\operatorname{CSP}(\Gamma)$;
  3. 3.
    $\mathcal{I}$ has a solution;
  4. 4.
    $B$ is a strong subset of $D_{x}$, where $x$ is a variable of $\mathcal{I}$.
Then $\mathcal{I}$ has a solution with $x\in B$.
Informal Claim 3.
Suppose
  1. 1.
    $\Gamma$ is a constraint language preserved by a WNU operation $w$;
  2. 2.
    $\mathcal{I}$ is a consistent enough (cycle-consistent + irreducible + another one) instance of $\operatorname{CSP}(\Gamma)$ with variables $x_{1},\dots,x_{n}$;
  3. 3.
    $D_{x_{i}}$ has no strong subsets for any $i$;
  4. 4.
    $\mathcal{I}$ is linked, i.e. the following graph is connected: the vertices are all pairs $(x_{i},a)$, where $a\in D_{x_{i}}$, and vertices $(x_{i},a)$ and $(x_{j},b)$ are adjacent whenever there is a constraint in $\mathcal{I}$ whose projection onto $x_{i},x_{j}$ contains $(a,b)$;
  5. 5.
    $\sigma_{x_{i}}$ is the minimal equivalence relation on each $D_{x_{i}}$ such that $D_{x_{i}}/\sigma_{x_{i}}\cong\mathbb{Z}_{q_{1}}\times\dots\times\mathbb{Z}_{q_{n_{i}}}$;
  6. 6.
    $\varphi:\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to D_{x_{1}}/\sigma_{x_{1}}\times\dots\times D_{x_{n}}/\sigma_{x_{n}}$ is a linear map;
  7. 7.
    if we remove any constraint from $\mathcal{I}$ then the obtained instance has a solution inside $\varphi(\alpha)$ for every $\alpha\in\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$.
Then $\{(a_{1},\dots,a_{m})\mid\mathcal{I}\text{ has a solution in }\varphi(a_{1},\dots,a_{m})\}$ is either empty, or full, or an affine subspace of $\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ of dimension $m-1$.
Let us explain how Zhuk’s algorithm works (for the precise algorithm see [33, 32]). The main function Solve takes a CSP instance $\mathcal{I}$ with variables $x_{1},\dots,x_{n}$ as an input (see the pseudocode). First, it forces a sufficient level of consistency by function ForceConsistency. If we cannot achieve this, then the instance has no solutions, and we answer “No”. Then, if there exists a strong subset $B$ of the domain $D_{x_{i}}$ of some variable $x_{i}$, it reduces the domain of $x_{i}$ to $B$ by ReduceDomain, and forces the consistency again. This procedure is justified by Informal Claim 2, which guarantees that we cannot lose all the solutions when reduce to a strong subset. If there are no strong subsets, then Informal Claim 1 implies that for every domain $D_{x_{i}}$ of size at least 2 there exists an equivalence relation $\sigma_{x_{i}}$ such that $D_{x_{i}}/\sigma_{x_{i}}\cong\mathbb{Z}_{q_{1}}\times\dots\times\mathbb{Z}_{q_{n_{i}}}$ for some $n_{i}\geqslant 1$. Choose $\sigma_{x_{i}}$ to be minimal and therefore $n_{i}$ to be maximal for every $i$. This case is solved by a separate function SolveLinear.
1:function $\mathcal{I}$)
2:    repeat
3:         $\mathcal{I}:=\mbox{{ForceConsistency}}(\mathcal{I})$
4:         if $\mathcal{I}=false$ then return “No”
5:         if $D_{x_{i}}$ has a strong subset $B$ then
6:             $\mathcal{I}:=\mbox{{ReduceDomain}}(\mathcal{I},x_{i},B)$
7:    until nothing changed
8:    return $\mbox{{SolveLinear}}(\mathcal{I})$
Let $\varphi:\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to D_{x_{1}}/\sigma_{x_{1}}\times\dots\times D_{x_{n}}/\sigma_{x_{n}}$ be a linear map. By $\varphi^{-1}(\mathcal{I})$ we denote $\{\alpha\in\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\mid\mathcal{I}\text{ has a solution in }\varphi(\alpha)\}.$ Calculating $\varphi^{-1}(\mathcal{I})$ would solve the instance $\mathcal{I}$ because $\mathcal{I}$ has a solution if and only if $\varphi^{-1}(\mathcal{I})$ is not empty. We do not know how to calculate $\varphi^{-1}(\mathcal{I})$ but we can do the following calculations:
  1. (p0)
    For a concrete $\alpha\in\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ check whether $\varphi^{-1}(\mathcal{I})$ contains $\alpha$:
    1. (a)
      reduce each domain $D_{x_{i}}$ to the $i$-th element of $\varphi(\alpha)$, which is a block of $\sigma_{x_{i}}$, and solve CSP on a smaller domain by recursion.
  2. (p1)
    Check whether $\varphi^{-1}(\mathcal{I})=\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$:
    1. (a)
      using (p0) check that $(0,\dots,0)\in\varphi^{-1}(\mathcal{I})$;
    2. (b)
      using (p0) check that $(\underbrace{0,\dots,0,1}_{i},0,\dots,0)\in\varphi^{-1}(\mathcal{I})$ for every $i$.
  3. (p2)
    Calculate $\varphi^{-1}(\mathcal{I})$ if $\mathcal{I}$ is not linked (see condition 4 in Informal Claim 3):
    1. (a)
      split $\mathcal{I}$ into linked instances $\mathcal{I}_{1},\dots,\mathcal{I}_{r}$ on smaller domains;
    2. (b)
      using recursion calculate $\varphi^{-1}(\mathcal{I}_{i})$ for every $i$;
    3. (c)
      $\varphi^{-1}(\mathcal{I})=\varphi^{-1}(\mathcal{I}_{1})\cup\dots\cup\varphi^{-1}(\mathcal{I}_{r})$.
  4. (p3)
    Calculate $\varphi^{-1}(\mathcal{I})$ if the dimension of $\varphi^{-1}(\mathcal{I})$ is $m-1$ or $\varphi^{-1}(\mathcal{I})$ is empty.
    1. (a)
      using (p1) find $(a_{1},\dots,a_{m})\notin\varphi^{-1}(\mathcal{I})$;
    2. (b)
      for every $i$ using (p0) find $b_{i}$ such that $(a_{1},\dots,a_{i-1},b_{i},a_{i+1},\dots,a_{m})\in\varphi^{-1}(\mathcal{I})$ if it exists;
    3. (c)
      let $J$ be the set of all $i\in\{1,2,\dots,m\}$ such that $b_{i}$ exists;
    4. (d)
      the equation defining $\varphi^{-1}(\mathcal{I})$ is $\sum_{i\in J}(y_{i}-a_{i})/(b_{i}-a_{i})=1$.  [^1] The fact that different variables $y_{i}$ take on values from different fields $\mathbb{Z}_{p_{i}}$ is not a problem as $J$ may contain only variables on the same field.
1:function $\mathcal{I}$)
2:    $m:=$ the dimension of $D_{x_{1}}/\sigma_{x_{1}}\times\dots\times D_{x_{n}}/\sigma_{x_{n}}$triangleright $m\geqslant 0$
3:    $\varphi:=$ a bijective linear map $\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to D_{x_{1}}/\sigma_{x_{1}}\times\dots\times D_{x_{n}}/\sigma_{x_{n}}$
4:    while $\varphi^{-1}(\mathcal{I})\neq\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ do $\triangleright$ using (p1)
5:         $\mathcal{I}^{\prime}:=\mathcal{I}$
6:         for $C\in\mathcal{I}^{\prime}$ do $\triangleright$ remove unnecessary constraint from $\mathcal{I}^{\prime}$
7:             if $\varphi^{-1}(\mathcal{I}^{\prime}\setminus\{C\})\neq\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ then $\triangleright$ using (p1)
8:                 :=$\mathcal{I}^{\prime}:=\mathcal{I}^{\prime}\setminus\{C\}$
9:         $F:=\varphi^{-1}(\mathcal{I}^{\prime})$triangleright using (p2) or (p3)
10:         if $F=\varnothing$ then return “No”
11:         $m:=$ the dimension of $F$
12:         $\psi:=$ a bijective linear map $\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to F$$\triangleright$ $p_{1},\dots,p_{m}$ are also updated
13:         $\varphi:=\varphi\circ\psi$
14:    return “Yes”
The function SolveLinear solving the remaining case works as follows (see the pseudocode). We start with a bijective linear map $\varphi:\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to D_{x_{1}}/\sigma_{x_{1}}\times\dots\times D_{x_{n}}/\sigma_{x_{n}}$. We gradually reduce the dimension $m$ maintaining the property that $\operatorname{Im}{\varphi}$ contains all the solutions of $\mathcal{I}$. We stop when $\varphi^{-1}(\mathcal{I})=\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ or $\varphi^{-1}(\mathcal{I})$ is empty. First, we make a copy $\mathcal{I}^{\prime}$ of the instance $\mathcal{I}$ and remove all the constraints from $\mathcal{I}^{\prime}$ that can be removed so that $\mathcal{I}^{\prime}$ maintains the property $\varphi^{-1}(\mathcal{I}^{\prime})\neq\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$. This property can be checked in polynomial time using (p1). If we cannot remove any other constraint, by Informal Claim 3 we have one of the following cases: either $\mathcal{I}^{\prime}$ is not linked and we can calculate $\varphi^{-1}(\mathcal{I}^{\prime})$ using (p2); or $\mathcal{I}^{\prime}$ satisfies all the conditions of Informal Claim 3, and $\varphi^{-1}(\mathcal{I}^{\prime})$ has dimension $m-1$ or is empty. In the second case we can calculate $\varphi^{-1}(\mathcal{I}^{\prime})$ using (p3). Since $\mathcal{I}^{\prime}$ was obtained from $\mathcal{I}$ by removing some constraints, we have $\varphi^{-1}(\mathcal{I})\subseteq\varphi^{-1}(\mathcal{I}^{\prime})$. Thus, we found a smaller affine subspace $\varphi^{-1}(\mathcal{I}^{\prime})$ that still covers all the solutions of $\mathcal{I}$. It remains to replace $m$ with the dimension of $\varphi^{-1}(\mathcal{I}^{\prime})$ and update the linear map $\varphi$. Since we cannot reduce the dimension $m$ forever, we will eventually stop in one of the two cases: $\varphi^{-1}(\mathcal{I})=\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ or $\varphi^{-1}(\mathcal{I})$ is empty. First of them implies the existence of a solution for $\mathcal{I}$, the second implies that no solutions exist.
1.2 Existence of XY-symmetric operations
The second main result of the paper is a proof of the fact that the existence of a WNU term operation (polymorphism) implies the existence of a much stronger term operation (polymorphism).
An $n$-ary operation $f$ is called symmetric on a tuple of variables $(x_{i_{1}},\dots,x_{i_{n}})$ if it satisfies the identity $f(x_{i_{1}},\dots,x_{i_{n}})=f(x_{i_{\sigma(1)}},\dots,x_{i_{\sigma(n)}})$ for every permutation $\sigma$ on $\{1,2,\dots,n\}$. For instance, an operation $f$ is symmetric on $(x,\dots,x,y)$ if and only if $f$ is a WNU operation. An operation is called XY-symmetric if it is symmetric on $(\underbrace{x,\dots,x}_{i},y,\dots,y)$ for any $i$. An operation $f$ is called idempotent if $f(x,x,\dots,x)=x$.
As it follows from the definition, an XY-symmetric operation satisfies much more identities than a WNU operation. Nevertheless, we managed to prove that an XY-symmetric operation can always be derived from a WNU operation. To formulate the precise statement we will need a definition of a clone. A set of operations is called a clone if it is closed under composition and contains all projections. For a set of operations $M$ by $\operatorname{Clo}(M)$ we denote the clone generated by $M$.
Theorem 2.
Suppose $f$ is a WNU operation of an odd arity $n$ on a finite set. Then there exists an XY-symmetric operation $f^{\prime}\in\operatorname{Clo}(\{f\})$ of arity $n$.
Theorem 2 extends known characterization of finite Taylor algebras.
Corollary 3.
Suppose $\mathbf{A}$ is a finite idempotent algebra. Then the following conditions are equivalent:
  1. 1.
    $\mathbf{A}$ is a Taylor algebra (satisfies nontrivial identities);
  2. 2.
    there does not exist an algebra $\mathbf{B}\in\mathrm{HS}(\mathbf{A})$ of size 2 whose operations are projections [13];
  3. 3.
    $\mathbf{A}$ has a WNU term operation of any prime arity $p>|A|$ [29];
  4. 4.
    $\mathbf{A}$ has a cyclic term operation of any prime arity $p>|A|$, i.e. an operation $c_{p}$ satisfying
[…]
  5. 5.
    $\mathbf{A}$ has a Siggers term operation, i.e an operation $f$ satisfying
[…]
  6. 6.
    $\mathbf{A}$ has an XY-symmetric term operation of any prime arity $p>|A|$.
Composing a cyclic operation $c_{n}$ of arity $n$ and an XY-symmetric operation $f$ of arity $n$ we can get an operation which is simultaneously cyclic and XY-symmetric:
$f^{\prime}(x_{1},\dots,x_{n}):=c_{n}(f(x_{1},\dots,x_{n}),f(x_{2},\dots,x_{n},x_{1}),\dots,f(x_{n},x_{1},\dots,x_{n})).$
Hence, conditions 4 and 6 of Corollary 3 give an infinite sequence of cyclic XY-symmetric operations, which are the most symmetric operations known to be in every finite Taylor algebra.
This result cannot be generalized to XYZ-symmetric operations as witnessed by the following operation on $\{0,1,2\}$:
$f(x_{1},x_{2},x_{3})=\begin{cases}x_{1}+x_{2}+x_{3}\;(\mathrm{mod}\;\;2),&\text{if $x_{1},x_{2},x_{3}\in\{0,1\}$}\\ 2,&\text{if $x_{1}=x_{2}=x_{3}=2$}\\ \text{first element different from 2 in $x_{1},x_{2},x_{3}$},&\text{otherwise}\end{cases}$
Note that the clone generated by $f$ is a minimal Taylor clone and its operations were completely described in [22]. The following lemma shows that even if we try to generalize XY-symmetric operations to some tuples with $x$, $y$, and $z$ we fail.
Lemma 4.
$\operatorname{Clo}(\{f\})$ has a WNU operation of any odd arity but $\operatorname{Clo}(\{f\})$ has no operation that is symmetric on $(\underbrace{x,\dots,x}_{k},\underbrace{y,\dots,y}_{\ell},\underbrace{z,\dots,z}_{j})$ for some $k,\ell,j\geqslant 1$.
Proof.
The operation $f$ is conservative (always returns one of the coordinates), and behaves as a linear sum modulo 2 on $\{0,1\}$, as $\mathrm{min}$ on $\{0,2\}$ and $\{1,2\}$. Let us show how to derive a WNU (and even XY-symmetric) operation of any odd arity. Put
$f_{3}:=f,\;\;\;\;\;\;\;\;\;f_{2n+1}(x_{1},\dots,x_{2n+1}):=f_{2n-1}(f(x_{1},x_{2},x_{3}),x_{4},\dots,x_{2n+1}).$
It follows immediately from the definition that $f_{2n+1}$ is symmetric on $\{0,1\}$, on $\{0,2\}$, and on $\{1,2\}$.
Also, $f$ has the following properties. Whenever we substitute 00 or $1$ in $f$ we obtain either 00, or $1$. Whenever we substitute 2 into some arguments, we get an operation whose restriction to $\{0,1\}$ is a linear operation (in fact a projection). Finally, the operation preserves the sets $\{0,2\}$ and $\{1,2\}$. These three properties imply that if we put 2 to some arguments of a term operation $g\in\operatorname{Clo}(\{f\})$ and restrict the obtained operation to $\{0,1\}$ we get an idempotent linear operation on $\{0,1\}$.
Assume that $g$ is symmetric on $(\underbrace{x,\dots,x}_{k},\underbrace{y,\dots,y}_{\ell},\underbrace{z,\dots,z}_{j})$ for some $k,\ell,j\geqslant 1$. Without loss of generality assume that $j$ is odd and the first variable of $g$ is not dummy. Substituting 2 for the last $j$ coordinates of $g$ and restricting the obtained operation to $\{0,1\}$ we must get an idempotent linear operation $h$ of an even arity $k+\ell$. Since $h$ must return the same value on all the tuples with $k$ 0s and $\ell$ 1s, all the variables of $h$ are not dummy. Since $h$ has even number of arguments, it cannot be idempotent, which gives a contradiction and completes the proof. ∎
Notice that both Zhuk’s and Bulatov’s algorithms for the CSP are not universal in the sense that the algorithms work only if the domain is fixed and, therefore, all the algebraic properties are known. It would be great to find a universal algorithm for all tractable CSPs. Recently the importance of symmetric operations was rediscovered while studying the limits of universal algorithms for the CSP and its variation, called the Promise CSP [17, 2, 12]. For instance, the algorithm known as BLP+AIP solves $\operatorname{CSP}(\Gamma)$ if and only if $\Gamma$ has infinitely many symmetric polymorphisms [12].
We believe that Theorem 2 can be further generalized, and finally we will get enough symmetric operations to make some universal algorithm work. Theorem 2 already gives us some implications that can be viewed as a tiny step in this direction:
  • $\operatorname{CSP}(\Gamma)$ is solvable by BLP+AIP for any multi-sorted language $\Gamma$ on a two-element domain.
  • if $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is solvable by reducing to a tractable $\operatorname{CSP}(\mathbb{C})$, where $\mathbb{A}\to\mathbb{C}\to\mathbb{B}$, $\mathbb{C}$ is finite, and $|A|=2$, then $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is solvable by BLP+AIP (see [26, 2] for more information about Promise CSP).
To show the second claim we apply Theorem 2 to $\operatorname{Pol}(\mathbb{C})$ and obtain infinitely many XY-symmetric operations on $C$. Composing them with the homomorphisms $\mathbb{A}\to\mathbb{C}$ and $\mathbb{C}\to\mathbb{B}$ we obtain infinitely many symmetric polymorphisms $\mathbb{A}\to\mathbb{B}$, which implies that BLP+AIP solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ [12].
One of the reasons why these two independent results (proof of the CSP Dichotomy Conjecture and the existence of an XY-symmetric operation) appeared in one paper is that their proofs have the same flavour. Even though, the second result has a purely algebraic formulation, it is strongly connected to the CSP. Let us consider the matrix whose rows are all the tuples of length $n$ having exactly two different elements. We apply a WNU operation to columns of this matrix coordinate-wise deriving new columns till we cannot derive anything new. The set of all the derived columns can be viewed as a relation $R$ of some big arity $N$. To prove that an XY-symmetric operation can be derived from a WNU we need to show that $R$ contains a tuple whose elements corresponding to permutations of the same tuple are equal. This can be written as a CSP instance with the constraint $R(x_{1},\dots,x_{N})$ and many equality constraints $(x_{i}=x_{j})$, and we need to prove that it has a solution. Then the proofs of Informal Claims 2 and 3 are similar to the proof of Theorem 2, only sufficient level of consistency is replaced by symmetries of the relation $R$.
1.3 History and acknowledgements
The first symmetric operations (WNU) that exist in every Taylor algebra appeared in [29], and the idea was to show that every symmetric invariant relation has a constant tuple. In [34] I showed the existence of a constant tuple in a symmetric relation gradually reducing the domain to strong subalgebras and keeping the property that the relation is symmetric. It turned out that the only reason why a constant tuple does not exist in a symmetric relation is a linear essence inside, for instance the relation $x_{1}+\dots+x_{p}=1$ does not have a constant tuple in $\mathbb{Z}_{p}$. Thus, the existence of a WNU term operation of an arity $n$ is reduced to a pure linear algebra question: does every affine symmetric subspace of $\mathbb{Z}_{p}^{n}$ have a constant tuple. Similarly, we could try to show the existence of a 2-WNU operation (symmetric on $(x,x,y,y,\dots,y)$). For the proof to work we need to show that every relation of arity (${n\choose 2}$ with symmetries coming from $n$-permutations has a constant tuple. This question is again reduced to a pure linear algebra question: does every affine (weak) symmetric subspace of $\mathbb{Z}_{p}^{n\choose 2}$ have a constant tuple. We worked on it together with Libor Barto, Michael Pinsker, and their students Johanna Brunar and Martin Boroš. Martin Boroš in his master thesis [9] proved that a constant tuple always exists if and only if $n\cdot{n\choose 2}$ is co-prime with $p$. Unfortunately this beautiful approach did not lead to a proof of existence of 2-WNU or XY-symmetric operations as it turned out that an XY-symmetric operation may exist even if the condition on the arity is not satisfied. Nevertheless, I am very thankful to Libor Barto and Michael Pinsker for the exciting play with a beautiful linear algebra and the ideas I took from this play.
I would also like to thank Stanislav Živný, Lorenzo Ciardo, and Tamio-Vesa Nakajima for very fruitful discussions about algorithms for the CSP based on linear programming and their limits. My understanding of what operations we need for the algorithms to work came to me during my visit of Oxford University. The impotence of symmetric operations for these algorithms motivated me to finish my research on XY-symmetric operations.
1.4 Structure of the paper
In Section 2 we give definitions and statements of the new theory of strong subalgebras. In Section 3 we use this theory to prove that the algorithm for the CSP presented in [33, 32] works. In Section 4 we show that an XY-symmetric operation can be derived from a WNU operation of an odd arity. In Section 5 we prove all the statements formulated in Section 2.
The main goal of the paper is to show the power of the new theory of strong subalgebras but not to provide a shortest proof of the CSP Dichotomy Conjecture. That is why, we formulate and prove all the properties of strong/linear subalgebras for arbitrary finite idempotent algebras, even though in Sections 3 and 4 we only consider algebras with a special WNU operation. Moreover, many definitions and statements could be simplified if we consider only Taylor minimal algebras (see [1]), which would be sufficient to prove two main results of this paper. Also, for better readability we always duplicate statements if the proof appears in a later section. For instance, all the statements from Section 2 are formulated again in Section 5.
2 Strong/Linear subuniverses
In this section we define six types of subuniverses and formulate all the necessary properties of these subuniverses. We start with auxiliary definitions and notations, then we define two types of irreducible congruences, and introduce notations for all types of subuniverses. In Subsection 2.3 we give their properties without a proof. We conclude the section with a few auxiliary statements.
2.1 Auxiliary definitions
For a positive integer $k$ by $[k]$ we denote the set $\{1,2,\dots,k\}$. An idempotent WNU $w$ is called special if
$w(x,\dots,x,y)=w(x,\dots,x,w(x,\dots,x,y)).$
It is not hard to show that for any idempotent WNU $w$ on a finite set there exists a special WNU $w^{\prime}\in\operatorname{Clo}(w)$ (see Lemma 26).
Algebras. We denote algebras by bold letters $\mathbf{A},\mathbf{B},\mathbf{C},\dots$, their domains by $A,B,C,\dots$, and the basic operations by $f^{\mathbf{A}},f^{\mathbf{B}},g^{\mathbf{C}},\dots$. We use standard universal algebraic notions of term operation, subalgebra, factor algebra, product of algebras, see [6]. We write $\mathbf{B}\leqslant\mathbf{A}$ if $\mathbf{B}$ is a subalgebra of $\mathbf{A}$. A congruence is called nontrivial if it is not the equality relation and not $A^{2}$. By $0_{\mathbf{A}}$ we denote the equality relation on $A$, which is the 0-congruence on $\mathbf{A}$. To avoid overusing of bold symbols sometimes we write capital symbol meaning the algebra. An algebra $(A;F_{A})$ is called polynomially complete (PC) if the clone generated by $F_{A}$ and all constants on $A$ is the clone of all operations on $A$ (see [21, 27]).
By $\mathcal{V}_{n}$ we denote the class of finite algebras $\mathbf{A}=(A;w^{\mathbf{A}})$ whose basic operation $w^{\mathbf{A}}$ is an idempotent special WNU operation. Since we only consider finite algebras, $\mathcal{V}_{n}$ is not a variety. For a prime $p$ by $\mathbf{Z}_{p}$ we denote the algebra whose domain is $\{0,1,\dots,p-1\}$ and whose basic operation $w^{\mathbf{Z}_{p}}$ is $x_{1}+\dots+x_{n}(mod\;\;p)$. In the paper every algebra $\mathbf{Z}_{p}$ belongs to $\mathcal{V}_{n}$ for a fixed $n$, hence the algebra $\mathbf{Z}_{p}$ is uniquely defined. In this paper we assume that every algebra is a finite idempotent algebra having a WNU term operation. Moreover, in Sections 3 and 4 we usually consider algebras from $\mathcal{V}_{n}$.
Notations. A relation $R\subseteq A_{1}\times\dots\times A_{n}$ is called subdirect if for every $i$ the projection of $R$ onto the $i$-th coordinate is $A_{i}$. A relation $R\subseteq A^{n}$ is called reflexive if it contains $(a,a,\dots,a)$ for every $a\in A$. For a relation $R$ by $\operatorname{pr}_{i_{1},\ldots,i_{s}}(R)$ we denote the projection of $R$ onto the coordinates $i_{1},\ldots,i_{s}$. We write $\mathbf{R}\leqslant_{sd}\mathbf{A}_{1}\times\dots\times\mathbf{A}_{m}$ (say that $R$ is a subdirect subalgebra) if $R$ is a subdirect relation and $\mathbf{R}\leqslant\mathbf{A}_{1}\times\dots\times\mathbf{A}_{m}$. For $R\subseteq A^{n}$ by $\operatorname{Sg}_{\mathbf{A}}(R)$ we denote the minimal subalgebra of $\mathbf{A}^{n}$ containing $R$, that is the subalgebra of $\mathbf{A}^{n}$ generated from $R$.
For an equivalence relation $\sigma$ on $A$ and $a\in A$ by $a/\sigma$ we denote the equivalence class containing $a$. For an equivalence relation $\sigma$ on $A$ and $B\subseteq A$ denote $B/\sigma=\{b/\sigma\mid b\in B\}$. Similarly, for a relation $R\subseteq A^{n}$ denote $R/\sigma=\{(b_{1}/\sigma,\dots,b_{n}/\sigma)\mid(b_{1},\dots,b_{n})\in R\}$. For a binary relation $\sigma$ and $n\geqslant 2$ by $\sigma^{[n]}$ we denote the $n$-ary relation $\{(a_{1},\dots,a_{n})\mid\forall i,j\in[n]\colon(a_{i},a_{j})\in\sigma\}$.
For two binary relations $\delta_{1}\subseteq A_{1}\times A_{2}$ and $\delta_{2}\subseteq A_{2}\times A_{3}$ by $\delta_{1}\circ\delta_{2}$ we denote the binary relation $\{(a,b)\mid\exists c\colon(a,c)\in\delta_{1}\wedge(c,b)\in\delta_{2}\}$. Similarly, for $B\subseteq A_{1}$ and $\delta\subseteq A_{1}\times A_{2}$ put $B\circ\delta=\{c\mid\exists b\in B\colon(b,c)\in\delta\}$. For a binary relation $\delta$, we denote $\delta^{-1}(x,y)=\delta(y,x)$. A binary subdirect relation $\delta\subseteq A\times B$ is called linked if the bipartite graph corresponding to $\delta$ is connected. A binary subdirect relation $\delta\subseteq A_{1}\times A_{2}$ is called bijective if $|\delta|=|A_{1}|=|A_{2}|$.
Parallelogram property and rectangularity. We say that an $n$-relation $R$ has the parallelogram property if any permutation of its variables gives a relation $R^{\prime}$ satisfying
$\forall\ell\in\{1,2,\dots,n-1\}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;$ $(a_{1},\dots,a_{\ell},b_{\ell+1},\dots,b_{n})\in R^{\prime}$
$\forall a_{1},\dots,a_{n},b_{1},\dots,b_{n}\colon\;\;\;\;$ $(b_{1},\dots,b_{\ell},a_{\ell+1},\dots,a_{n})\in R^{\prime}\;\;\;\Rightarrow\;\;\;(a_{1},\dots,a_{n})\in R^{\prime}.$
$(b_{1},\dots,b_{\ell},b_{\ell+1},\dots,b_{n})\in R^{\prime}$
Note that the parallelogram property plays an important role in universal algebra (see [23] for more details). We say that the $i$-th variable of a relation $R$ is rectangular, if
$(a_{1},\dots,a_{i-1},b_{i},a_{i+1},\dots,a_{n})\in R$
$\forall a_{1},\dots,a_{n},b_{1},\dots,b_{n}\colon\;\;\;\;$ $(b_{1},\dots,b_{i-1},a_{i},b_{i+1},\dots,b_{n})\in R\;\;\;\Rightarrow\;\;\;(a_{1},\dots,a_{n})\in R.$
$(b_{1},\dots,b_{i-1},b_{i},b_{i+1},\dots,b_{n})\in R$
As it follows from the definitions, if a relation has the parallelogram property then it is rectangular. The rectangular closure of a relation $R$ is the minimal rectangular relation $R^{\prime}$ containing $R$.
Irreducible congruences. For a relation $R\subseteq A_{1}\times\dots\times A_{n}$ and a congruence $\sigma$ on $A_{i}$, we say that the $i$-th variable of the relation $R$ is stable under $\sigma$ if $(a_{1},\ldots,a_{n})\in R$ and $(a_{i},b_{i})\in\sigma$ imply $(a_{1},\ldots,a_{i-1},b_{i},a_{i+1},\ldots,a_{n})\in R$. We say that a relation is stable under $\sigma$ if every variable of this relation is stable under $\sigma$. We say that a congruence $\sigma$ on $\mathbf{A}$ is irreducible if it cannot be represented as an intersection of other binary subalgebras of $\mathbf{A}\times\mathbf{A}$ that are stable under $\sigma$. Equivalently, a congruence is irreducible if there are no subalgebras $\mathbf{S}_{1},\mathbf{S}_{2},\dots,\mathbf{S}_{k}\leqslant\mathbf{A}/\sigma\times\mathbf{A}/\sigma$ such that $0_{\mathbf{A}/\sigma}=S_{1}\cap S_{2}\cap\dots\cap S_{k}$ and $0_{\mathbf{A}/\sigma}\neq S_{i}$ for every $i\in[k]$. Then for an irreducible congruence $\sigma$ on $\mathbf{A}$ by ${{\sigma}^{*}}$ we denote the minimal $\delta\leqslant\mathbf{A}\times\mathbf{A}$ such that $\delta\supsetneq\sigma$ and $\delta$ is stable under $\sigma$.
Bridges. Suppose $\sigma_{1}$ and $\sigma_{2}$ are congruences on $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$, respectively. A relation $\delta\leqslant\mathbf{D}_{1}^{2}\times\mathbf{D}_{2}^{2}$ is called a bridge from $\sigma_{1}$ to $\sigma_{2}$ if the following conditions hold:
  1. 1.
    the first two variables of $\delta$ are stable under $\sigma_{1}$,
  2. 2.
    the last two variables of $\delta$ are stable under $\sigma_{2}$,
  3. 3.
    $\operatorname{pr}_{1,2}(\delta)\supsetneq\sigma_{1}$, $\operatorname{pr}_{3,4}(\delta)\supsetneq\sigma_{2}$,
  4. 4.
    $(a_{1},a_{2},a_{3},a_{4})\in\delta$ implies $(a_{1},a_{2})\in\sigma_{1}\Leftrightarrow(a_{3},a_{4})\in\sigma_{2}.$
[…]
For a bridge $\delta$ by $\widetilde{\delta}$ we denote the binary relation defined by $\widetilde{\delta}(x,y)=\delta(x,x,y,y)$.
We can compose a bridge $\delta_{1}$ from $\sigma_{0}$ to $\sigma_{1}$ and a bridge $\delta_{2}$ from $\sigma_{1}$ to $\sigma_{2}$ using the following formula:
$\delta(x_{1},x_{2},z_{1},z_{2})=\exists y_{1}\exists y_{2}\;\delta_{1}(x_{1},x_{2},y_{1},y_{2})\wedge\delta_{2}(y_{1},y_{2},z_{1},z_{2}).$
We can prove (Lemma 28) that $\delta$ is a bridge from $\sigma_{0}$ to $\sigma_{2}$ whenever the congruences $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$ are irreducible. Moreover, $\widetilde{\delta}=\widetilde{\delta_{1}}\circ\widetilde{\delta_{2}}$.
A congruence $\sigma$ on $\mathbf{A}=(A;w)$ is called a perfect linear congruence if it is irreducible and there exists $\zeta\leqslant\mathbf{A}\times\mathbf{A}\times\mathbf{Z}_{p}$ such that $\operatorname{pr}_{1,2}\zeta=\sigma^{*}$ and $(a_{1},a_{2},b)\in\zeta$ implies that $(a_{1},a_{2})\in\sigma\Leftrightarrow(b=0)$. Such congruences are important for us because we can control relaxation of $\sigma$ to $\sigma^{*}$ by an additional element from $\mathbf{Z}_{p}$.
In our proofs we compose bridges to get a bridge $\delta$ whose binary relation $\widetilde{\delta}$ is linked and then apply the following lemma that will be proved in Subection 2.4.
Lemma 5.
Suppose $\sigma$ is a irreducible congruence on $\mathbf{A}\in\mathcal{V}_{n}$, $\delta$ is a bridge from $\sigma$ to $\sigma$ such that $\widetilde{\delta}$ is linked. Then $\sigma$ is a perfect linear congruence.
2.2 Definition of strong subuniverses
(Binary) absorbing subuniverse. We say $B$ is an absorbing subuniverse of an algebra ${\mathbf{A}}$ if there exists $t\in\operatorname{Clo}(\mathbf{A})$ such that $t(B,B,\dots,B,A,B,\dots,B)\subseteq B$ for any position of $A$. Also in this case we say that $B$ absorbs ${\mathbf{A}}$ with a term $t$.
If the operation $t$ can be chosen binary then we say that $B$ is a binary absorbing subuniverse of $\mathbf{A}$. To shorten sometimes we will write BA instead of binary absorbing. If $t$ can be chosen ternary the we call $B$ a ternary absorbing subuniverse. For more information about absorption and its connection to CSP see [5].
Central subuniverse. A subuniverse $C$ of ${\mathbf{A}}$ is called central if it is an absorbing subuniverse and for every $a\in A\setminus C$ we have $(a,a)\notin\operatorname{Sg}_{\mathbf{A}}((\{a\}\times C)\cup(C\times\{a\}))$.
Central subuniverses are strongly connected with ternary absorption.
Lemma 6 ([34], Corollary 6.11.1).
Suppose $B$ is a central subuniverse of $\mathbf{A}$, then $B$ is a ternary absorbing subuniverse of $\mathbf{A}$.
In general ternary absorption does not imply central subuniverse, but they are equivalent for minimal Taylor algebras (see [1]). We say that an algebra $\mathbf{A}$ is BA and center free if $\mathbf{A}$ has no proper nonempty binary absorbing subuniverse or proper nonempty central subuniverse.
Linear and PC congruences. There are two different types of irreducible congruences. A congruence $\sigma$ on $\mathbf{A}$ is called linear if
  1. 1.
    $\sigma$ is irreducible
  2. 2.
    $\sigma^{*}$ is a congruence
  3. 3.
    there exist prime $p$ and $S\leqslant(\sigma^{*})^{[4]}$ such that for any block $B$ of $\sigma^{*}$ there exists $n\geqslant 0$ with $(B/\sigma;S\cap(B/\sigma)^{4})\cong(\mathbb{Z}_{p}^{n};x_{1}-x_{2}=x_{3}-x_{4})$.
Notice that the relation $S$ above is a bridge from $\sigma$ to $\sigma$ such that $\widetilde{S}=\operatorname{pr}_{1,2}(S)=\operatorname{pr}_{3,4}(S)=\sigma^{*}$.
An irreducible congruence is called a PC congruence if it is not linear. Notice that a congruence $\sigma$ is an irreducible/PC/linear congruence if and only if $0_{\mathbf{A}/\sigma}$ is an irreducible/PC/Linear congruence.
Lemma 7.
Suppose $\sigma$ is an irreducible congruence on $\mathbf{A}$. Then the following conditions are equivalent:
  1. 1.
    $\sigma$ is a linear congruence;
  2. 2.
    there exists a bridge $\delta$ from $\sigma$ to $\sigma$ such that $\widetilde{\delta}\supsetneq\sigma$.
Another important fact is that there cannot be a bridge between PC and linear congruences.
Lemma 8.
Suppose $\sigma_{1}$ is a linear congruence, $\sigma_{2}$ is an irreducible congruence, $\delta$ is a bridge from $\sigma_{1}$ to $\sigma_{2}$. Then $\sigma_{2}$ is a also linear congruence.
Unlike bridges for linear congruences, bridges from PC congruences are trivial.
Lemma 9.
Suppose $\sigma$ is a PC congruence on $A$. Then any reflexive bridge $\delta$ from $\sigma$ to $\sigma$ such that $\operatorname{pr}_{1,2}(\delta)=\operatorname{pr}_{3,4}(\delta)=\sigma^{*}$ can be represented as $\delta(x_{1},x_{2},x_{3},x_{4})=\sigma(x_{1},x_{3})\wedge\sigma(x_{2},x_{4})$ or $\delta(x_{1},x_{2},x_{3},x_{4})=\sigma(x_{1},x_{4})\wedge\sigma(x_{2},x_{3})$.
Lemma 10.
Suppose $\delta$ is a bridge from a PC congruence $\sigma_{1}$ on $\mathbf{A}_{1}$ to an irreducible congruence $\sigma_{2}$ on $\mathbf{A}_{2}$, $\operatorname{pr}_{1,2}(\delta)=\sigma_{1}^{*}$, and $\operatorname{pr}_{3,4}(\delta)=\sigma_{2}^{*}$. Then
  1. 1.
    $\sigma_{2}$ is a PC congruence;
  2. 2.
    $\mathbf{A}_{1}/\sigma_{1}\cong\mathbf{A}_{2}/\sigma_{2}$;
  3. 3.
    $\{(a/\sigma_{1},b/\sigma_{2})\mid(a,b)\in\widetilde{\delta}\}$ is bijective;
  4. 4.
    $\delta(x_{1},x_{2},x_{3},x_{4})=\widetilde{\delta}(x_{1},x_{3})\wedge\widetilde{\delta}(x_{2},x_{4})$ or $\delta(x_{1},x_{2},x_{3},x_{4})=\widetilde{\delta}(x_{1},x_{4})\wedge\widetilde{\delta}(x_{2},x_{3})$.
The following claims show the connection of the new definitions with the linear and PC subuniverses from the original proof of the CSP Dichotomy Conjecture [33].
Lemma 11.
Suppose $\sigma$ is a linear congruence on $\mathbf{A}\in\mathcal{V}_{n}$ such that $\sigma^{*}=A^{2}$. Then $\mathbf{A}/\sigma\cong\mathbf{Z}_{p}$ for some prime $p$.
Lemma 12.
Suppose $\sigma$ is a PC congruence on $\mathbf{A}$ and $\sigma^{*}=A^{2}$. Then $\mathbf{A}/\sigma$ is a PC algebra.
All types of subuniverses. Suppose $\varnothing\neq C\lneq B\leqslant A$. We write
  • $C<_{\mathcal{BA}}^{A}B$ if $C$ is a BA subuniverse of $B$.
  • $C<_{\mathcal{C}}^{A}B$ if $C$ is a central subuniverse of $B$.
  • $C<_{\mathcal{D}}^{A}B$ if there exists an irreducible congruence $\sigma$ such that
    1. 1.
      $B^{2}\subseteq\sigma^{*}$;
    2. 2.
      $C=B\cap E$ for some block $E$ of $\sigma$;
    3. 3.
      $B/\sigma$ is BA and center free.
  • $C<_{\mathcal{L}}^{A}B$ if $C<_{\mathcal{D}}^{A}B$ and the congruence $\sigma$ from the definition of $<_{\mathcal{D}}^{A}$ is linear.
  • $C<_{\mathcal{PC}}^{A}B$ if $C<_{\mathcal{D}}^{A}B$ and the congruence $\sigma$ from the definition of $<_{\mathcal{D}}^{A}$ is a PC congruence.
  • $C<_{\mathcal{S}}^{A}B$ if there exists a BA and central (simultaneously) subuniverse $D$ in $\mathbf{B}$ such that $D\leqslant C$.
When we want to specify what congruence was used in the definition we write $C<_{T(\sigma)}^{A}B$. Sometimes, we also put a congruence there even if $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$, which means that $\sigma$ is a full congruence. If $C<_{T}^{A}B$ then we say that $C$ is a subuniverse of $B$ of type $T$. Sometimes we also call $B$ a dividing subuniverse for the type $\mathcal{D}$, a linear subuniverse for the type $\mathcal{L}$, and a PC subuniverse for the type $\mathcal{PC}$. Also, we say that $\sigma$ is a dividing/linear/PC congruence for $B\leqslant A$ if $C<_{\mathcal{T}(\sigma)}^{A}B$ for some $C$ and $\mathcal{T}=\mathcal{D}/\mathcal{L}/\mathcal{PC}$. We say that an algebra $\mathbf{A}$ is S-free if there is no $D\leqslant A$ such that $D<_{\mathcal{BA}}A$ and $D<_{\mathcal{C}}A$. Equivalently, an algebra $\mathbf{A}$ is S-free if there does not exist $C<_{\mathcal{S}}\mathbf{A}$.
Sometimes instead of $C<_{\mathcal{BA}}^{A}B$, $C<_{\mathcal{C}}^{A}B$, and $C<_{\mathcal{S}}^{A}B$ we write $C<_{\mathcal{BA}}B$, $C<_{\mathcal{C}}B$, and $C<_{\mathcal{S}}B$, which is justified because $A$ is irrelevant to the definition. Also, we write $C<_{\mathcal{BA},\mathcal{C}}B$ meaning that $C<_{\mathcal{BA}}B$ and $C<_{\mathcal{C}}B$.
We write $C\lll^{A}B$ if there exist $B_{0},B_{1},\dots,B_{n}\subseteq B$ and $T_{1},\dots,T_{n}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{D}\}$ such that $C=B_{n}<_{T_{n}}^{A}B_{n-1}<_{T_{n-1}}^{A}<\dots<_{T_{2}}^{A}<B_{1}<_{T_{1}}^{A}B_{0}=B$. Notice that $n$ can be 0 and the relation $\lll^{A}$ is reflexive. We say that a congruence comes from $C\lll^{A}B$ if it is one of the dividing congruences used in the sequence $C\lll^{A}B$. We usually write $B\lll A$ instead of $B\lll^{A}A$. We write $C\leqslant_{T(\sigma)}^{A}B$ if $C=B$ or $C<_{T(\sigma)}^{A}B$.
Let us introduce the types $\mathcal{ML},\mathcal{MPC},\mathcal{MD}$ of subuniverses. Suppose $T\in\{\mathcal{L},\mathcal{PC},\mathcal{D}\}$. We write $C<_{\mathcal{M}T}^{A}B$ if $C\neq\varnothing$ and $C=C_{1}\cap\dots\cap C_{t}$, where $C_{i}<_{T}^{A}B$ for every $i\in[t]$.
Notice that we do not allow empty subuniverses and the condition $\varnothing\lll A$ never holds. Nevertheless, sometimes we need to allow an empty set. In this case we add a dot above and write $B\dot{\lll}A$ meaning that $B\lll A$ or $B=\varnothing$. With the same meaning we use dots in the following notations $C\dot{<}_{T}^{A}B$ or $C\dot{\leqslant}_{T}^{A}B$.
2.3 Properties of strong subuniverses
Recall that all the algebras in the following statements are assumed finite idempotent algebras having a WNU term operation (Taylor). To avoid listing all the possible types in the following lemmas we assume that if the type $T$ is not specified then $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{PC},\mathcal{L},\mathcal{D}\}$. If we write the type $\mathcal{M}T$ then we assume that $T\in\{\mathcal{PC},\mathcal{L},\mathcal{D}\}$ and, therefore, $\mathcal{M}T\in\{\mathcal{MPC},\mathcal{ML},\mathcal{MD}\}$.
Lemma 13.
Suppose $B\lll A$ and $|B|>1$. Then there exists $C<_{T}^{A}B$, where $T\in\{\mathcal{BA},\mathcal{C},\mathcal{L},\mathcal{PC}\}$.
Lemma 14.
Suppose $f\colon\mathbf{A}\to\mathbf{A}^{\prime}$ is a surjective homomorphism, then
  1. (f)
    $C\lll^{A}B\Rightarrow f(C)\lll f(B)$;
  2. (b)
    $C^{\prime}\lll^{A^{\prime}}B^{\prime}\Rightarrow f^{-1}(C^{\prime})\lll f^{-1}(B)$;
  3. (ft)
    $C<_{T(\sigma)}^{A}B\lll A\Longrightarrow(f(C)=f(B)\text{ or }f(C)<_{S}f(B)\text{ or }f(C)<_{T}^{A^{\prime}}f(B))$;
  4. (bt)
    $C^{\prime}<_{T(\sigma)}^{A^{\prime}}B^{\prime}\Rightarrow f^{-1}(C^{\prime})<_{T(f^{-1}(\sigma))}^{A^{\prime}}f^{-1}(B^{\prime})$;
  5. (fs)
    $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$ and $C<_{T}B$ $\Longrightarrow f(C)\leqslant_{T}f(B)$;
  6. (fm)
    $C\leqslant_{\mathcal{M}T}^{A}B\lll A$ and $f(B)$ is S-free $\Longrightarrow f(C)\leqslant_{\mathcal{M}T}^{A^{\prime}}f(B)$;
  7. (bm)
    $C^{\prime}\leqslant_{\mathcal{M}T}^{A^{\prime}}B^{\prime}\lll A^{\prime}\Longrightarrow f^{-1}(C)\leqslant_{\mathcal{M}T}^{A}f^{-1}(B)$.
Corollary 15.
Suppose $\delta$ is a congruence on $\mathbf{A}$. Then
  1. (f)
    $C\lll^{A}B\Rightarrow C/\delta\lll^{A/\delta}B/\delta$;
  2. (t)
    $C<_{T(\sigma)}^{A}B\lll A\Longrightarrow(C/\delta=B/\delta\text{ or }C/\delta<_{\mathcal{S}}B/\delta\text{ or }C/\delta<_{T}^{A/\delta}B/\delta)$;
  3. (s)
    $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$ and $C<_{T}B$ $\Longrightarrow C/\delta\leqslant_{T}B/\delta$;
  4. (m)
    $C\leqslant_{\mathcal{M}T}^{A}B\lll A$ and $B/\delta$ is S-free $\Longrightarrow C/\delta\leqslant_{\mathcal{M}T}^{A/\delta}B/\delta$.
Corollary 16.
Suppose $\delta$ is a congruence on $\mathbf{A}$, $\mathbf{B},\mathbf{C}\leqslant\mathbf{A}$. Then
  1. (f)
    $C/\delta\lll^{A/\delta}B/\delta\Longleftrightarrow C\circ\delta\lll^{A}B\circ\delta$;
  2. (t)
    $C/\delta<_{T}^{A/\delta}B/\delta\Longleftrightarrow C\circ\delta<_{T}^{A}B\circ\delta$.
Corollary 17.
Suppose $\delta$ is a congruence on $\mathbf{A}$. Then
  1. (f)
    $C\lll^{A}B\Rightarrow C\circ\delta\lll^{A}B\circ\delta$;
  2. (t)
    $C<_{T(\sigma)}^{A}B\lll^{A}A\Longrightarrow(C\circ\delta=B\circ\delta\text{ or}C\circ\delta<_{\mathcal{S}}^{A}B\circ\delta\text{ or }C\circ\delta<_{T}^{A}B\circ\delta)$;
  3. (e)
    $\delta\subseteq\sigma$ and $C<_{T(\sigma)}^{A}B\lll^{A}A\Longrightarrow C\circ\delta<_{T}^{A}B\circ\delta$.
Corollary 18.
Suppose $R\leqslant_{sd}A_{1}\times\dots\times A_{n}$, $B_{i}\lll A_{i}$ for $i\in[n]$. Then
  1. (r)
    $R\cap(B_{1}\times\dots\times B_{n}))\dot{\lll}R$;
  2. (r1)
    $\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))\dot{\lll}A_{1}$;
  3. (b)
    $\forall i\colon C_{i}\lll^{A_{i}}B_{i}\Longrightarrow(R\cap(C_{1}\times\dots\times C_{n}))\dot{\lll}^{R}(R\cap(B_{1}\times\dots\times B_{n}))$;
  4. (b1)
    $\forall i\colon C_{i}\lll^{A_{i}}B_{i}\Longrightarrow\operatorname{pr}_{1}(R\cap(C_{1}\times\dots\times C_{n}))\dot{\lll}^{A_{1}}\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))$;
  5. (m)
    $\forall i\colon C_{i}\leqslant_{\mathcal{M}T}^{A_{i}}B_{i}\Longrightarrow R\cap(C_{1}\times\dots\times C_{n})\dot{\leqslant}_{\mathcal{M}T}^{R}R\cap(B_{1}\times\dots\times B_{n})$;
  6. (m1)
    $\forall i\colon C_{i}\leqslant_{\mathcal{M}T}^{A_{i}}B_{i}$, $\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))$ is S-free $\Longrightarrow$
    $\operatorname{pr}_{1}(R\cap(C_{1}\times\dots\times C_{n}))\dot{\leqslant}_{\mathcal{M}T}^{A_{1}}\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))$.
For binary absorbing and central subuniverses we can prove a stronger claim.
Lemma 19 ([34], Corollaries 6.1.2 and 6.9.2).
Suppose $R\leqslant A_{1}\times\dots\times A_{n}$, $C_{i}\leqslant_{T}A_{i}$ for every $i\in[n]$, where $T\in\{\mathcal{BA},\mathcal{C}\}$. Then $\operatorname{pr}_{1}(R\cap(C_{1}\times\dots\times C_{N}))\dot{\leqslant}_{T}A_{1}$.
Lemma 20.
Suppose $B\lll A$, $D\lll A$. Then
  1. (i)
    $B\cap D\dot{\lll}A$;
  2. (t)
    $C<_{T(\sigma)}^{A}B\Rightarrow C\cap D\dot{\leqslant}_{T(\sigma)}^{A}B\cap D$.
Theorem 21.
Suppose
  1. 1.
    $C_{i}<_{T_{i}(\sigma_{i})}^{A}B_{i}\lll A$, where $T_{i}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{PC}\}$ for $i=1,2,\dots,n$, $n\geqslant 2$;
  2. 2.
    $\bigcap\limits_{i\in[n]}C_{i}=\varnothing$;
  3. 3.
    $B_{j}\cap\bigcap\limits_{i\in[n]\setminus\{j\}}C_{i}\neq\varnothing$ for every $j\in[n]$.
Then one of the following conditions hold:
  1. (ba)
    $T_{1}=\dots=T_{n}=\mathcal{BA}$;
  2. (l)
    $T_{1}=\dots=T_{n}=\mathcal{L}$ and for every $k,\ell\in[n]$ there exists a bridge $\delta$ from $\sigma_{k}$ and $\sigma_{\ell}$ such that $\widetilde{\delta}=\sigma_{k}\circ\sigma_{\ell}$;
  3. (c)
    $n=2$ and $T_{1}=T_{2}=\mathcal{C}$;
  4. (pc)
    $n=2$, $T_{1}=T_{2}=\mathcal{PC}$, and $\sigma_{1}=\sigma_{2}$.
Corollary 22.
Suppose
  1. 1.
    $R\leqslant_{sd}A_{1}\times\dots\times A_{n}$;
  2. 2.
    $C_{i}<_{T_{i}(\sigma_{i})}^{A_{i}}B_{i}\lll A_{i}$, where $T_{i}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{PC}\}$ for $i=1,2,\dots,n$, $n\geqslant 2$;
  3. 3.
    $R\cap(C_{1}\times\dots\times C_{n})=\varnothing$;
  4. 4.
    $R\cap(C_{1}\times\dots\times C_{j-1}\times B_{j}\times C_{j+1}\times\dots\times C_{n})\neq\varnothing$ for every $j\in[n]$.
Then one of the following conditions hold:
  1. (ba)
    $T_{1}=\dots=T_{n}=\mathcal{BA}$;
  2. (l)
    $T_{1}=\dots=T_{n}=\mathcal{L}$ and for every $k,\ell\in[n]$ there exists a bridge $\delta$ from $\sigma_{k}$ and $\sigma_{\ell}$ such that $\widetilde{\delta}=\sigma_{k}\circ\operatorname{pr}_{k,\ell}(R)\circ\sigma_{\ell}$;
  3. (c)
    $n=2$ and $T_{1}=T_{2}=\mathcal{C}$;
  4. (pc)
    $n=2$, $T_{1}=T_{2}=\mathcal{PC}$, $A_{1}/\sigma_{1}\cong A_{2}/\sigma_{2}$, and the relation $\{(a/\sigma_{1},b/\sigma_{2})\mid(a,b)\in R\}$ is bijective.
Remark 1.
Notice that sometimes we want to have several restrictions on one coordinate of a relation. To keep the statement of Corollary 22 simple we do not add this possibility into the claim, but we can always duplicate the coordinate of the relation and apply restrictions separately on different coordinates.
Lemma 23.
Suppose $C\leqslant_{\mathcal{M}T}^{A}B$. Then $C<_{T}^{A}\dots<_{T}^{A}B$ and $C\lll^{A}B$.
Lemma 24.
Suppose $R\leqslant_{sd}\mathbf{A}_{1}\times\mathbf{A}_{2}$, $C_{i}\leqslant_{\mathcal{\mathcal{MD}}}^{A_{i}}B_{i}\lll A_{i}$ for $i\in\{1,2\}$, $S$ is a rectangular closure of $R$, $R\cap(B_{1}\times B_{2})\neq\varnothing$, $S\cap(C_{1}\times C_{2})\neq\varnothing$. Then $R\cap(C_{1}\times C_{2})\neq\varnothing$.
Lemma 25.
Suppose $C_{1}<_{\mathcal{M}T}^{A}B_{1}\lll A$, $B_{2}\lll A$, $C_{1}\cap B_{2}=\varnothing$, $B_{1}\cap B_{2}\neq\varnothing$, $\sigma$ is a maximal congruence on $\mathbf{A}$ such that $(C_{1}\circ\sigma)\cap B_{2}=\varnothing$. Then $\sigma=\omega_{1}\cap\dots\cap\omega_{s}$, where $\omega_{1},\dots,\omega_{s}$ are congruences of type $T$ on $\mathbf{A}$ such that $\omega_{i}^{*}\supseteq B_{1}^{2}$.
2.4 Auxiliary Statements
Lemma 26 ([28] Lemma 4.7).
Suppose $w$ is an idempotent WNU operation on $A$. Then there exists a special idempotent WNU operation $w^{\prime}\in\operatorname{Clo}(w)$ of arity $n^{n!}$.
Lemma 27 ([33] Corollary 8.17.1).
Suppose $\sigma$ is an irreducible congruence on $\mathbf{A}\in\mathcal{V}_{n}$ and $\delta$ is a bridge from $\sigma$ to $\sigma$ such that $\widetilde{\delta}=A^{2}$. Then $\sigma$ is a perfect linear congruence.
Lemma 28 ([33] Lemma 6.3).
Suppose $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ are irreducible congruences, $\rho_{1}$ is a bridge from $\sigma_{1}$ to $\sigma_{2}$, $\rho_{2}$ is a bridge from $\sigma_{2}$ to $\sigma_{3}$. Then the formula
$\rho(x_{1},x_{2},z_{1},z_{2})=\exists y_{1}\exists y_{2}\;\rho_{1}(x_{1},x_{2},y_{1},y_{2})\wedge\rho_{2}(y_{1},y_{2},z_{1},z_{2})$
defines a bridge from $\sigma_{1}$ to $\sigma_{3}$. Moreover, $\widetilde{\rho}=\widetilde{\rho_{1}}\circ\widetilde{\rho_{2}}$.
Lemma 5.
Suppose $\sigma$ is a irreducible congruence on $\mathbf{A}\in\mathcal{V}_{n}$, $\delta$ is a bridge from $\sigma$ to $\sigma$ such that $\widetilde{\delta}$ is linked. Then $\sigma$ is a perfect linear congruence.
Proof.
Let $\delta^{-1}$ be the bridge defined by $\delta^{-1}(y_{1},y_{2},x_{1},x_{2})=\delta(x_{1},x_{2},y_{1},y_{2})$. Since $\widetilde{\delta}$ is linked, $\underbrace{\widetilde{\delta}\circ\widetilde{\delta^{-1}}\circ\dots\widetilde{\delta}\circ\widetilde{\delta^{-1}}}_{2N}=A^{2}$ for sufficiently large $N$. Using Lemma 28 we compose $2N$ bridges $\delta,\delta^{-1},\dots,\delta,\delta^{-1}$ and obtain a new bride $\delta^{\prime}$ such that $\widetilde{\delta}^{\prime}=A^{2}$. By Lemma 27 $\sigma$ is a perfect linear congruence. ∎
Lemma 29.
Suppose $B\leqslant\mathbf{Z}_{p}\times\dots\times\mathbf{Z}_{p}$. Then there does not exist $C<_{\mathcal{T}}B$ such that $T\in\{\mathcal{BA},\mathcal{C}\}$.
Proof.
It is sufficient to check for any term $\tau$ that $\tau^{\mathbf{Z}_{p}}(a,\dots,a,x)$ takes all the values if the last variable is not dummy. Hence, there cannot be an absorbing subuniverse in $\mathbf{Z}_{p}\times\dots\times\mathbf{Z}_{p}$. ∎
3 Proof of the CSP Dichotomy Conjecture
In this section we prove Theorems 43 and 44 that show the correctness of Zhuk’s algorithm for the CSP. We start with Subsection 3.1, where we give additional definitions such as irreducible, linked, and crucial instances of the CSP. Crucial instances are the instances that have no solutions but any weakening of the instance (like removing a constraint) gives an instance with a solution. We will show any constraint in a crucial and consistent enough instance has the parallelogram property. This allows us to define a congruence for every constraint and its variable and talk about connectedness of the variables by bridges. Also, we explain how we weaken the instance: usually we just replace a constraint by a weaker constraint but sometimes we also need to disconnect two constraints by adding an additional variable, which leads us to the notion of Expanded Coverings.
In the next subsection we give all the auxiliary statements necessary for the main proof. Mainly we explain how our new theory works for the CSP and it works especially well if the solution set of the instance is subdirect.
The core of the proof of both main theorems is Theorem 41, which states that all constraints in a crucial instance have the parallelogram property, and there exists a crucial expanded covering with a connected subinstance. Additionally it states that a restriction of the domains to strong subalgebras cannot destroy all the solutions. As in the original proof, Theorem 41 is proved by induction on the size of the domain but this time we connect the variables using dividing congruences coming from the reductions, which significantly simplifies the whole argument.
3.1 Additional definitions
CSP Instances. An instance $\mathcal{I}$ of $\operatorname{CSP}(\Gamma)$ is a list (or conjunction) of constraints of the form $R(x_{1},\dots,x_{m})$, where $R\in\Gamma$. We write $C\in\mathcal{I}$ meaning that $C$ is a constraint of $\mathcal{I}$. For an instance $\mathcal{I}$ and a constraint $C$ by $\operatorname{Var}(\mathcal{I})$ and $\operatorname{Var}(C)$ we denote the set of variables appearing in $\mathcal{I}$ and $C$, respectively. Every variable $x$ appearing in an instance has its domain, which we denote by $D_{x}$. Every domain can be viewed as an algebra $\mathbf{D}_{x}=(D_{x};w^{\mathbf{D}_{x}})\in\mathcal{V}_{m}$. A subset of constraints of an instance $\mathcal{I}$ is called a subinstance of $\mathcal{I}$. Then for every constraint $R(x_{1},\dots,x_{h})$ the relation $R$ is a subuniverse of $\mathbf{D}_{x_{1}}\times\dots\times\mathbf{D}_{x_{h}}$. We say that a solution set of an instance $\mathcal{I}$ is subdirect if for every $x$ and every $a\in D_{x}$ the instance has a solution with $x=a$.
Reductions. A reduction $D^{(\top)}$ for a CSP instance $\mathcal{I}$ is mapping that assign a subuniverse $D_{x}^{(\top)}\leqslant\mathbf{D}_{x}$ to every variable $x$ of $\mathcal{I}$. $D$ can be viewed as a trivial reduction. For two reductions $D^{(\bot)}$ and $D^{(\top)}$ we write $D^{(\bot)}\lll D^{(\top)}$ and $D^{(\bot)}\leqslant_{T}D^{(\top)}$ whenever $D_{i}^{(\bot)}\lll D_{i}^{(\top)}$ for every $i\in I$ and $D_{i}^{(\bot)}\leqslant_{T}D_{i}^{(\top)}$ for every $i\in I$, respectively. For an instance $\mathcal{I}$ and a reduction $D^{(\top)}$ by $\mathcal{I}^{(\top)}$ we denote the instance whose variables $x$ are restricted to $D_{x}^{(\top)}$. A reduction $D^{(\top)}$ is called nonempty if $D_{x}^{(\top)}\neq\varnothing$ for every $x$.
Induced congruences. For a relation $R$ of arity $n$ and $i\in[n]$ by $\operatorname{Con}(R,i)$ we denote the binary relation $\sigma(y,y^{\prime})$ defined by
$\exists x_{1}\dots\exists x_{i-1}\exists x_{i+1}\dots\exists x_{n}\;R(x_{1},\ldots,x_{i-1},y,x_{i+1},\ldots,x_{n})\wedge R(x_{1},\ldots,x_{i-1},y^{\prime},x_{i+1},\ldots,x_{n}).$
For a constraint $C=R(x_{1},\ldots,x_{n})$ by $\operatorname{Con}(C,x_{i})$ we denote $\operatorname{Con}(R,i)$. For an instance $\mathcal{I}$ by $\operatorname{Con}(\mathcal{I},x)$ we denote the set $\{\operatorname{Con}(C,x)\mid C\in\mathcal{I}\}$. By $\operatorname{Con}(\mathcal{I})$ we denote $\bigcup\limits_{x\in\operatorname{Var}(\mathcal{I})}\operatorname{Con}(\mathcal{I},x)$. Notice that the $i$-th variable of a relation $R$ is rectangular if and only if $R$ is stable under $\operatorname{Con}(R,i)$. Moreover, if the $i$-th variable of a subdirect relation $R$ is rectangular then $\operatorname{Con}(R,i)$ is a congruence;
Linear-type and PC-type. We say that a relation $R$ is of the PC/Linear type if $R$ is rectangular and each congruence $\operatorname{Con}(R,i)$ is a PC/Linear congruence. We say that an instance has the PC/Linear type if all of its constraints are of the PC/Linear type.
A path and a tree-covering. We say that $z_{1}-C_{1}-z_{2}-\dots-C_{l-1}-z_{l}$ is a path in a CSP instance $\mathcal{I}$ if $z_{i},z_{i+1}\in\operatorname{Var}(C_{i})$. We say that a path $z_{1}-C_{1}-z_{2}-\dots-C_{l-1}-z_{l}$ connects $b$ and $c$ if there exists $a_{i}\in D_{z_{i}}$ for every $i$ such that $a_{1}=b$, $a_{l}=c$, and the projection of $C_{i}$ onto $z_{i},z_{i+1}$ contains the tuple $(a_{i},a_{i+1})$. We say that an instance is a tree-instance if there is no a path $z_{1}-C_{1}-z_{2}-\dots-z_{l-1}-C_{l-1}-z_{l}$ such that $l\geqslant 3$, $z_{1}=z_{l}$, and all the constraints $C_{1},\ldots,C_{l-1}$ are different.
Consistency conditions. A CSP instance $\mathcal{I}$ is called 1-consistent if $\operatorname{pr}_{z}(C)=D_{z}$ for any constraint $C$ of $\mathcal{I}$ and any variable $z$ of $C$. A reduction $D^{(\top)}$ is called 1-consistent for an instance $\mathcal{I}$ if the instance $\mathcal{I}^{(\top)}$ is 1-consistent. An instance $\mathcal{I}$ is called cycle-consistent if it is 1-consistent and for every variable $z$ and $a\in D_{z}$ any path starting and ending with $z$ in $\mathcal{I}$ connects $a$ and $a$. Other types of local consistency and its connection with the complexity of the CSP are considered in [25, 10].
Linkedness and irreducibility. An instance $\mathcal{I}$ is called linked if for every variable $z\in\operatorname{Var}(\mathcal{I})$ and every $a,b\in D_{z}$ there exists a path starting and ending with $z$ in $\mathcal{I}$ that connects $a$ and $b$. We say that an instance $\mathcal{I}$ is fragmented if $\operatorname{Var}(\mathcal{I})$ can be divided into 2 disjoint nonempty sets $\mathbf{X_{1}}$ and $\mathbf{X_{2}}$ such that $\operatorname{Var}(C)\subseteq\mathbf{X}_{1}$ or $\operatorname{Var}(C)\subseteq\mathbf{X}_{1}$ for any $C\in\mathcal{I}$. An instance $\mathcal{I}$ is called irreducible if there is no instance $\mathcal{I}^{\prime}$ satisfying the following conditions:
  1. 1.
    $\operatorname{Var}(\mathcal{I}^{\prime})\subseteq\operatorname{Var}(\mathcal{I})$,
  2. 2.
    each constraint of $\mathcal{I}^{\prime}$ is a projection of a constraint of $\mathcal{I}$ on some variables,
  3. 3.
    $\mathcal{I}^{\prime}$ is not fragmented,
  4. 4.
    $\mathcal{I}^{\prime}$ is not linked,
  5. 5.
    the solution set of $\mathcal{I}^{\prime}$ is not subdirect.
Weakening of an instance. We say that a constraint $R_{1}(y_{1},\ldots,y_{t})$ is weaker or equivalent to a constraint $R_{2}(z_{1},\ldots,z_{s})$ if $\{y_{1},\ldots,y_{t}\}\subseteq\{z_{1},\ldots,z_{s}\}$ and $R_{2}(z_{1},\ldots,z_{s})$ implies $R_{1}(y_{1},\ldots,y_{t})$. We say that $C_{1}$ is weaker than $C_{2}$ if $C_{1}$ is weaker or equivalent to $C_{2}$ but $C_{1}$ does not imply $C_{2}$. The weakening of a constraint $C$ in an instance $\mathcal{I}$ is the replacement of $C$ by all weaker constraints. An instance $\mathcal{I}^{\prime}$ is called a weakening of an instance $\mathcal{I}$ if $\operatorname{Var}(I^{\prime})\subseteq\operatorname{Var}(I)$ are every constraint of $\mathcal{I}^{\prime}$ is weaker or equivalent to a constraint of $\mathcal{I}$.
Crucial instance. We say that a variable $y_{i}$ of the constraint $R(y_{1},\ldots,y_{t})$ is dummy if ${\mathbb{R}}$ does not depend on its $i$-th variable. Let $D_{i}^{\prime}\subseteq D_{i}$ for every $i$. Suppose $D^{(\top)}$ is a reduction for an instance $\mathcal{I}$. A constraint $C$ of $\mathcal{I}$ is called crucial in $D^{(\top)}$ if it has no dummy variables, $\mathcal{I}^{(\top)}$ has no solutions but the weakening of $C\in\Theta$ gives an instance $\mathcal{I}^{\prime}$ with a solution in $D^{(\top)}$. An instance $\mathcal{I}$ is called crucial in $D^{(\top)}$ if it has at least one constraint and all its constraints are crucial in $D^{(\top)}$.
Remark 2.
Suppose $\mathcal{I}^{(\top)}$ has no solutions. Then we can iteratively replace every constraint by all weaker constraints having no dummy variables until it is crucial in $D^{(\top)}$. Notice that $R\leqslant\mathbf{D}_{x_{1}}\times\dots\times\mathbf{D}_{x_{n}}$ for any weaker constraint $R(x_{1},\dots,x_{n})$ we introduce.
Relations defined by instances. For an instance $\mathcal{I}$ and $x_{1},\dots,x_{n}\in\operatorname{Var}(\mathcal{I})$ by $\mathcal{I}(x_{1},\dots,x_{n})$ we denote the set of all tuples $(a_{1},\dots,a_{n})$ such that $\mathcal{I}$ has a solution with $x_{i}=a_{i}$ for every $i$. Thus, $\mathcal{I}(x_{1},\dots,x_{n})$ defines an $n$-ary relation. Note that the obtained relation is a subuniverse of $\mathbf{D}_{x_{1}}\times\dots\times\mathbf{D}_{x_{n}}$ as it is defined by a primitive positive formula over the relations in $\mathcal{I}$ (see [19, 7, 8]).
Expanded coverings. For an instance $\mathcal{I}$ by $\operatorname{ExpCov}(\mathcal{I})$ (Expanded Coverings) we denote the set of all instances $\mathcal{I}^{\prime}$ such that there exists a mapping $S:\operatorname{Var}(\Omega^{\prime})\to\operatorname{Var}(\Omega)$ satisfying the following conditions:
  1. 1.
    if $x\in\operatorname{Var}(\mathcal{I})\cap\operatorname{Var}(\mathcal{I}^{\prime})$ then $S(x)=x$;
  2. 2.
    $D_{x}=D_{S(x)}$ for every $x\in\operatorname{Var}(\mathcal{I}^{\prime})$;
  3. 3.
    for every constraint $R(x_{1},\ldots,x_{n})$ of $\mathcal{I}^{\prime}$ either the variables $S(x_{1}),\ldots,S(x_{n})$ are different and the constraint $R(S(x_{1}),\ldots,S(x_{n}))$ is weaker or equivalent to some constraint of $\Omega$, or $S(x_{1})=\dots=S(x_{n})$ and $\{(a,a,\ldots,a)\mid a\in D_{x_{1}}\}\subseteq R$;
An expanded covering $\mathcal{I}^{\prime}$ of $\mathcal{I}$ is called a covering if for every constraint $R(x_{1},\ldots,x_{n})$ of $\mathcal{I}^{\prime}$ the constraint $R(S(x_{1}),\ldots,S(x_{n}))$ is in $\mathcal{I}$. An instance is called a tree-covering if it is a covering and also a tree-instance. For a variable $x$ we say that $S(x)$ is the parent of x and $x$ is a child of $S(x)$. The same child/parent terminology will also be applied to constraints.
The following easy facts can be derived from the definition.
  1. (p1)
    If we replace every variable $x$ by $S(x)$ in an expanded covering of $\mathcal{I}$ (and remove all the constraints $R(x,x,\dots,x)$) we get a weakening of $\mathcal{I}$;
  2. (p2)
    A weakening is an expanded covering such that $S(x)=x$ for every $x$;
  3. (p3)
    any solution of an instance can be naturally expanded to a solution of its expanded covering;
  4. (p4)
    if an instance is 1-consistent and its expanded covering is a tree-covering, then the solution set of the covering is subdirect;
  5. (p5)
    the union (union of all constraints) of two expanded coverings is also a expanded covering;
  6. (p6)
    an expanded covering of an expanded covering is an expanded covering.
  7. (p7)
    an expanded covering of a cycle-consistent irreducible instance is cycle-consistent and irreducible (see Lemma 30).
  8. (p8)
    any reduction of an instance can be naturally extended to its expanded covering; moreover, if the reduction was 1-consistent for the instance, it is 1-consistent for the covering.
Connected instances. A bridge $\delta\subseteq D^{4}$ is called reflexive if $(a,a,a,a)\in\delta$ for every $a\in D$. We say that two congruences $\sigma_{1}$ and $\sigma_{2}$ on $\mathbf{D}_{x}$ are adjacent if there exists a reflexive bridge from $\sigma_{1}$ to $\sigma_{2}$. Since we can always put $\delta(x_{1},x_{2},x_{3},x_{4})=\sigma(x_{1},x_{3})\wedge\sigma(x_{2},x_{4})$, any proper congruence $\sigma$ is adjacent with itself. We say that two rectangular constraints $C_{1}$ and $C_{2}$ are adjacent in a common variable $x$ if $\operatorname{Con}(C_{1},x)$ and $\operatorname{Con}(C_{2},x)$ are adjacent. An instance $\mathcal{I}$ is called connected if all its constraints are rectangular, all the congruences of $\operatorname{Con}(\mathcal{I})$ are irreducible, and the graph, whose vertices are constraints and edges are adjacent constraints, is connected.
3.2 Auxiliary statements
Lemma 30 ([33], Lemma 6.1).
Suppose $\mathcal{I}$ is a cycle-consistent irreducible CSP instance and $\mathcal{I}^{\prime}\in\operatorname{ExpCov}(\mathcal{I})$. Then $\mathcal{I}^{\prime}$ is cycle-consistent and irreducible.
Lemma 31.
Suppose
  1. 1.
    $D^{(1)}$ is a 1-consistent reduction for an instance $\mathcal{I}$,
  2. 2.
    $D_{x}^{(1)}$ is S-free for every $x\in\operatorname{Var}(\mathcal{I})$,
  3. 3.
    $T\in\{\mathcal{PC},\mathcal{L},\mathcal{D}\}$,
  4. 4.
    $D^{(1)}\lll D$,
  5. 5.
    $D_{x}^{(2)}\leqslant_{\mathcal{M}T}D_{x}^{(1)}$ is a minimal $\mathcal{M}T$ subuniverse for every $x\in\operatorname{Var}(\mathcal{I})$.
Then either there exists a constraint $C$ such that $C^{(2)}$ is empty, or $\mathcal{I}^{(2)}$ is 1-consistent.
Proof.
If $C^{(2)}$ is empty for some constraint $C$ then we are done. Otherwise, consider some constraint $R(x_{1},\dots,x_{n})$. By Lemma 14(fm) $\operatorname{pr}_{i}(R^{(2)})\leqslant_{\mathcal{MD}}^{D_{x_{i}}}D_{x_{i}}^{(1)}$ for every $i\in[n]$. Since $D_{x_{i}}^{(2)}$ is a minimal subuniverse $B$ such that $B\leqslant_{\mathcal{MD}}^{D_{x_{i}}}D_{x_{i}}^{(1)}$ we have $\operatorname{pr}_{i}(R^{(2)})=D_{x_{i}}^{(2)}$. Hence $\mathcal{I}^{(2)}$ is 1-consistent. ∎
Lemma 32.
Suppose $R(x_{1},\dots,x_{n})$ is a rectangular constraint of a 1-consistent instance $\mathcal{I}$, $R(x_{1},\dots,x_{n})$ is crucial in $D^{(\top)}$. Then $\operatorname{Con}(R,i)$ is an irreducible congruence for every $i\in[n]$.
Proof.
To simplify notations assume that $i=1$. Assume the converse, then $\operatorname{Con}(R,1)=\omega_{1}\cap\omega_{2}$ for some $\operatorname{Con}(R,1)\lneq\omega_{1},\omega_{2}\leqslant\mathbf{D}_{x_{1}}\times\mathbf{D}_{x_{1}}$. Define the relation $R_{j}$ for $j\in\{1,2\}$ by
$R_{j}(x_{1},x_{2},\dots,x_{n})=\exists y(R(y,x_{2},\dots,x_{n})\wedge\omega_{1}(y,x_{1})).$
Since $\omega_{i}\supsetneq\operatorname{Con}(R,1)$ we have $R_{j}\supsetneq R$ for each $j\in\{1,2\}$. Since $\omega_{1}\cap\omega_{2}=\operatorname{Con}(R,1)$ we have $R=R_{1}\cap R_{2}$. Thus $R(x_{1},\dots,x_{n})$ could be replaced by two weaker constraints $R_{1}(x_{1},\dots,x_{n})$ and $R_{2}(x_{1},\dots,x_{n})$ and still be without a solution in $D^{(1)}$. This contradicts the cruciality. ∎
Lemma 33.
Suppose $R\leqslant_{sd}\mathbf{A}_{1}\times\dots\times\mathbf{A}_{n}$, the first and the last variables of $R$ are rectangular, and there exist $(b_{1},a_{2},\ldots,a_{n}),(a_{1},\ldots,a_{n-1},b_{n})\in R$ such that $(a_{1},a_{2},\ldots,a_{n})\notin R$. Then there exists a bridge $\delta$ from $\operatorname{Con}(R,1)$ to $\operatorname{Con}(R,n)$ such that $\widetilde{\delta}=\operatorname{pr}_{1,n}(R)$.
Proof.
The required bridge can be defined by
$\delta(x_{1},x_{2},y_{1},y_{2})=\exists z_{2}\dots\exists z_{n-1}\;R(x_{1},z_{2},\ldots,z_{n-1},y_{1})\wedge R(x_{2},z_{2},\ldots,z_{n-1},y_{2}).$
In fact, since the first and the last variables of $R$ are rectangular, we have $(x_{1},x_{2})\in\operatorname{Con}(R,1)$ if and only if $(y_{1},y_{2})\in\operatorname{Con}(R,n)$. It remains to notice that $(b_{1},a_{1},a_{n},b_{n})\in\delta$, $(b_{1},a_{1})\notin\operatorname{Con}(R,1)$, and $\widetilde{\delta}=\operatorname{pr}_{1,n}(R)$. ∎
Lemma 34.
Suppose $\mathcal{I}$ is a cycle-consistent connected instance. Then
  1. (a)
    any two constraints with a common variable are adjacent;
  2. (b)
    for any constraints $C_{1},C_{2}\in\mathcal{I}$, variables $x_{1}\in\operatorname{Var}(C_{1})$, $x_{2}\in\operatorname{Var}(C_{2})$, and any path from $x_{1}$ to $x_{2}$, there exists a bridge $\delta$ from $\operatorname{Con}(C_{1},x_{1})$ to $\operatorname{Con}(C_{2},x_{2})$ such that $\widetilde{\delta}$ contains all pairs connected by this path;
  3. (p)
    if $\mathcal{I}$ is linked then $\operatorname{Con}(C,x)$ is a perfect linear congruence for every constraint $C\in\mathcal{I}$ and $x\in\operatorname{Var}(C)$.
Proof.
Let us prove (a) for two constraints $C_{1}$ and $C_{2}$ with a common variable $x$. Since $\mathcal{I}$ is connected, there exists a path $z_{1}-C_{1}^{\prime}-z_{2}-C_{2}^{\prime}-\dots-z_{\ell}-C_{\ell}^{\prime}-z_{\ell+1}$ such that $z_{1}=z_{\ell+1}=x$, $C_{1}^{\prime}=C_{1}$, $C_{\ell}^{\prime}=C_{2}$, $C_{j}^{\prime}$ and $C_{j+1}$ are adjacent in a common variable $z_{j+1}$ for each $j\in[\ell-1]$. Let $\omega_{j}$ be a reflexive bridge from $\operatorname{Con}(C_{j}^{\prime},z_{j+1})$ to $\operatorname{Con}(C_{j+1}^{\prime},z_{j+1})$. By Lemma 33 for every $i\in[\ell]$ there exists a bridge $\delta_{i}$ from $\operatorname{Con}(C_{i}^{\prime},z_{i})$ to $\operatorname{Con}(C_{i}^{\prime},z_{i+1})$ such that $\widetilde{\delta_{i}}=\operatorname{pr}_{z_{i},z_{i+1}}(C_{i}^{\prime})$. Since $\mathcal{I}$ is connected, all the congruences $\operatorname{Con}(C_{i}^{\prime},z_{i}),\operatorname{Con}(C_{i}^{\prime},z_{i+1})$ are irreducible. Composing bridges $\delta_{1},\omega_{1},\delta_{2},\omega_{2},\dots,\delta_{\ell-1},\omega_{\ell-1},\delta_{\ell}$ using Lemma 28 we get the required bridge from $\operatorname{Con}(C_{1}^{\prime},z_{1})$ to $\operatorname{Con}(C_{\ell}^{\prime},z_{\ell+1})$. Since $\mathcal{I}$ is cycle-consistent, the bridge is reflexive, and therefore $C_{1}$ and $C_{2}$ are adjacent.
To prove (b) we repeat the whole argument of (a) for the path in $\mathcal{I}$. Since we already proved (a), $C_{j}^{\prime}$ and $C_{j+1}^{\prime}$ are adjacent in a common variable $z_{j+1}$ for any path. As a result we obtain the required bridge $\operatorname{Con}(C_{1}^{\prime},z_{1})$ to $\operatorname{Con}(C_{\ell}^{\prime},z_{\ell+1})$.
Let us prove (p). Since $\mathcal{I}$ is connected, for any $a,b\in D_{x}$ there exists a path from $x$ to $x$ connecting $a$ and $b$. Let us build a bridge $\delta_{a,b}$ using (b) for this path. Since $\mathcal{I}$ is cycle-consistent, $\delta_{a,b}$ is reflexive. Composing all the bridges $\delta_{a,b}$ for $a,b\in D_{x}$ we get a bridge $\delta$ from $\operatorname{Con}(C,x)$ to $\operatorname{Con}(C,x)$ such that $\widetilde{\delta}=D_{x}^{2}$. By Lemma 27 $\operatorname{Con}(C,x)$ is a perfect linear congruence. ∎
Lemma 35 ([34], Lemma 5.6).
Suppose $D^{(\top)}$ is a reduction for an instance $\mathcal{I}$, $D^{(\bot)}$ is an inclusion maximal 1-consistent reduction for $\mathcal{I}$ such that $D^{(\bot)}\leqslant D^{(\top)}$. Then for every variable $y\in\operatorname{Var}(\mathcal{I})$ there exists a tree-covering $\Upsilon_{y}$ of $\mathcal{I}$ such that $\Upsilon_{y}^{(\top)}(y)$ defines $D_{y}^{(\bot)}$.
Corollary 36.
Suppose $D^{(\top)}$ is a reduction of a 1-consistent instance $\mathcal{I}$, $D^{(\top)}\lll D$, $D^{(\bot)}$ is an inclusion-maximal nonempty 1-consistent reduction of $\mathcal{I}$ such that $D^{(\bot)}\leqslant D^{(\top)}$. Then $D^{(\bot)}\lll^{D}D^{(\top)}\lll D$.
Proof.
By Lemma 35 for every variable $y\in\operatorname{Var}(\mathcal{I})$ there exists a tree-covering $\Upsilon_{y}$ of $\mathcal{I}$ such that $\Upsilon_{y}^{(\top)}(y)$ defines $D_{y}^{(\bot)}$. Since $\mathcal{I}$ is 1-consistent, the solution set of $\Upsilon_{y}$ can be viewed as a subdirect relation. By Corollary 18(r1) we obtain $\Upsilon_{y}^{(\top)}(y)=D_{y}^{(\bot)}\lll^{D_{y}}D_{y}^{(\top)}$
Lemma 37.
Suppose $D^{(1)}$ is a 1-consistent reduction of a cycle-consistent instance $\mathcal{I}$, $D^{(1)}\lll D$, $B<_{T}^{D_{x}}D_{x}^{(1)}$ for some variable $x$, and $T\in\{\mathcal{BA},\mathcal{C},\mathcal{PC}\}$. Then there exists a nonempty 1-consistent reduction $D^{(2)}\lll D^{(1)}$ such that $D_{x}^{(2)}\leqslant B$. Moreover,
  1. 1.
    if $T\in\{\mathcal{BA},\mathcal{C}\}$ then $D^{(2)}\leqslant_{T}D^{(1)}$;
  2. 2.
    if $T=\mathcal{PC}$ and $D_{y}^{(1)}$ is S-free for every $y\in\mathcal{I}$ then $D^{(2)}\leqslant_{\mathcal{MPC}}D^{(1)}$.
Proof.
Define the reduction $D^{(\top)}$ by $D_{x}^{(\top)}=B$ and $D_{y}^{(\top)}=D_{y}^{(1)}$ for every $y\neq x$. Let $D^{(2)}$ be an inclusion maximal 1-consistent reduction for $\mathcal{I}$ such that $D^{(2)}\leqslant D^{(\top)}$. By Lemma 35 for every variable $y\in\operatorname{Var}(\mathcal{I})$ there exists a tree-covering $\Upsilon_{y}$ of $\mathcal{I}$ such that $\Upsilon_{y}^{(\top)}(y)$ defines $D_{y}^{(2)}$.
Assume that $D_{y}^{(2)}=\varnothing$ and $\Upsilon_{y}^{(\top)}$ has no solutions for some $y$. Since $\mathcal{I}$ is 1-consistent, the solution set of $\Upsilon_{y}$ can be viewed as a subdirect relation. By Corollary 22 there should two children of $x$ in $\Upsilon_{y}$ such that if we restrict them to $D_{x}^{(\top)}$ we kill all the solutions of $\Upsilon_{y}$. Since $\Upsilon_{y}$ is a tree-covering of $\mathcal{I}$ and $\mathcal{I}$ is cycle-consistent, this cannot happen.
Thus, $D^{(2)}$ is a nonempty reduction. Assume that $\mathcal{T}\in\{\mathcal{BA},\mathcal{C}\}$. Again considering the solution set $\Upsilon_{y}$ and applying Lemma 19 we derive that $D_{y}^{(2)}\leqslant_{\mathcal{T}}D_{y}^{(1)}$, and therefore $D^{(2)}\leqslant_{\mathcal{T}}D^{(1)}$. For $\mathcal{T}=\mathcal{PC}$ we do the same but apply Corollary 18(rm) instead and obtain $D^{(2)}\leqslant_{\mathcal{MPC}}^{D}D^{(1)}$. ∎
Lemma 38.
Suppose
  1. 1.
    $D^{(1)}$ is a 1-consistent reduction for a constraint $R(x_{1},\dots,x_{n})$,
  2. 2.
    $T\in\{\mathcal{L},\mathcal{PC},\mathcal{D}\}$,
  3. 3.
    $D^{(2)}\leqslant_{\mathcal{M}T}^{D}D^{(1)}\lll D$,
  4. 4.
    $R(x_{1},\dots,x_{n})$ is crucial (as the whole instance) in $D^{(2)}$.
Then $R$ has the parallelogram property and $\operatorname{Con}(R,i)$ is a congruence of type $T$ such that $\operatorname{Con}(R,i)^{*}\supseteq(D_{x_{i}}^{(1)})^{2}$ for every $i\in[n]$. Moreover, if $T=\mathcal{PC}$ then $n=2$.
Proof.
First, let us prove that $R$ has the parallelogram property. We need to check the parallelogram property for any splitting of the variables of $R$ into two disjoint sets. Without loss of generality we assume that this splitting is $\{x_{1},\dots,x_{k}\}$ and $\{x_{k+1},\dots,x_{n}\}$. Let us define a binary relation $R^{\prime}\leqslant_{sd}\mathbf{E}_{1}\times\mathbf{E}_{2}$ by $((a_{1},\dots,a_{k}),(a_{k+1},\dots,a_{n}))\in R^{\prime}\Leftrightarrow(a_{1},\dots,a_{n})\in R$, where $E_{1}=\operatorname{pr}_{1,\dots,k}(R)$, $E_{2}=\operatorname{pr}_{k+1,\dots,n}(R)$. Let us define $E_{1}^{(1)}$, $E_{2}^{(1)}$, $E_{1}^{(2)}$, $E_{2}^{(2)}$, naturally (we just reduce the corresponding coordinates to $D^{(1)}$ or $D^{(2)}$). By Corollary 18(r) and (m) we have $E_{i}^{(2)}\leqslant_{\mathcal{M}T}^{E_{i}}E_{i}^{(1)}\lll E_{i}$ for each $i\in\{1,2\}$. Put $S^{\prime}=R^{\prime}\circ R^{\prime-1}\circ R^{\prime}$ and $S=\{(a_{1},\dots,a_{n})\mid((a_{1},\dots,a_{k}),(a_{k+1},\dots,a_{n}))\in S^{\prime}\}$. Since $R^{\prime(2)}=\varnothing$, Lemma 24 implies that $S^{\prime(2)}=\varnothing$. Since $R(x_{1},\dots,x_{n})$ is crucial and $S\supseteq R$, we obtain $S=R$. Hence $R$ has the parallelogram property.
Put $E=\operatorname{pr}_{1}(R\cap(D_{x_{1}}\times D_{x_{2}}^{(2)}\times\dots\times D_{x_{n}}^{(2)}))$. If $E=\varnothing$ then the constraint could be weakened to $R_{0}(x_{2},\dots,x_{n})$, where $R_{0}(x_{2},\dots,x_{n})=\exists x_{1}R(x_{1},\dots,x_{n})$, which contradicts the cruciality. Hence $E\neq\varnothing$ and by Corollary 18(r1) $E\lll D_{x_{1}}$. Since $R(x_{1},\dots,x_{n})$ is crucial in $D^{(2)}$ we have $E\cap D_{x_{1}}^{(2)}=\varnothing$.
If $E\cap D_{x_{1}}^{(1)}=\varnothing$, choose $C$, $B$, and $\mathcal{T}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{PC},\mathcal{D}\}$ such that $D_{x_{1}}^{(1)}\lll^{D_{x_{1}}}C<_{\mathcal{T}}^{D_{x_{1}}}B\lll D_{x_{1}}$, $E\cap C=\varnothing$, and $E\cap B\neq\varnothing$. Since $R^{(1)}$ is not empty and $D^{(2)}\leqslant_{\mathcal{M}T}^{D}D^{(1)}$, Corollary 22 implies that $\mathcal{T}=T$. If $E\cap D_{x_{1}}^{(1)}\neq\varnothing$ put $C=D_{x_{1}}^{(2)}$ and $B=D_{x_{1}}^{(1)}$. Thus, in both cases we have $C<_{\mathcal{M}T}^{D_{x_{1}}}B\lll D_{x_{1}}$.
Notice that $E\circ\operatorname{Con}(R,1)=E$, hence $(E\circ\operatorname{Con}(R,1))\cap C=\varnothing$. Let $\sigma\supseteq\operatorname{Con}(R,1)$ be a maximal congruence such that $(E\circ\sigma)\cap C=\varnothing$. If $\sigma\supsetneq\operatorname{Con}(R,1)$ then we weaken the constraint $R(x_{1},\dots,x_{n})$ to $R_{0}(x_{1},\dots,x_{n})$, where $R_{0}(x_{1},\dots,x_{n})=\exists zR(z,x_{2},\dots,x_{n})\wedge\sigma(z,x_{1})$. The obtained constraint must have a solution in $D^{(2)}$, which means that $(E\circ\sigma)\cap D_{x_{1}}^{(2)}\neq\varnothing$ and contradicts $(E\circ\sigma)\cap C=\varnothing$. Thus, $\sigma=\operatorname{Con}(R,1)$. By Lemma 25 $\operatorname{Con}(R,1)=\omega_{1}\cap\dots\cap\omega_{s}$ for some congruences $\omega_{1},\dots,\omega_{s}$ of type $T$ such that $\omega_{i}^{*}\supseteq B^{2}$. By Lemma 32 $\operatorname{Con}(R,1)$ is irreducible, hence $\operatorname{Con}(R,1)$ is a congruence of type $T$ satisfying $\operatorname{Con}(R,1)^{*}\supseteq B^{2}\supseteq(D_{x_{1}}^{(1)})^{2}$.
It remains to show that $n=2$ for $T=\mathcal{PC}$. By Corollary 22 there exist $i,j\in[n]$, $B_{i}<_{PC(\sigma_{i})}^{D_{x_{i}}}D_{x_{i}}^{(1)}$, and $B_{j}<_{PC(\sigma_{j})}^{D_{x_{j}}}D_{x_{j}}^{(1)}$ such that $R$ has no tuples whose $i$-th element is from $B_{i}$ and $j$-th element is from $B_{j}$. If $n\geqslant 3$ the we can existentially quantify all the variables of $R$ but $i$-th and $j$-th and obtain a weaker constraint without a solution in $D^{(2)}$, which contradicts cruciality. ∎
Lemma 39.
Suppose
  1. 1.
    $\mathcal{I}$ is an instance having a subdirect solution set,
  2. 2.
    $D^{(1)}$ is a reduction for $\mathcal{I}$ such that $D_{x}^{(1)}\lll D_{x}$ for every $x$,
  3. 3.
    $C$ is a constraint in $\mathcal{I}$ of type $T\in\{\mathcal{PC},\mathcal{L}\}$,
  4. 4.
    $B<_{\mathcal{T}(\xi)}^{D_{z}}D_{z}^{(1)}$ for some variable $z$, where $\mathcal{T}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{PC},\mathcal{L}\}$,
  5. 5.
    if $T=\mathcal{PC}$ then $\mathcal{T}\in\{\mathcal{PC},\mathcal{L}\}$,
  6. 6.
    $\mathcal{I}^{(1)}$ has a solution,
  7. 7.
    $\mathcal{I}^{(1)}$ has no solutions with $z\in B$,
  8. 8.
    weakening of $C$ in $\mathcal{I}$ gives an instance with a solution in $D^{(1)}$ and $z\in B$.
Then $\mathcal{T}=T$ and for any variable $x$ of $C$ there exists a bridge $\delta$ from $\xi$ to $\operatorname{Con}(C,x)$ such that $\widetilde{\delta}$ contains $\mathcal{I}(z,x)$.
Proof.
By $D^{(2)}$ we denote the reduction that differs from $D^{(1)}$ only on the variable $z$ and $D_{z}^{(2)}=B$.
Choose some variable $x_{0}$ in $C$. Let $\omega=\operatorname{Con}(C,x_{0})$. By condition 3, $\omega$ is either a PC, or linear congruence. We take $\mathcal{I}$, replace the variable $x_{0}$ in $C$ by $x_{0}^{\prime\prime}$, all the other variables $x_{i}$ by $x_{i}^{\prime}$, and add a new constraint $\omega^{*}(x_{0}^{\prime},x_{0}^{\prime\prime})$. The obtained instance we denote by $\Theta$. Extend our reduction $D^{(1)}$ and $D^{(2)}$ to $\Theta$ by $D_{x_{0}^{\prime\prime}}^{(1)}=D_{x_{0}^{\prime\prime}}^{(2)}=D_{x_{0}^{\prime\prime}}=D_{x_{0}}$, $D_{x_{i}^{\prime}}^{(2)}=D_{x_{i}}^{(2)}$, and $D_{x_{i}^{\prime}}^{(1)}=D_{x_{i}}^{(1)}$. Notice that the solution set of $\Theta$ is still subdirect and $\Theta^{(2)}$ has a solution.
Let us consider a minimal reduction $D^{(\top)}$ for $\mathcal{I}$ such that $D^{(1)}_{x}\lll^{D_{x}}D^{(\top)}_{x}\lll D_{x}$ for every $x$ and $\Theta^{(2)}\wedge\mathcal{I}^{(\top)}\wedge\omega(x_{0},x_{0}^{\prime\prime})$ has a solution. If $D^{(\top)}\neq D^{(1)}$, choose a variable $y$ and $G<_{\mathcal{T}_{0}(\nu)}^{D_{y}}D_{y}^{(\top)}$ such that $D_{y}^{(1)}\lll^{D_{y}}G$. If $D^{(\top)}=D^{(1)}$ then put $G=B$, $y=z$, $\mathcal{T}_{0}=\mathcal{T}$, and $\nu=\xi$.
Define a new reduction $D^{(\bot)}$ by $D_{y}^{(\bot)}=G$ and $D_{x}^{(\bot)}=D_{x}^{(\top)}$ for every $x\neq y$. We extend the reduction $D^{(\top)}$ and $D^{(\bot)}$ so that the reductions on $x_{i}$ and $x_{i}^{\prime}$ coincide and $D_{x_{0}^{\prime\prime}}^{(\top)}=D_{x_{0}^{\prime\prime}}^{(\bot)}=D_{x_{0}}$. Since the instance $\Theta\wedge\mathcal{I}\wedge\omega(x_{0},x_{0}^{\prime\prime})$ has a subdirect solution set, by Corollary 22 the types $\mathcal{T}$ and $\mathcal{T}_{0}$ are the same. Moreover, if $\mathcal{T}\in\{\mathcal{L},\mathcal{PC}\}$, there exists a bridge $\delta^{\prime}$ from $\xi$ to $\nu$ such that $\widetilde{\delta}^{\prime}\supseteq\Theta(z^{\prime},x_{0}^{\prime\prime})\circ\mathcal{I}(x_{0},y)\supseteq\mathcal{I}(z,y)$ (if $z=y$ then it is just a trivial reflexive bridge). Let $F$ be set of possible values of $x_{0}^{\prime\prime}$ in the solutions of $\Theta^{(2)}\wedge\mathcal{I}^{(\top)}\wedge\omega(x_{0},x_{0}^{\prime\prime})$. In other words, $F=\Theta^{(2)}(x_{0}^{\prime\prime})\cap(\mathcal{I}^{(\top)}(x_{0})\circ\omega)$. Since the variable $x_{0}^{\prime\prime}$ only appears in the constraint $C$ and $\omega^{*}$ we have $F\circ\omega=F$. By Corollaries 18(r1) $F\lll D_{x_{1}}$. By Lemma 13 we find a single block $E$ of $\omega$ such that $\{E\}\lll^{D_{x_{0}}/\omega}F/\omega$. By Corollary 17(f) $E\lll F$.
As $\mathcal{I}^{(\top)}$ has a solution with $x_{0}\in E\subseteq F$, the instance $\Theta^{(\top)}$ has a solution with $x_{0}^{\prime},x_{0}^{\prime\prime}\in E$. As $\Theta^{(2)}$ has a solution with $x_{0}^{\prime\prime}\in E\subseteq F$, the instance $\Theta^{(\bot)}$ has a solution with $x_{0}^{\prime\prime}\in E$. As $\mathcal{I}^{(\bot)}$ has no solutions with $x_{0}\in F\supseteq E$, the instance $\Theta^{(\bot)}$ has no solutions with $x_{0}^{\prime},x_{0}^{\prime\prime}\in E$. Consider two cases:
Case 1. $T=\mathcal{L}$ and $\mathcal{T}\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$. Let $G_{1}$ be the set of all values of $x_{0}^{\prime}$ in solutions of $\Theta^{(\top)}$ with $x_{0}^{\prime\prime}\in E$, and $G_{2}$ be the set of all values of $x_{0}^{\prime}$ in solutions of $\Theta^{(\bot)}$ with $x_{0}^{\prime\prime}\in E$. By Lemma 19 $G_{2}\leqslant_{\mathcal{T}}G_{1}$. By Corollary 15(s) $G_{2}/\omega\leqslant_{\mathcal{T}}G_{1}/\omega$. Since $E\subseteq G_{1}$ and $E\not\subseteq G_{2}$, we have $G_{2}/\omega<_{\mathcal{T}}G_{1}/\omega$. By the construction of $\Theta$, $G_{1}$ and $G_{2}$ are from the same block of $\omega^{*}$. Hence we obtained a BA or central subuniverse in a block of $\omega^{*}$, which contradicts the properties of a linear congruence.
Case 2. $\mathcal{T}\in\{\mathcal{PC},\mathcal{L}\}$. Choose $E^{\prime}$ and $E^{\prime\prime}$ such that $E\lll^{D_{x_{0}}}E^{\prime}<_{T_{0}(\zeta)}^{D_{x_{0}}}E^{\prime\prime}\lll D_{x_{0}}$ and $\Theta^{(\bot)}$ has a solution with $x_{0}^{\prime\prime}\in E,x_{0}^{\prime}\in E^{\prime\prime}$ but has no solutions with $x_{0}^{\prime\prime}\in E,x_{0}^{\prime}\in E^{\prime}$. Notice that we can choose $E^{\prime}$, $E^{\prime\prime}$, and $\zeta$ stable under $\omega$ as $E^{\prime}$ and $E^{\prime\prime}$ may come from $\{E\}\lll D_{x_{0}}/\omega$ and Corollary 17(f). By Corollary 22 we derive that $T_{0}=\mathcal{T}_{0}$ and there exists a bridge $\delta^{\prime\prime}$ from $\nu$ to $\zeta$ such that $\widetilde{\delta}^{\prime\prime}=\Theta(y^{\prime\prime},x_{0}^{\prime\prime})\supseteq\mathcal{I}(y,x_{0})$. Notice that $\zeta$ must be equal to $\omega$ as otherwise $\omega^{*}\subseteq\zeta$ and any solution of $\Theta^{(\bot)}$ with $x_{0}\in E$ and $x_{0}^{\prime}\in E^{\prime\prime}$ also satisfies $x_{0}^{\prime}\in E^{\prime}$. It remains to compose bridges $\delta^{\prime}$ and $\delta^{\prime\prime}$ to obtain a bridge $\delta$ from $\xi$ to $\omega$ such that $\widetilde{\delta}\supseteq\mathcal{I}(z,y)\circ\mathcal{I}(y,x_{0})\supseteq\mathcal{I}(z,x_{0})$. ∎
Corollary 40.
Suppose $\mathcal{I}$ is an instance having subdirect solution set, $D^{(1)}$ and $D^{(2)}$ are reductions for $\mathcal{I}$, $\mathcal{I}^{(1)}$ has a solution, $D^{(2)}\leqslant_{T}^{D}D^{(1)}\lll D$, where $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$, $C$ is a constraint of $\mathcal{I}$ of type $\mathcal{L}$. Then $C$ is not crucial in $D^{(2)}$.
Proof.
We take a minimal reduction $D^{(\top)}$ such that $D_{x}^{(\top)}\in\{D_{x}^{(1)},D_{x}^{(2)}\}$ for every $x$ and $\mathcal{I}^{(\top)}$ has a solution. Take some variable $z$ such that $D_{z}^{(\top)}=D_{z}^{(1)}$, take $B=D_{z}^{(2)}$, and apply Lemma 39 for the reduction $D^{(\top)}$. ∎
3.3 Main Statements
Theorem 41.
Suppose
  • $D^{(1)}$ is a 1-consistent reduction of an irreducible, cycle-consistent instance $\mathcal{I}$;
  • $D^{(1)}\lll D$.
If $\mathcal{I}$ is crucial in $D^{(1)}$ then (1a) and ((1b) or (1c)).
  1. (1a)
    every constraint of $\mathcal{I}$ has the parallelogram property;
  2. (1b)
    $\mathcal{I}$ is a connected linear-type instance having a subdirect solution set;
  3. (1c)
    there exists a expanded covering $\mathcal{J}$ of $\mathcal{I}$ with a linked connected subinstance $\Upsilon$ such that the solution set of $\Upsilon$ is not subdirect and $\mathcal{J}$ is crucial in $D^{(1)}$.
If $D^{(2)}\leqslant_{\mathcal{T}}D^{(1)}$ is a 1-consistent reduction of $\mathcal{I}$, where $\mathcal{T}\in\{\mathcal{BA},\mathcal{C}\}$, and $\mathcal{I}^{(1)}$ has a solution, then
  1. (2)
    $\mathcal{I}^{(2)}$ has a solution.
Proof.
We prove the claim by induction on the size of $D^{(1)}$.
Let us prove (2) first. Assume that $\mathcal{I}^{(2)}$ has no solutions. Weaken $\mathcal{I}^{(2)}$ to make it crucial in $D^{(2)}$ and denote the obtained instance by $\mathcal{I}^{\prime}$. By the inductive assumption for $\mathcal{I}^{\prime}$ and $D^{(2)}$ the instance $\mathcal{I}^{\prime}$ satisfies (1a) and also (1b) or (1c). Assume that $\mathcal{I}^{\prime}$ satisfies (1c), then there exists an expanded covering $\mathcal{J}$ of $\mathcal{I}^{\prime}$ with a linked connected subinstance $\Upsilon$ such that $\mathcal{J}$ is crucial in $D^{(2)}$ Let $x$ be a variable of a constraint $C\in\Upsilon$. By Lemma 34(p) $\operatorname{Con}(C,x)$ is a perfect linear congruence. Choose $\zeta\subseteq\mathbf{D}_{x}\times\mathbf{D}_{x}\times\mathbf{Z}_{p}$ such that $(y_{1},y_{2},0)\in\zeta\Leftrightarrow(y_{1},y_{2})\in\operatorname{Con}(C,x)$ and $\operatorname{pr}_{1,2}(\zeta)={{\operatorname{Con}(C,x)}^{*}}$. Let us replace the variable $x$ of $C$ in $\mathcal{J}$ by $x^{\prime}$ and add the constraint $\zeta(x,x^{\prime},z)$. The obtained instance we denote by $\Theta$. We extend the reductions $D^{(1)}$ and $D^{(2)}$ to $x^{\prime}$ by $D_{x^{\prime}}^{(1)}=D_{x^{\prime}}^{(2)}=D_{x^{\prime}}$. Let $E_{1}$ and $E_{2}$ be the set of all $z$ such that $\Theta$ has a solution in $D^{(1)}$ and in $D^{(2)}$, respectively. By Lemma 19 $E_{2}\leqslant_{\mathcal{T}}E_{1}$. Since $\mathcal{J}$ is crucial in $D^{(2)}$, $\Theta^{(2)}$ must have some solution but not a solution with $z=0$. Since $\mathcal{I}^{(1)}$ has a solution, $\mathcal{J}^{(1)}$ also has a solution and $\Theta^{(1)}$ has a solution with $z=0$. Hence, $E_{1}$ contains at least two different elements, $0\in E_{1}$ and $0\notin E_{2}$. Since $\mathbf{Z}_{p}$ does not have proper subalgebras of size greater than 1, we have $E_{1}=\mathbf{Z}_{p}$, which contradicts the fact that $\mathbf{Z}_{p}$ has no BA or central subuniverses (Lemma 29).
Assume that $\mathcal{I}^{\prime}$ satisfies (1b). Applying Corollary 40 we derive a contradiction.
Let us prove (1a) and ((1b) or (1c)). Notice that $|D_{x}^{(1)}|>1$ for some $x$ as otherwise we would get a contradiction of 1-consistency of $\mathcal{I}$ and its cruciality in $D^{(1)}$. Consider two cases.
Case 1. There exists a nontrivial $\mathcal{BA}$ or central subuniverse on some $D^{(1)}_{x}$. By Lemma 37 there exists a 1-consistent reduction $D^{(2)}\leqslant_{T}D^{(1)}$ for $\mathcal{I}$ such that $T\in\{BA,C\}$. Let us show that $\mathcal{I}$ is crucial in $D^{(2)}$. Let $\mathcal{J}$ be obtained from $\mathcal{I}$ by a weakening of some constraint $C\in\mathcal{I}$. Then $\mathcal{J}^{(1)}$ has a solution. By the inductive assumption for $D^{(2)}$ we derive that $\mathcal{J}^{(2)}$ has a solution. Thus, any weakening of $\mathcal{I}$ has a solution in $D^{(2)}$ and $\mathcal{I}$ is crucial in $D^{(2)}$. Again applying the inductive assumption to $D^{(2)}$ we derive the required conditions, which completes this case.
Case 2. Otherwise. By Lemma 13, there exists $E<_{T(\sigma)}^{D_{z}}D^{(1)}_{z}$ for some $z$ and $T\in\{\mathcal{PC},\mathcal{L}\}$. Let us prove (1a) first. Choose some constraint $C$ in $\mathcal{I}$. Since $\mathcal{I}$ is crucial, a weakening of this constraint gives a solution in $D^{(1)}$. By $s(x)$ we denote the value of $x$ in this solution. For every variable $x$ choose the minimal $D^{(2)}_{x}\leqslant_{\mathcal{M}T}D^{(1)}$ containing $s(x)$. By Lemma 23 $D^{(2)}\lll^{D}D^{(1)}$ and by Lemma 31 we have two subcases.
Subcase 1. $\mathcal{I}^{(2)}$ is 1-consistent. Weaken the instance to make it crucial in $D^{(2)}$. Notice that the constraint $C$ must be in there because weakening of $C$ gives a solution in $D^{(2)}$. Then applying the inductive assumption to the obtained instance (crucial in $D^{(2)}$) we obtain the required property (1a) for $C$.
Subcase 2. There exists some constraint $C^{\prime}$ in $\mathcal{I}$ such that ${C^{\prime}}^{(2)}$ is empty. Since the weakening of $C$ gives a solution in $D^{(2)}$, $C^{\prime}$ must be $C$. By Lemma 38 $C$ has the parallelogram property, which is the property (1a).
Let us prove that (1b) or (1c) holds. Recall that we have $E<_{T(\sigma)}^{D_{z}}D^{(1)}_{z}$ for some $z$ and $T\in\{\mathcal{PC},\mathcal{L}\}$. Let =$\mathcal{B}=\{B\mid B<_{T(\sigma)}^{D_{z}}D^{(1)}_{z}\}$. For every $B\in\mathcal{B}$ we do the following. Let us consider the reduction $D^{(B,\top)}$ such that $D_{x}^{(B,\top)}=D_{x}^{(1)}$ if $x\neq z$ and $D_{z}^{(B,\top)}=B$. Let $D^{(B,\bot)}$ be the maximal 1-consistent (probably empty) reduction for $\mathcal{I}$ such that $D^{(B,\bot)}\leqslant D^{(B,\top)}$. By Lemma 35, for every variable $x$ and $B\in\mathcal{B}$ there exists a tree-covering $\Upsilon_{B,x}$ such that $\Upsilon_{B,x}^{(B,\top)}(x)$ defines $D_{x}^{(B,\bot)}$. By $\Upsilon_{x}=\bigwedge_{B\in\mathcal{B}}\Upsilon_{B,x}$ we define one universal tree-covering, that is, $\Upsilon_{x}^{(B,\top)}(x)$ defines $D_{x}^{(B,\bot)}$ for every $B\in\mathcal{B}$. We extend this definition to variables from an expanded covering of $\mathcal{I}$. Precisely, for a variable $x^{\prime}$ that is a child of $x$ by $\Upsilon_{x^{\prime}}$ we denote $\Upsilon_{x}$ whose variable $x$ is replaced by $x^{\prime}$. Let $\mathcal{B}_{0}$ be the set of all $B\in\mathcal{B}$ such that $D_{x}^{(B,\bot)}$ is not empty. Let us consider two cases:
Case 1. $\mathcal{B}_{0}$ is empty. Consider a tree-covering $\Upsilon$ such that $\Upsilon^{(B,\top)}$ has no solutions for every $B\in\mathcal{B}$. Since $\Upsilon$ is a tree-covering, its solution set is subdirect. Notice that $T$ cannot be the $\mathcal{PC}$ type, because Lemma 37 guarantees the existence of a nonempty reduction $D_{x}^{(B,\bot)}$ for every $B\in\mathcal{B}$. Hence $T=\mathcal{L}$.
Then weaken $\Upsilon$ while we can keep the property that $\Upsilon^{(B,\top)}$ has no solutions for every $B\in\mathcal{B}$. The obtained instance we denote by $\Upsilon^{\prime}$. Since $\mathcal{I}$ is crucial in $D^{(1)}$, $\Upsilon^{\prime}$ must contain every constraint relation that appeared in $\mathcal{I}$. Let us prove that $\mathcal{I}$ is connected. Take two constraints $C_{1}$ and $C_{2}$ of $\mathcal{I}$ having a common variable $x$. Applying Lemma 39 to $\Upsilon^{\prime}$ we obtain a bridge from $\operatorname{Con}(C_{1},x)$ to $\sigma$ and a bridge from $\sigma$ to $\operatorname{Con}(C_{2},x)$. Composing these bridges we obtain a bridge from $\operatorname{Con}(C_{1},x)$ to $\operatorname{Con}(C_{2},x)$. This bridge is reflexive because $\Upsilon$ a tree-covering and the path from a child of $x$ to a child of $z$ and back is just a path in the cycle-consistent instance $\mathcal{I}$. Additionally, we derived from Lemma 39 that all the congruences of $\operatorname{Con}(\mathcal{I})$ are of the linear type. Thus, we proved that any two constraints of $\mathcal{I}$ with a common variable are adjacent, which means that $\mathcal{I}$ is connected and satisfies (1b) if its solution set is subdirect or (1c) otherwise.
Case 2. $\mathcal{B}_{0}$ is not empty. For every expanded covering $\mathcal{J}$ of $\mathcal{I}$ by $\operatorname{Sol}(\mathcal{J})$ we denote the set of all $B\in\mathcal{B}_{0}$ such that $\mathcal{J}^{(B,\bot)}$ has a solution.
We want to find a set of instances $\Omega$ satisfying the following conditions:
  1. 1.
    Every instance in $\Omega$ is a weakening of $\mathcal{I}$.
  2. 2.
    $\bigcap_{\mathcal{J}\in\Omega}\operatorname{Sol}(\mathcal{J})=\varnothing$.
  3. 3.
    If we replace any instance in $\Omega$ by all weaker instances then 2 is not satisfied.
  4. 4.
    For every $\mathcal{J}\in\Omega$ there exists $B\in\mathcal{B}_{0}$ such that
    1. (a)
      $\mathcal{J}$ is crucial in $D^{(B,\bot)}$, and
    2. (b)
      $B\in\operatorname{Sol}(\mathcal{J}^{\prime})$ for every $J^{\prime}\in\Omega\setminus\{\mathcal{J}\}$.
We start with $\Omega=\{\mathcal{I}\}$. It already satisfies conditions 1 and 2. If 3 is not satisfied then we replace the corresponding instance by all weaker instances and get a new $\Omega$. We cannot weaken forever, that is why at some moment conditions 1-3 will be satisfied. Let us show that it also satisfies condition 4. Take some $\mathcal{J}\in\Omega$. For every constraint $C$ in $\mathcal{J}$ by $\mathcal{J}_{C}$ we denote the instance obtained from $\mathcal{J}$ by weakening $C$. By condition 3, $\Omega\cup\{\mathcal{J}_{C}\mid C\in\mathcal{J}\}\setminus\{\mathcal{J}\}$ cannot satisfy condition 2, which means that there exists $B\in\mathcal{B}_{0}$ such that $\mathcal{J}$ is crucial in $D^{(B,\bot)}$ and $B\in\operatorname{Sol}(\mathcal{J}^{\prime})$ for every $\mathcal{J}^{\prime}\in\Omega\setminus\{J\}$. Thus, we have $\Omega$ satisfying conditions 1-4.
For an expanded covering $\mathcal{J}$ of $\mathcal{I}$ by $\bot(\mathcal{J})$ we denote the instance $\mathcal{J}\wedge\bigwedge_{x\in\operatorname{Var}(\mathcal{J})}\Upsilon_{x}$, where we rename the variables so that the only common variable of $\Upsilon_{x}$ and $\mathcal{J}$ is $x$. Also, by $\Delta(\mathcal{J})$ we denote the instance that is obtained from $\mathcal{J}$ by adding the constraints $\sigma(z^{\prime},z^{\prime\prime})$ for every pair of variables whose parent is $z$.
For any weakening $\mathcal{J}$ of $\mathcal{I}$ and any $B\in\mathcal{B}_{0}$ such that $\mathcal{J}$ is crucial in $D^{(B,\bot)}$ we can apply the inductive assumption, which proves that either $\mathcal{J}$ satisfies (1b) or $\mathcal{J}$ satisfies (1c). Let us consider two subcases.
Subcase 1. Some instance $\mathcal{I}^{\prime}\in\Omega$ does not satisfy (1b). Let $\Omega^{\prime}$ be the set of all instances that are weaker than $\mathcal{I}^{\prime}$ joined with the instances from $\Omega\setminus\{\mathcal{I}^{\prime}\}$. Put $\mathcal{B}_{1}=\bigcap_{\mathcal{J}\in\Omega^{\prime}}\operatorname{Sol}(\mathcal{J})$. Notice that condition 3 implies that $\mathcal{B}_{1}$ is not empty. It follows from the definition that $\mathcal{I}^{\prime}$ is crucial in $D^{(B,\bot)}$ for every $B\in\mathcal{B}_{1}$. We want to build a sequence $\mathcal{J}_{1},\dots,\mathcal{J}_{s}$ of expanded coverings of $\mathcal{I}^{\prime}$ such that $\mathcal{B}_{1}\cap\bigcap_{i\in[s]}\operatorname{Sol}(\mathcal{J}_{i})=\varnothing$, some $B$ belongs to $\mathcal{B}_{1}\cap\bigcap_{i\in[s-1]}\operatorname{Sol}(\mathcal{J}_{i})$, $\mathcal{J}_{s}$ is crucial in $D^{(B,\bot)}$ and has a connected subinstance whose solution set is not subdirect. Take some $B\in\mathcal{B}_{1}$ and apply the inductive assumption to $\mathcal{I}^{\prime}$ and $D^{(B,\bot)}$. Since $\mathcal{I}^{\prime}$ does not satisfy (1b) there exists an expanded covering $\mathcal{J}_{1}$ of $\mathcal{I}^{\prime}$ such that $\mathcal{J}_{1}$ is crucial in $D^{(B,\bot)}$ and $\mathcal{J}_{1}$ has a connected subinstance whose solution set is not subdirect. If $\operatorname{Sol}(\mathcal{J}_{1})\cap\mathcal{B}_{1}=\varnothing$, then we are done. Otherwise, put $\mathcal{B}_{2}=\operatorname{Sol}(\mathcal{J}_{1})\cap\mathcal{B}_{1}$, choose some $B\in\mathcal{B}_{2}$ and apply the inductive assumption to $\mathcal{I}^{\prime}$ and $D^{(B,\bot)}$ to obtain $\mathcal{J}_{2}$. Since $\mathcal{B}_{1}$ is finite, and the sequence $\mathcal{B}_{1},\mathcal{B}_{2},\dots$ is decreasing, at some moment the required condition $\mathcal{B}_{1}\cap\bigcap_{i\in[s]}\operatorname{Sol}(\mathcal{J}_{s})=\varnothing$, will be satisfied.
Put $\Theta=\Delta((\bigwedge_{\mathcal{J}\in\Omega^{\prime}}\bot(\mathcal{J}))\wedge(\bigwedge_{i=1}^{s}\bot(\mathcal{J}_{s})))$. It follows from the definition that $\Theta$ is an expanded covering of $\mathcal{I}$ not having a solution in $D^{(1)}$. Let $\Theta^{\prime}$ be the weakening of $\Theta$ such that $\Theta^{\prime}$ is crucial in $D^{(1)}$. Notice that all the constraints of $\Theta$ that came from $\mathcal{J}_{s}$ are crucial in $D^{(1)}$, which means that they stay in $\Theta^{\prime}$. Therefore, $\Theta^{\prime}$ has a connected subinstance whose solution set is not subdirect. Thus, we proved that $\mathcal{I}$ satisfies (1c).
Subcase 2. Every instance $\mathcal{J}\in\Omega$ satisfies (1b). This implies that each instance $\bot(\mathcal{J})$ has a subdirect solution set. Notice that if $\Omega=\{\mathcal{I}\}$ then $\mathcal{I}$ satisfies (1b), which completes this case. Otherwise, both $\mathcal{J}^{(1)}$ and $(\bot(\mathcal{J}))^{(1)}$ have a solution for every $\mathcal{J}\in\Omega$. Since $\bot(\mathcal{J})$ is crucial in $D^{(B,\top)}$ for some $B\in\mathcal{B}_{0}$ (property 4(a) of $\Omega$), Lemma 39 implies that $T=\mathcal{L}$. Let us show that $\mathcal{I}$ is connected. Take some constraint $C\in\mathcal{I}$ and a variable $x$ of $C$. Put $\Theta=\Delta(\bigwedge_{\mathcal{J}\in\Omega}\bot(\mathcal{J}))$. Let us weaken $\Theta$ to make all the constraints except for the constraints $\sigma(z^{\prime},z^{\prime\prime})$ crucial in $D^{(1)}$. We do not weaken the constrains $\sigma(z^{\prime},z^{\prime\prime})$. The obtained instance we denote by $\Theta^{\prime}$. Condition 4 for $\Omega$ implies that all the constraints of $\Theta$ coming from some $\mathcal{J}\in\Omega$ are still in $\Theta^{\prime}$. Since $\mathcal{I}$ is crucial in $D^{(1)}$, a child of $C$ must appear in $\Theta^{\prime}$. Let the child appear in $\bot(\mathcal{J})$-part of $\Theta^{\prime}$. Denote the $\bot(\mathcal{J})$-part of $\Theta^{\prime}$ by $\Theta_{\mathcal{J}}^{\prime}$. By the cruciality of $C$, there exists $B\in\mathcal{B}_{0}$ such that $\Theta_{\mathcal{J}}^{\prime(B,\top)}$ has no solutions but a weakening of the child of $C$ gives a solution inside $D^{(B,\top)}$. Since $\bot(\mathcal{J})$ has a solution in $D^{(1)}$, $\Theta_{\mathcal{J}}^{\prime}$ also has a solution in $D^{(1)}$. Let $\mathcal{M}$ be a minimal set of children of $z$ we need to restrict to $B$ in $\Theta_{\mathcal{J}}^{\prime}$ to kill all the solutions of $\Theta_{\mathcal{J}}^{\prime(1)}$. We will build a bridge $\delta$ from $\sigma$ to $\operatorname{Con}(C,x)$ such that $\widetilde{\delta}$ is larger than a binary relation coming from some path from $z$ to $x$ in $\mathcal{I}$. Consider two subsubcases.
Subsubcase 1. The child of $C$ appears in a child of $\Upsilon_{y}$. Since $\Upsilon_{y}$ is a tree-covering and $\mathcal{I}^{(1)}$ is 1-consistent, the set $\mathcal{M}$ contains at least one variable from the child of $\Upsilon_{y}$. Applying Lemma 39 to the solution set of $\Theta_{\mathcal{J}}^{\prime}$, we get a required bridge from $\sigma$ to $\operatorname{Con}(C,x)$ and also prove that $\operatorname{Con}(C,x)$ is a linear congruence.
Subsubcase 2. The child of $C$ appears in $\mathcal{J}\in\Omega$. Let $y$ be the variable of $\Theta_{\mathcal{J}}^{\prime}$ such that some variable from $\mathcal{M}$ appears in a child of $\Upsilon_{y}$ in $\Theta_{\mathcal{J}}^{\prime}$. Applying Lemma 39 to the solution set of $\Theta_{\mathcal{J}}^{\prime}$, we get a bridge from $\sigma$ to $\xi$ for some $\xi\in\operatorname{Con}(\mathcal{J},y)$. Since $\mathcal{J}$ is connected, by Lemma 34(b) there is a corresponding bridge from $\xi$ to $\operatorname{Con}(C,x)$. Composing these bridges we get a required bridge $\delta$ from $\sigma$ to $\operatorname{Con}(C,x)$. Notice that we could also build a bridge $\delta$ from $\sigma$ to $\operatorname{Con}(C,x)$ without an intermediate step but $\widetilde{\delta}$ would not satisfy the required property.
Thus, for every constraint $C$ and every variable $x$ in $\mathcal{I}$ we have a bridge $\delta$ from $\sigma$ to $\operatorname{Con}(C,x)$ such that $\widetilde{\delta}$ is larger than a binary relation coming from some path from $z$ to $x$ in $\mathcal{I}$. To prove that $\mathcal{I}$ is connected we do the following. We take two constraints $C_{1}$ and $C_{2}$ with a common variables $x$. We proved that there is a bridge from $\operatorname{Con}(C_{1},x)$ to $\sigma$, and a bridge from $\operatorname{Con}(C_{2},x)$ to $\sigma$. Composing these bridges (and using cycle-consistency of $\mathcal{I}$) we obtain a required reflexive bridge. Hence $C_{1}$ and $C_{2}$ are adjacent, $\mathcal{I}$ is connected, and $\mathcal{I}$ satisfies (1b) if it has subdirect solution set, or (1c) otherwise. ∎
Theorem 42.
Suppose $\mathcal{I}$ is a cycle-consistent irreducible instance, $B<_{PC(\sigma)}^{D_{y}}D_{y}$ for some $y\in\operatorname{Var}(\mathcal{I})$, $\mathcal{I}$ has a solution. Then $\mathcal{I}$ has a solution with $y\in B$.
Proof.
For any $G<_{PC(\sigma)}^{D_{y}}D_{y}$ by $D^{(G,\top)}$ we denote the reduction of $\mathcal{I}$ such that $D_{z}^{(G,\top)}=G$ and $D_{x}^{(G,\top)}=D_{x}$ if $x\neq y$. For an expanded covering $\mathcal{J}$ of $\mathcal{I}$ by $\operatorname{Sol}(\mathcal{J})$ we denote the set of all $G\in D_{y}/\sigma$ such that $\mathcal{J}^{(G,\top)}$ has a solution.
Assume that $\mathcal{I}$ has no solutions with $y\in B$. Let $\mathcal{B}\subsetneq D_{y}/\sigma$ be an inclusion-maximal set such that $\operatorname{Sol}(\mathcal{J})=\mathcal{B}$ for some expanded covering $\mathcal{J}$ of $\mathcal{I}$. Let $\mathcal{J}$ be the expanded covering witnessing this. Choose $G\in(D_{y}/\sigma)\setminus\mathcal{B}$.
By Lemma 37 there exists a 1-consistent reduction for $\mathcal{I}$ smaller than $D^{(G,\top)}$. Since $\mathcal{J}$ is an expanded covering, the maximal 1-consistent reduction $D^{(G,\bot)}$ for $\mathcal{J}$ such that $D^{(G,\bot)}\leqslant D^{(G,\top)}$ is also nonempty. By Lemma 35 for every $x\in\operatorname{Var}(\mathcal{J})$ there exists a tree-covering $\Upsilon_{x}$ of $\mathcal{J}$ such that $\Upsilon_{x}^{(G,\top)}(x)$ defines $D_{x}^{(G,\bot)}$. Notice that the reduction $D^{(G,\top)}$ was defined for $\mathcal{I}$ and then extended to $\mathcal{J}$ but $D^{(G,\bot)}$ was originally defined for $\mathcal{J}$ and does not exist for $\mathcal{I}$.
Weaken $\mathcal{J}$ to make it crucial in $D^{(G,\bot)}$ and denote the obtained instance by $\mathcal{J}^{\prime}$. By Theorem 41 applied to $\mathcal{J}^{\prime}$ and $D^{(G,\bot)}$, $\mathcal{J}^{\prime}$ satisfies (1b) or (1c).
Assume that $\mathcal{J}^{\prime}$ satisfies (1b). Then the solution set of $\mathcal{J}^{\prime}$ is subdirect. Put $\mathcal{J}^{\prime\prime}=\mathcal{J}^{\prime}\wedge\bigwedge_{x\in\operatorname{Var}(\mathcal{J}^{\prime})}\Upsilon_{x}$. Notice that $\mathcal{J}^{\prime\prime}$ is an expanded covering of $\mathcal{I}$ with a subdirect solution set. Since $\mathcal{J}^{\prime\prime(G,\top)}$ has no solutions, $\mathcal{J}^{\prime\prime}$ has a solution (as $\mathcal{I}$ has a solution), and any weakening of a constraint from $\mathcal{J}^{\prime}$ inside $\mathcal{J}^{\prime\prime}$ gives an instance with a solution in $D^{(G,\top)}$, Lemma 39 implies that the type $\mathcal{PC}$ coincides with the type of the crucial constraints, which is linear by (1b). This contradiction completes this case.
Assume that $\mathcal{J}^{\prime}$ satisfies (1c). Let $\Theta$ be the expanded covering of $\mathcal{J}^{\prime}$ that is crucial in $D^{(G,\bot)}$ and $\Upsilon$ be the linked connected subinstance of $\Theta$. Put $\Theta^{\prime}=\Theta\wedge\bigwedge_{x\in\operatorname{Var}(\Theta)}\Upsilon_{x}$. Notice that $\Theta^{\prime}$ is an expanded covering of $\mathcal{I}$ with a subdirect solution set and $\Theta^{\prime}$ has no solutions in $D^{(G,\top)}$. Let $x$ be a variable of a constraint $C\in\Upsilon$. By Lemma 34(p), $\operatorname{Con}(C,x)$ is a perfect linear congruence and there exists $\zeta\leqslant\mathbf{D}_{x}\times\mathbf{D}_{x}\times\mathbf{Z}_{p}$ such that $(y_{1},y_{2},0)\in\zeta\Leftrightarrow(y_{1},y_{2})\in\operatorname{Con}(C,x)$ and $\operatorname{pr}_{1,2}(\zeta)={{\operatorname{Con}(C,x)}^{*}}$. Let us replace the variable $x$ of $C$ in $\Theta^{\prime}$ by $x^{\prime}$ and add the constraint $\zeta(x,x^{\prime},z)$. The obtained instance we denote by $\Theta^{\prime\prime}$. Another instance we build from $\Theta^{\prime\prime}$ by replacing $\zeta(x,x^{\prime},z)$ by ${{\omega}^{*}}(x,x^{\prime})$, where $\omega=\operatorname{Con}(C,x)$. We denote it by $\Theta^{\prime\prime\prime}$. Since $\Theta^{\prime\prime\prime}$ is an expanded covering of $\mathcal{J}$ we have $\operatorname{Sol}(\Theta^{\prime\prime\prime})\supseteq\operatorname{Sol}(\mathcal{J})$. Since the weakening of $C$ in $\Theta$ gives an instance with a solution in $D^{(G,\bot)}$ and $\mathcal{B}$ was chosen maximal, we have $\operatorname{Sol}(\Theta^{\prime\prime\prime})=D_{y}/\sigma$. Let $R\leqslant D_{y}/\sigma\times\mathbf{Z}_{p}$ be the set of all pairs $(F,j)$ such that $\Theta^{\prime\prime}$ has a solution in $D^{(F,\top)}$ with $z=j$. We know that $R$ is subdirect, $(G,0)\notin R$, $(G,j)\in R$ for some $j\in Z_{p}$. Applying Corollary 22 to $R$, $\{G\}<_{\mathcal{PC}}D_{y}/\sigma$, and $\{0\}<_{\mathcal{L}}\mathbf{Z}_{p}$, we get a contradiction as we mixed linear and PC types. ∎
3.4 Statements sufficient to prove that Zhuk’s algorithm works
Theorem 43.
Suppose $\Theta$ is a cycle-consistent irreducible CSP instance, and $B<_{T}^{D_{x}}D_{x}$, where $T\in\{BA,C,PC\}$. Then $\Theta$ has a solution if and only if $\Theta$ has a solution with $x\in B$.
Proof.
For $T=PC$ it follows from Theorem 42. Assume that $T\in\{BA,C\}$. By Lemma 37 there exists a 1-consistent reduction $D^{(1)}\leqslant_{T}D$ such that $D_{x}^{(1)}\leqslant B$. By Theorem 41 $\Theta^{(1)}$ has a solution, and therefore $\Theta$ has a solution with $x\in B$. ∎
Theorem 44.
Suppose the following conditions hold:
  1. 1.
    $\mathcal{I}$ is a linked cycle-consistent irreducible CSP instance with $\operatorname{Var}(\mathcal{I})=\{x_{1},\dots,x_{n}\}$;
  2. 2.
    $D_{x_{i}}$ is S-free for every $i\in[n]$;
  3. 3.
    if we weaken all the constraints of $\Theta$, we get an instance whose solution set is subdirect.
  4. 4.
    $\sigma_{x_{i}}$ is the intersection of all the linear congruences $\sigma$ on $D_{x_{i}}$ such that $\sigma^{*}=D_{x_{i}}\times D_{x_{i}}$.
  5. 5.
    $L_{x_{i}}=D_{x_{i}}/\sigma_{x_{i}}$ for every $i\in[n]$;
  6. 6.
    $\phi:\mathbf{Z}_{q_{1}}\times\dots\times\mathbf{Z}_{q_{k}}\to L_{x_{1}}\times\dots\times L_{x_{n}}$ is a homomorphism, where $q_{1},\dots,q_{k}$ are prime numbers;
  7. 7.
    if we weaken any constraint of $\mathcal{I}$ then for every $(a_{1},\ldots,a_{k})\in\mathbf{Z}_{q_{1}}\times\dots\times\mathbf{Z}_{q_{k}}$ there exists a solution of the obtained instance in $\phi(a_{1},\ldots,a_{k})$.
Then $\{(a_{1},\dots,a_{k})\mid\Theta\text{ has a solution in }\phi(a_{1},\dots,a_{k})\}$ is either empty, or is full, or is an affine subspace of $\mathbf{Z}_{q_{1}}\times\dots\times\mathbf{Z}_{q_{k}}$ of codimension 1 (the solution set of a single linear equation).
Proof.
Put $\Delta=\{(a_{1},\dots,a_{k})\mid\Theta\text{ has a solution in }\phi(a_{1},\dots,a_{k})\}$. If $\Delta$ is full then we are done. Otherwise, consider $(b_{1},\ldots,b_{k})\in(\mathbf{Z}_{q_{1}}\times\dots\times\mathbf{Z}_{q_{k}})\setminus\Delta$. Notice that $\phi(b_{1},\ldots,b_{k})$ can be viewed as a reduction for $\mathcal{I}$. We denote this reduction by $D^{(1)}$. It follows from condition (7) that $\mathcal{I}$ is crucial in $D^{(1)}$.
Let us prove that there exists a constraint $C\in\mathcal{I}$ and its variable $x$ such that $\operatorname{Con}(C,x)$ is a perfect linear congruence. By Lemma 31 either $C_{0}^{(1)}$ is empty for some $C_{0}\in\mathcal{I}$, or the reduction $D^{(1)}$ is 1-consistent for $\mathcal{I}$. Consider two cases.
Case 1. $C_{0}^{(1)}$ is empty. Since $\mathcal{I}$ is crucial in $D^{(1)}$, it consists of just one constraint $C_{0}$. Let $C_{0}=R(y_{1},\dots,y_{t})$. By Lemma 38 $R$ has the parallelogram property and $\operatorname{Con}(R,1)$ is a linear congruence such that $\operatorname{Con}(R,1)^{*}=D_{y_{1}}^{2}$. By Lemma 11 $\mathbf{D}_{y_{1}}/\delta\cong\mathbf{Z}_{p}$. Let $\psi\colon\mathbf{D}_{y_{1}}\to\mathbf{Z}_{p}$ be the homomorphism. Then the required ternary relation $\zeta\leqslant\mathbf{D}_{y_{1}}\times\mathbf{D}_{y_{1}}\times\mathbf{Z}_{p}$ can be defined by $\{(a_{1},a_{2},b)\mid\psi(a_{1})-\psi(a_{2})=b\}$. Hence $\operatorname{Con}(R,1)$ is a perfect linear congruence.
Case 2. The reduction $D^{(1)}$ is 1-consistent. By Theorem 41, every constraint of $\mathcal{I}$ has the parallelogram property and satisfies condition (1b) or (1c). If $\mathcal{I}$ satisfies (1c) then there exists an instance $\Theta\in\operatorname{ExpCov}(\mathcal{I})$ that is crucial in $D^{(1)}$ and contains a linked connected subinstance $\Upsilon$ such that the solution set of $\Upsilon$ is not subdirect. By condition 4, since the solution set of $\Upsilon$ is not subdirect, $\Upsilon$ must contain a constraint relation from the original instance $\mathcal{I}$. Applying Lemma 34(p), we derive that $\operatorname{Con}(C,x)$ is a perfect linear congruence for the corresponding child of the original constraint and its variable. If $\mathcal{I}$ satisfies (1b), then $\mathcal{I}$ is linked connected itself and the existence of a perfect linear congruence again follows from Lemma 34(p).
Thus, $\operatorname{Con}(C,x)$ is a perfect linear congruence for some $C\in\mathcal{I}$ and its variable $x$. Let $\zeta$ be the corresponding ternary relation. We add a new variable $z$ with domain $\mathbf{Z}_{p}$, replace the variable $x$ in $C$ by $x^{\prime}$, and add the constraint $\zeta(x,x^{\prime},z)$. We denote the obtained instance by $\mathcal{I}^{\prime}$. Let $L$ be the set of all tuples $(a_{1},\ldots,a_{k},b)\in\mathbf{Z}_{q_{1}}\times\dots\times\mathbf{Z}_{q_{k}}\times\mathbf{Z}_{p}$ such that $\mathcal{I}^{\prime}$ has a solution with $z=b$ in $\phi(a_{1},\ldots,a_{k})$. Notice that $L\leqslant\mathbf{Z}_{q_{1}}\times\dots\times\mathbf{Z}_{q_{k}}\times\mathbf{Z}_{p}$. By condition 7, the projection of $L$ onto the first $k$ coordinates is a full relation and $(b_{1},\dots,b_{k},0)\notin L$. Therefore $L$ has dimension $k$ and can be defined by one linear equation. If this equation is $z=b$ for some $b\neq 0$, then $\Delta$ is empty. Otherwise, we put $z=0$ in this equation and get an equation describing all $(a_{1},\ldots,a_{k})$ such that $\mathcal{I}$ has a solution in $\phi(a_{1},\ldots,a_{k})$. Hence the dimension of $\Delta$ is $k-1$. ∎
4 XY-symmetric operations
In this section we prove that a weak near unanimity operation of an odd arity implies an operation that is symmetric on all tuples having exactly two different elements. The idea of the proof is to generate a relation such that the existence of an XY-symmetric operation was equivalent to existence of a block-constant tuple. We gradually reduce coordinates of this relation to strong subalgebras trying to achieve this tuple. If we cannot make the next reduction, it means that we found a linear congruence such that there is no block-constant tuple even modulo this congruence. Since this linear congruence lives on the original domain, and the generated relation must be linked, we immediately obtain a perfect linear congruence. This allows us to represent the domain as a product of a smaller domain $B$ (where we have an XY-symmetric operation by the inductive assumption) and $\mathbf{Z}_{p}$. The rest of the proof is purely operational: we start with an XY-symmetric operation on $B$ and show how composing this operation with itself we can gradually increase the number of tuples where it behaves well on the whole domain.
This section is organised as follows. First, we explain how we define the relation for a tuple of algebras, how we apply and denote reductions. We also define symmetries this relation has and $\boxtimes$-product of $\mathbf{B}$ and $\mathbf{Z}_{p}$. In Subsection 4.2 we show how to derive the main result from three theorems that are proved later. In the next section we prove two out of three theorems explaining how to build a smaller reduction if possible, and how to build a reduction if it is known that an XY-symmetric operation exists. Finally, in Subsection 4.4 we show how to improve an operation gradually to make it XY-symmetric on $\mathbf{B}\boxtimes\mathbf{Z}_{p}$ even if originally it was XY-symmetric only on $B$.
4.1 Definitions
The free generated relation $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. For a tuple of algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$ by $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ we denote the relation of arity $N:=(2^{n-1}-1)\cdot\sum\limits_{i=1}^{s}|\mathbf{A}_{i}|\cdot(|\mathbf{A}_{i}|-1)$ defined as follows. Coordinates of the relation are indexed by $(\mathbf{A}_{i},\alpha)$, where $\alpha\in\{a,b\}^{n}$ for some $a,b\in A_{i}$, $a\neq b$. The set of all indexes denote by $I$. For a set of tuples $S$ by $\operatorname{TwoTuples}(S)$ we denote the set of tuples from $S$ having exactly 2 different elements. Then $I=\{(\mathbf{A}_{i},\alpha)\mid i\in[s],\alpha\in\operatorname{TwoTuples}(A^{n})\}$. For $i\in[n]$ by $\gamma_{i}$ we denote the tuple of length $N$ whose $(\mathbf{A}_{i},\alpha)$-th element is equal to $\alpha(i)$ for every $(\mathbf{A}_{i},\alpha)\in I$. Then $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{n}}$ is the minimal subuniverse of $\prod\limits_{i\in[s]}\mathbf{A}_{i}^{(2^{n-1}-1)|A_{i}|\cdot|A_{i}-1|}$ containing $\gamma_{1},\dots,\gamma_{n}$. We also say that $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{n}}$ is the subalgebra generated by $\gamma_{1},\dots,\gamma_{n}$, and the tuples $\gamma_{1},\dots,\gamma_{n}$ are called the generators of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{n}}$.
We will use terminology similar to the one we used in the previous section. For every $(\mathbf{A}_{i},\alpha)\in I$ by $\mathbf{D}_{(\mathbf{A}_{i},\alpha)}^{(0)}$ we denote the subalgebra of $\mathbf{A}_{i}$ generated by elements of $\alpha$. Notice that $\operatorname{pr}_{i}(R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}})=D_{i}^{(0)}$ for every $i\in I$. By $\mathcal{R}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ we denote the set of all relations $R$ of arity $N$ whose coordinates are indexed by $I$ such that the domain of the $i$-th coordinate of $R$ is $D_{i}^{(0)}$ for every $i\in I$.
Reductions. In our proof we reduce the relation $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ by reducing their coordinates. A reduction $D^{(\top)}$ for $R\in\mathcal{R}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ is a mapping that assigns a subuniverse $D_{i}^{(\top)}\leqslant D_{i}^{(0)}$ to every $i\in I$. $D^{(0)}$ can be viewed as a trivial reduction. As in the previous section we write $D^{(\bot)}\lll D^{(\top)}$ and $D^{(\bot)}\leqslant_{T}D^{(\top)}$ whenever $D_{i}^{(\bot)}\lll D_{i}^{(\top)}$ for every $i\in I$ and $D_{i}^{(\bot)}\leqslant_{T}D_{i}^{(\top)}$ for every $i\in I$, respectively. Notice that any reduction $D^{(\top)}$ can be viewed as a relation from $\mathcal{R}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Then for any $R\in\mathcal{R}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and a reduction $D^{(\bot)}$ by $R^{(\bot)}$ we denote $R\cap D^{(\bot)}$. A reduction $D^{(\bot)}$ is called 1-consistent for $R\in\mathcal{R}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ if $\operatorname{pr}_{i}(R^{(\bot)})=D_{i}^{(\bot)}$ for every $i\in I$.
$k$-WNU. An operation $f:A^{n}\to A$ is called a $k$-WNU operation if it is symmetric on $(\underbrace{x,x,\dots,x}_{k},y,y,\dots,y)$. Then, 1-WNU is just a usual WNU.
Permutations and symmetries. For a tuple $\alpha\in A^{n}$ by $\operatorname{Perm}(\alpha)$ we denote the set of all tuples that can be obtained from $\alpha$ by a permutation of elements. For an index $i=(\mathbf{A}_{j},\alpha)$ by $\operatorname{Perm}(i)$ we denote the set of indexes $(\mathbf{A}_{j},\beta)$ with $\beta\in\operatorname{Perm}(\alpha)$. For a tuple $\alpha\in A^{n}$ and a permutation on $[n]$ by $\sigma(\alpha)$ we denote the tuple $\alpha^{\prime}$ such that $\alpha^{\prime}(j)=\alpha(\sigma(j))$ for every $j\in[n]$. For a tuple $\gamma$ of arity $N$ whose coordinates are indexed by elements from $I$ and a permutation $\sigma$ on $[n]$ by $\gamma^{\sigma}$ we denote the tuple $\gamma^{\prime}$ such that $\gamma^{\prime}((\mathbf{A}_{i},\alpha))=\gamma((\mathbf{A}_{i},\sigma(\alpha)))$ for any $(\mathbf{A}_{i},\alpha)\in I$. Similarly, for a relation $R\in\mathcal{R}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ put $R^{\sigma}=\{\gamma^{\sigma}\mid\gamma\in R\}$. A relation $R$ is called $\sigma$-symmetric if $R^{\sigma}=R$. A relation $R$ is called symmetric if it is $\sigma$ symmetric for every permutation $\sigma$ on $[n]$. Similarly, a reduction $D^{(\top)}$ is called symmetric if $D_{i}^{(\top)}=D_{j}^{(\top)}$ for any $j\in\operatorname{Perm}(i)$.
$\boxtimes$-product of $\mathbf{B}$ and $\mathbf{Z}_{p}$. For $x=(a,b)$ by $x^{(1)}$ and $x^{(2)}$ we denote $a$ and $b$ respectively. For an algebra $\mathbf{B}=(B;w^{\mathbf{B}})$ by $\mathbf{B}\boxtimes\mathbf{Z}_{p}$ we denote the set of algebras $\mathbf{A}$ such that $A=B\times Z_{p}$, $(w^{\mathbf{A}}(x_{1},\dots,x_{n}))^{(1)}=w^{\mathbf{B}}(x_{1}^{(1)},\dots,x_{n}^{(1)})$ and $(w^{\mathbf{A}}(x_{1},\dots,x_{n}))^{(2)}=f(x_{1}^{(1)},\dots,x_{n}^{(1)})+a_{1}x_{1}^{(2)}+\dots+a_{n}x_{n}^{(2)}$ for some mapping $f\colon B^{n}\to\mathbf{Z}_{p}$ and $a_{1},\dots,a_{n}\in\mathbf{Z}_{p}$.
4.2 Proof of the main result
Theorem 45.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$, $n$ is odd, $D^{(1)}$ is a 1-consistent symmetric reduction of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, $D^{(1)}\lll D^{(0)}$. Then one of the following conditions hold
  1. 1.
    $|D^{(1)}_{(\mathbf{A}_{i},\alpha)}|=1$ for all $(\mathbf{A}_{i},\alpha)$.
  2. 2.
    there exists a 1-consistent symmetric reduction $D^{(2)}$ for $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ such that $D^{(2)}\lll D^{(1)}$ and $D^{(2)}\neq D^{(1)}$.
  3. 3.
    there exists a perfect linear congruence $\sigma$ on some $D^{(0)}_{(\mathbf{A}_{i},\alpha)}$ such that
    1. (a)
      $D_{(\mathbf{A}_{i},\alpha)}^{(1)}\times D_{(\mathbf{A}_{i},\alpha)}^{(1)}\not\subseteq\sigma$
    2. (b)
      $D_{(\mathbf{A}_{i},\alpha)}^{(1)}\times D_{(\mathbf{A}_{i},\alpha)}^{(1)}\subseteq\sigma^{*}$
Theorem 46.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$, $n$ is odd, there exists an $n$-ary term $\tau_{0}$ such that $\tau_{0}^{\mathbf{A}_{i}}$ is XY-symmetric for every $i$. Then there exists a 1-consistent symmetric reduction $D^{(\triangle)}\lll D^{(0)}$ of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and an $n$-ary term $\tau$ such that $\tau^{\mathbf{A}_{i}}$ is XY-symetric and $D_{(\mathbf{A}_{i},\alpha)}^{(\triangle)}=\{\tau(\alpha)\}$ for every $i$ and $\alpha\in\operatorname{TwoTuples}(A_{i}^{n})$.
Theorem 47.
Suppose $\mathbf{A},\mathbf{B}\in\mathcal{V}_{n}$, $0_{\mathbf{A}}$ is a perfect linear congruence, $\mathbf{A}/{0_{\mathbf{A}}}^{*}\times\mathbf{B}$ has an XY-symmetric term operation of arity $n$. Then $\mathbf{A}\times\mathbf{B}$ has an XY-symmetric term operation.
Theorem 48.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$, $n$ is odd. Then there exists a term $\tau$ such that $\tau^{\mathbf{A}_{i}}$ is an XY-symmetric operation for every $i$.
Proof.
First, we reorder algebras so that $|A_{1}|\geqslant|A_{2}|\geqslant\dots\geqslant|A_{s}|$. We prove the claim by induction on the size of algebras. Precisely, we assign an infinite tuple $(|A_{1}|,|A_{2}|,\dots,|A_{s}|,0,0,\dots)$ to the sequence of algebras, and our inductive assumption is that the statement holds for algebras $\mathbf{A}_{1}^{\prime},\dots,\mathbf{A}_{t}^{\prime}\in\mathcal{V}_{n}$ such that $(|A_{1}^{\prime}|,|A_{2}^{\prime}|,\dots,|A_{t}^{\prime}|,0,0,\dots)<(|A_{1}|,|A_{2}|,\dots,|A_{s}|,0,0,\dots)$ (lexicographic order).
The base of our induction is the case when $|A_{1}|=|A_{2}|=\dots=|A_{s}|=1$, which is obvious.
Let us prove the inductive step. First, we add all nontrivial subalgebras of $\mathbf{A}_{1}$ to the list $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ and prove even stronger claim. We do not want to introduce new notations that is why we assume that $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ contains all nontrivial subalgebras of $\mathbf{A}_{1}$. Consider two cases.
Case 1. Suppose $\mathbf{A}_{1}$ has two nontrivial congruences $\sigma$ and $\delta$ such that $\sigma\cap\delta$ is the equality relation (0-congruence) on $\mathbf{A}_{1}$. Consider algebras $\mathbf{A}_{1}/\sigma,\mathbf{A}_{1}/\delta,\mathbf{A}_{2},\mathbf{A}_{3},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$ and apply the inductive assumption. Then there exists a term $t$ such that $t^{\mathbf{A}_{i}}$ is XY-symmetric for every $i\geqslant 2$, $t^{\mathbf{A}_{1}/\sigma}$ and $t^{\mathbf{A}_{1}/\delta}$ are XY-symmetric. Therefore $t^{\mathbf{A}_{1}}$ is also XY-symmetric, which completes the proof.
Case 2. There exists a unique minimal nontrivial congruence $\delta$ on $\mathbf{A}_{1}$. By the inductive assumption, there exists a term $\tau_{0}$ such that $\tau_{0}^{\mathbf{A}_{1}/\delta}$ and $\tau_{0}^{\mathbf{A}_{i}}$ for $i\geqslant 2$ are XY-symmetric. By Theorem 46 there exists a 1-consistent symmetric reduction $D^{(\top)}$ for $R_{\mathbf{A}_{1}/\delta,\mathbf{A}_{2},\dots,\mathbf{A}_{s}}$ and an $n$-ary term $\tau_{1}$ satisfying the corresponding condition. We define a new reduction for $R_{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{s}}$ as follows. We put $D^{(1)}_{(\mathbf{A}_{i},\alpha)}=D^{(\top)}_{(\mathbf{A}_{i},\alpha)}$ for $i\geqslant 2$, and $D^{(1)}_{(\mathbf{A}_{1},\alpha)}=E$ whenever $D^{(\top)}_{(\mathbf{A}_{1},\alpha/\delta)}=\{E\}$. Applying term $\tau_{1}$ to the generators of $R_{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{s}}$ we obtain a tuple $\gamma\in R_{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{s}}^{(1)}.$ To make the reduction $D^{(1)}$ 1-consistent, put $D^{(2)}_{(\mathbf{A}_{i},\alpha)}=\operatorname{pr}_{(\mathbf{A}_{i},\alpha)}R_{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{s}}^{(1)}$. By Corollary 17(t) and Corollary 18(r1) we have $D^{(2)}\lll D^{(0)}$.
Notice that $|D^{(2)}_{(\mathbf{A}_{i},\alpha)}|=1$ for $i\geqslant 2$. By Theorem 45 we have one of the three cases. In case 2, we can apply Theorem 45 again and obtain even smaller reduction. Since we cannot reduce forever, we end up with one of the two subcases.
Subcase 1. There exists a symmetric 1-consistent reduction $D^{(3)}\lll D^{(2)}$ such that $|D_{(\mathbf{A}_{i},\alpha)}^{(3)}|=1$. Take the tuple $\gamma\in R_{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{s}}^{(3)}$ and a term $\tau$ giving $\gamma$ on the generators of $R_{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{s}}$. It follows from the symmetricity of $D^{(3)}$ that $\tau^{\mathbf{A}_{i}}$ is XY-symmetric for every $i$.
Subcase 2. There exists a perfect linear congruence $\sigma$ on some $\mathbf{D}^{(0)}_{(\mathbf{A}_{1},\alpha)}$ such that $D^{(2)}_{(\mathbf{A}_{1},\alpha)}\times D^{(2)}_{(\mathbf{A}_{1},\alpha)}\not\subseteq\sigma$. Assume that $D^{(0)}_{(\mathbf{A}_{1},\alpha)}\neq A_{1}$. Since we assumed that all subalgebras of $\mathbf{A}_{1}$ are in the list, there exists $k$ such that $\mathbf{A}_{k}=\mathbf{D}^{(0)}_{(\mathbf{A}_{1},\alpha)}$. By the definition of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ we have $\gamma(\mathbf{A}_{k},\alpha)=\gamma(\mathbf{A}_{1},\alpha)$ for all $\gamma\in R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Since the reduction $D^{(2)}$ is 1-consistent we obtain that $|\mathbf{D}^{(2)}_{(\mathbf{A}_{1},\alpha)}|=|\mathbf{D}^{(2)}_{(\mathbf{A}_{k},\alpha)}|=1$, which contradicts $D^{(2)}_{(\mathbf{A}_{1},\alpha)}\times D^{(2)}_{(\mathbf{A}_{1},\alpha)}\not\subseteq\sigma$. Thus, $D^{(0)}_{(\mathbf{A}_{1},\alpha)}=A_{1}$ and $\sigma$ is a perfect linear congruence on $\mathbf{A}_{1}$. Since $D^{(2)}_{(\mathbf{A}_{1},\alpha)}\times D^{(2)}_{(\mathbf{A}_{1},\alpha)}\not\subseteq\sigma$ and $D^{(2)}_{(\mathbf{A}_{1},\alpha)}$ is smaller than or equal to an equivalence block of $\delta$, we have $\delta\not\subseteq\sigma$. Since $\delta$ is the minimal nontrivial congruence, we obtain that $\sigma=0_{\mathbf{A}_{1}}$. Applying Theorem 47 to $\mathbf{A}_{1}/0_{\mathbf{A}_{1}}^{*}\times\mathbf{A}_{2}\times\mathbf{A}_{3}\times\dots\times\mathbf{A}_{s}$ we obtain a term $\tau$ such that $\tau^{\mathbf{A}_{i}}$ is XY-symmetric for every $i\in[s]$. ∎
Theorem 2.
Suppose $f$ is a WNU of an odd arity $n$ on a finite set. Then there exists an XY-symmetric operation $f^{\prime}\in\operatorname{Clo}(\{f\})$ of arity $n$.
Proof.
Let $f$ be an operation on a finite set $A$. By Lemma 26 there exists a special WNU $w\in\operatorname{Clo}(f)$ of arity $N=n^{n!}$. Consider the algebra $\mathbf{A}=(A;w)\in\mathcal{V}_{N}$. By Theorem 48 there exists an $N$-ary operation $w^{\prime}\in\operatorname{Clo}(w)$ such that $w^{\prime}$ is XY-symmetric. Then the required $n$-ary XY-symmetric operation can be defined by
$f^{\prime}(x_{1},\dots,x_{n})=w^{\prime}(\underbrace{x_{1},\dots,x_{1}}_{n^{n!-1}},\underbrace{x_{2},\dots,x_{2}}_{n^{n!-1}},\dots,\underbrace{x_{n},\dots,x_{n}}_{n^{n!-1}})$
4.3 Proof of Theorems 45 and 46 (Finding a reduction)
Theorem 49.
Suppose
  1. 1.
    $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$, where $n$ is odd.
  2. 2.
    $D^{(1)}$ is a 1-consistent symmetric reductions of a symmetric relation $R\in\mathcal{R}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$,
  3. 3.
    $D^{(1)}\lll D^{(0)}$,
  4. 4.
    $B<_{\mathcal{T}}^{D^{(0)}_{(\mathbf{A}_{j},\beta)}}D^{(1)}_{(\mathbf{A}_{j},\beta)}$, where $\mathcal{T}\in\{\mathcal{BA},\mathcal{C},\mathcal{PC}\}$.
Then there exists a 1-consistent symmetric reduction $D^{(2)}$ for $R$ such that $D^{(2)}\lll D^{(1)}$ and $D^{(2)}\neq D^{(1)}$. Moreover, $D^{(2)}\leqslant_{\mathcal{T}}D^{(1)}$ if $\mathcal{T}\neq\mathcal{PC}$.
Proof.
For a relation $S$ whose variables are indexed with $(\mathbf{A}_{j},\beta)$ by $S\downarrow^{(\mathbf{A}_{j},\beta)}_{B}$ denote $S$ whose coordinate $(\mathbf{A}_{j},\beta)$ is restricted to $S$.
Condition 4 just says that there exists $B$ of type $\mathcal{T}\in\{\mathcal{BA},\mathcal{C},\mathcal{PC}\}$. We want to choose BA or central subuniverse if possible and $\mathcal{PC}$, only if none of the domains $D^{(1)}_{(\mathbf{A}_{j},\beta)}$ have proper BA or central subuniverse. Thus, below we assume that $D^{(1)}_{(\mathbf{A}_{j},\beta)}$ is S-free whenever $\mathcal{T}=\mathcal{PC}$.
Choose an index $(\mathbf{A}_{j},\beta)$ and $B<_{\mathcal{T}}^{D^{(0)}_{(\mathbf{A}_{j},\beta)}}D^{(1)}_{(\mathbf{A}_{j},\beta)}$ such that $R\downarrow^{(\mathbf{A}_{j},\beta)}_{B}$ is inclusion maximal.
By Corollary 18(r) we have $R^{(1)}\lll^{R}R$. By Lemma 14(b,bt) we have $R\downarrow^{(\mathbf{A}_{j},\beta)}_{B}<_{\mathcal{T}}^{R}R\downarrow^{(\mathbf{A}_{j},\beta)}_{D^{(1)}_{(\mathbf{A}_{j},\beta)}}\lll R$. By Lemma 20(it) $R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B}\leqslant_{\mathcal{T}}^{R}R^{(1)}$. Choose $\alpha\in\operatorname{Perm}(\beta)$ and put $C=\operatorname{pr}_{(\mathbf{A}_{i},\alpha)}(R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B})$. Then Lemma 14(ft) implies $C\leqslant_{\mathcal{T}}^{D_{(\mathbf{A}_{i},\alpha)}^{(0)}}D_{(\mathbf{A}_{i},\alpha)}^{(1)}$. Because of the choice of $(\mathbf{A}_{j},\beta)$ and $B$, if $C\neq D_{(\mathbf{A}_{i},\alpha)}^{(1)}$, then $R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B}=R^{(1)}\downarrow^{(\mathbf{A}_{i},\alpha)}_{C}$. If additionally $\sigma(\alpha)=\alpha$ for some permutation on $[n]$, then from the symmetricity of $R$ we derive that
$(R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B})^{\sigma}=(R^{(1)}\downarrow^{(\mathbf{A}_{i},\alpha)}_{C})^{\sigma}=R^{(1)}\downarrow^{(\mathbf{A}_{i},\alpha)}_{C}=R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B}.$
Let us consider two cases.
Case 1. There exists $\alpha\in\operatorname{Perm}(\beta)$ such that $\alpha\neq\beta$ and $\operatorname{pr}_{(\mathbf{A}_{j},\alpha)}(R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B})\neq\operatorname{pr}_{(\mathbf{A}_{j},\alpha)}(R^{(1)})$. Then $R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B}$ is $\sigma$-symmetric for any permutation $\sigma$ preserving $\alpha$ and any permutation $\sigma$ preserving $\beta$. Since we can compose such permutations, $\alpha\neq\beta$, and $n$ is odd, we derive that $R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B}$ is $\sigma$-symmetric for any $\sigma$, hence $R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B}$ is just symmetric. Define a new reduction $D^{(2)}$ for $R$ by $D^{(2)}_{(\mathbf{A}_{i},\gamma)}:=\operatorname{pr}_{(\mathbf{A}_{i},\gamma)}(R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B})$. By Corollary 18(r1) $D^{(2)}\lll D^{(1)}$, and by Lemma 19 $D^{(2)}\leqslant_{\mathcal{T}}D^{(1)}$ for $\mathcal{T}\in\{\mathcal{BA},\mathcal{C}\}$, which completes this case.
Case 2. For any $\alpha\in\operatorname{Perm}(\beta)$ such that $\alpha\neq\beta$ we have $\operatorname{pr}_{(\mathbf{A}_{j},\alpha)}(R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B})=\operatorname{pr}_{(\mathbf{A}_{j},\alpha)}(R^{(1)})$. Put $S=\bigcap\limits_{\alpha\in\operatorname{Perm}(\beta)}R^{(1)}\downarrow^{(\mathbf{A}_{j},\alpha)}_{B}$. $S$ can be represented as an intersection of two symmetric relations, hence $S$ is also symmetric. If $S$ is not empty then we define a new reduction $D^{(2)}$ for $R$ by $D^{(2)}_{(\mathbf{A}_{i},\gamma)}:=\operatorname{pr}_{(\mathbf{A}_{i},\gamma)}(S)$. By Corollary 18(r1) $D^{(2)}\lll D^{(1)}$, and by Lemma 19 $D^{(2)}\leqslant_{\mathcal{T}}D^{(1)}$ if $\mathcal{T}\in\{\mathcal{BA},\mathcal{C}\}$. Thus, the only remaining case is when $S$ is empty. By Corollary 22, there should be $(\mathbf{A}_{j},\alpha_{1})$ and $(\mathbf{A}_{j},\alpha_{2})$ such that $\alpha_{1},\alpha_{2}\in\operatorname{Perm}(\beta)$ and $R^{(1)}\downarrow^{(\mathbf{A}_{j},\alpha_{1})}_{B}\cap R^{(1)}\downarrow^{(\mathbf{A}_{j},\alpha_{2})}_{B}=\varnothing$. Let $\beta=\sigma(\alpha_{1})$. Since $\operatorname{pr}_{(\mathbf{A}_{j},\alpha)}(R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B})=\operatorname{pr}_{(\mathbf{A}_{j},\alpha)}(R^{(1)})$ for any $\alpha\in\operatorname{Perm}(\beta)$ such that $\alpha\neq\beta$, we have
$(R^{(1)}\downarrow^{(\mathbf{A}_{j},\alpha_{1})}_{B}\cap R^{(1)}\downarrow^{(\mathbf{A}_{j},\alpha_{2})}_{B})^{\sigma}=(R^{(1)}\downarrow^{(\mathbf{A}_{j},\alpha_{1})}_{B})^{\sigma}\cap(R^{(1)}\downarrow^{(\mathbf{A}_{j},\alpha_{2})}_{B})^{\sigma}$ $=$
$R^{(1)}\downarrow^{(\mathbf{A}_{j},\beta)}_{B}\cap$ $R^{(1)}\downarrow^{(\mathbf{A}_{j},\sigma(\alpha_{2}))}_{B}\neq\varnothing.$
This contradiction completes the proof. ∎
Lemma 50.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$, $n$ is odd, $D^{(1)}$ is a 1-consistent symmetric reduction of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, $D^{(1)}\lll D^{(0)}$, $B<_{\mathcal{D}(\sigma)}^{D_{(\mathbf{A}_{i},\alpha)}^{(0)}}D_{(\mathbf{A}_{i},\alpha)}^{(1)}$ for some $(\mathbf{A}_{i},\alpha)$, $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(1)}$ has no tuple $\gamma$ such that $\gamma(\mathbf{A}_{i},\beta)\in B$ for every $\beta\in\operatorname{Perm}(\alpha)$. Then $\sigma$ is a perfect linear congruence $\sigma$ on $D^{(0)}_{(\mathbf{A}_{i},\alpha)}$.
Proof.
By Corollary 22 there must be a bridge $\delta$ from $\sigma$ to $\sigma$ such that $\widetilde{\delta}$ contains the relation $\operatorname{pr}_{(\mathbf{A}_{i},\alpha_{1}),(\mathbf{A}_{i},\alpha_{2})}R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ for some $\alpha_{1},\alpha_{2}\in\operatorname{Perm}(\alpha)$. Since $\operatorname{pr}_{(\mathbf{A}_{i},\alpha_{1}),(\mathbf{A}_{i},\alpha_{2})}R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ is linked, by Lemma 5 $\sigma$ is a perfect linear congruence, which completes the proof. ∎
Theorem 45.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$, $n$ is odd, $D^{(1)}$ is a 1-consistent symmetric reduction of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, $D^{(1)}\lll D^{(0)}$. Then one of the following conditions hold
  1. 1.
    $|D^{(1)}_{(\mathbf{A}_{i},\alpha)}|=1$ for all $(\mathbf{A}_{i},\alpha)$.
  2. 2.
    there exists a 1-consistent symmetric reduction $D^{(2)}$ for $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ such that $D^{(2)}\lll D^{(1)}$ and $D^{(2)}\neq D^{(1)}$.
  3. 3.
    there exists a perfect linear congruence $\sigma$ on some $D^{(0)}_{(\mathbf{A}_{i},\alpha)}$ such that
    1. (a)
      $D_{(\mathbf{A}_{i},\alpha)}^{(1)}\times D_{(\mathbf{A}_{i},\alpha)}^{(1)}\not\subseteq\sigma$
    2. (b)
      $D_{(\mathbf{A}_{i},\alpha)}^{(1)}\times D_{(\mathbf{A}_{i},\alpha)}^{(1)}\subseteq\sigma^{*}$
Proof.
If condition 1 holds then we are done. Otherwise, choose some $(\mathbf{A}_{i},\alpha)$ such that $|D_{(\mathbf{A}_{i},\alpha)}^{(1)}|>1$. By Lemma 13 there exists $B<_{\mathcal{T}(\sigma)}^{D_{(\mathbf{A}_{i},\alpha)}^{(0)}}D_{(\mathbf{A}_{i},\alpha)}^{(1)}$ for some $\mathcal{T}\in\{\mathcal{BA},\mathcal{C},\mathcal{D}\}$. If $\mathcal{T}$ can be chosen from $\{\mathcal{BA},\mathcal{C}\}$ then condition 2 follows from Theorem 49. Otherwise, we assume that $\mathcal{T}=\mathcal{D}$. Define a reduction $D^{(\top)}$ as follows. If $\beta\in\operatorname{Perm}(\alpha)$ put $D^{(\top)}_{(\mathbf{A}_{i},\beta)}=B$, for all other $(\mathbf{A}_{j},\beta)$ put $D^{(\top)}_{(\mathbf{A}_{j},\beta)}=D^{(1)}_{(\mathbf{A}_{j},\beta)}$. We consider two cases:
Case 1. $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(\top)}$ is not empty. Define a new reduction $D^{(2)}$ by $D^{(2)}_{(\mathbf{A}_{j},\beta)}=\operatorname{pr}_{(\mathbf{A}_{j},\beta)}(R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(\top)})$. By Corollary 18(r1) we have $D^{(2)}\lll^{D}D^{(1)}$. Since $D^{(\top)}$ and $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ are symmetric, $D^{(2)}$ is also symmetric. Thus, we satisfied condition 2.
Case 2. $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(\top)}$ is empty. Then by Lemma 50 condition 3 is satisfied. ∎
Theorem 46.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{n}$, $n$ is odd, there exists an $n$-ary term $\tau_{0}$ such that $\tau_{0}^{\mathbf{A}_{i}}$ is XY-symmetric for every $i$. Then there exists a 1-consistent symmetric reduction $D^{(\triangle)}\lll D^{(0)}$ of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and an $n$-ary term $\tau$ such that $\tau^{\mathbf{A}_{i}}$ is XY-symmetric and $D_{(\mathbf{A}_{i},\alpha)}^{(\triangle)}=\{\tau(\alpha)\}$ for every $i$ and $\alpha\in\operatorname{TwoTuples}(A_{i}^{n})$.
Proof.
Let $I$ be the set of all indices of the coordinates $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. That is, $I$ consists of pairs $(\mathbf{A}_{i},\alpha)$ and $|I|$ is the arity of $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$.
We build a sequence of symmetric reductions $D^{(s)}\lll^{D^{(0)}}D^{(s-1)}\lll^{D^{(0)}}\dots\lll^{D^{(0)}}D^{(1)}\lll D^{(0)}$ for $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. For every $j\in\{0,1,\dots,s\}$ and every $i\in I$ we define a congruence $\delta_{i}^{j}$ on $D_{i}^{(0)}$ such that the following conditions hold for every $j$:
  1. (1)
    $\delta_{(\mathbf{A}_{i},\alpha)}^{j}=\delta_{(\mathbf{A}_{i},\beta)}^{j}$ whenever $\beta\in\operatorname{Perm}(\alpha)$.
  2. (2)
    there exists a tuple $\gamma\in R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(j)}$ such that $(\gamma(\mathbf{A}_{i},\alpha),\gamma(\mathbf{A}_{i},\beta))\in\delta_{(\mathbf{A}_{i},\alpha)}^{j}$ for every $(\mathbf{A}_{i},\alpha)\in I$ and $\beta\in\operatorname{Perm}(\alpha)$.
  3. (3)
    $\delta_{i}^{j+1}\supseteq\delta_{i}^{j}$ for every $i\in I$.
  4. (4)
    if we make some congruences $\delta_{i}^{j}$ smaller then condition (2) or condition (3) will not hold.
Thus, for every $j$ we choose minimal congruences (but not smaller than the previous congruences) such that $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(j)}$ has a tuple whose corresponding elements are equivalent modulo these congruences. We start with $s=0$ and $\delta_{i}^{0}$ is the equality relation (0-congruence) for every $i\in I$.
We also need the following condition.
  1. (5)
    $\delta_{i}^{j}=\delta_{i}^{j+1}$ for every $i\in I$ or $D^{(j+1)}\leqslant_{\mathcal{T}}D^{(j)}$ for some $\mathcal{T}\in\{\mathcal{BA},\mathcal{C}\}$.
Suppose we already have $D^{(s)}$ and all $\delta_{i}^{s}$ are are defined. Let us show how to build $D^{(s+1)}$ and $\delta_{i}^{s+1}$. Using Lemma 13 we consider three cases:
Case 1. There exists $B<_{\mathcal{T}}D^{(s)}_{i}$ for some $i\in I$ and $\mathcal{T}\in\{\mathcal{BA},\mathcal{C}\}$. By Theorem 49 there exists a 1-consistent reduction $D^{(s+1)}\leqslant_{\mathcal{T}}D^{(s)}$. It remains to define $\delta_{i}^{s+1}$ for every $i\in I$. We choose them freely to satisfy the above condition (1)-(4).
Case 2. $|D^{(s)}_{i}|=1$ for every $i\in I$. Then we put $D^{(\triangle)}=D^{(s)}$ and choose an $n$-ary term $\tau$ such that $D_{(\mathbf{A}_{i},\alpha)}^{(\triangle)}=\{\tau(\alpha)\}$ for every $i$ and $\alpha\in\operatorname{TwoTuples}(A_{i}^{n})$. This completes the proof in this case.
Case 3. Otherwise, choose the maximal $\ell\in\{0,1,\dots,s\}$ such that$|D_{i}^{(s)}/\delta_{i}^{\ell}|>1$ for some $i\in I$. Since congruences $\delta_{i}^{0}$ are equalities and case 2 does not hold, such $\ell$ and $i$ always exist. By Lemma 13 choose $B^{\prime}<_{\mathcal{T}(\sigma^{\prime})}^{D_{i}^{(0)}/\delta_{i}^{\ell}}D_{i}^{(s)}/\delta_{i}^{\ell}$. If $\mathcal{T}\in\{\mathcal{BA},\mathcal{C}\}$ then by Corollary 16(t) and Lemma 20(t) we have $(\bigcup_{E\in B^{\prime}}E)\cap D_{i}^{(s)}<_{\mathcal{T}}D_{i}^{(s)}$, which means that it is case 1. Thus, $\mathcal{T}=\mathcal{D}$. By extending $\sigma^{\prime}$ to $D_{i}^{(0)}$ we get an irreducible congruence $\sigma\supseteq\delta_{i}^{\ell}$. Consider two subcases:
Subcase 3.1. $\ell=s$. Choose an equivalence class $B$ of $\sigma$ containing $\gamma(i)$ for $\gamma$ satisfying condition (2) for $j=s$. Define two reductions $D^{(\bot)}$ and $D^{(s+1)}$. Put $D_{j}^{(\bot)}=D_{j}^{(s)}$ for every $j\notin\operatorname{Perm}(i)$ and $D_{j}^{(\bot)}=B\cap D_{j}^{(s)}$ for $j\in\operatorname{Perm}(i)$. Put $D_{j}^{(s+1)}=\operatorname{pr}_{j}(R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(\bot)})$ for every $j\in I$. Notice that $D^{(\bot)}\leqslant_{\mathcal{D}}D^{(s)}$ and by Corollary 18(r1) we have $D^{(s+1)}\lll^{D^{(0)}}D^{(s)}$. Put $\delta_{i^{\prime}}^{s}=\delta_{i^{\prime}}^{s+1}$ for every $i^{\prime}\in I$. It is not hard to see that conditions (1)-(5) are satisfied.
Subcase 3.2. $\ell<s$. Choose some $B<_{\mathcal{D}(\sigma)}^{D_{i}^{(0)}}D_{i}^{(s)}$. Since $\ell$ was chosen maximal, $\sigma\not\supseteq\delta_{i}^{\ell+1}$. Since $\delta_{i}^{\ell+1}$ was chosen minimal, $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}^{(s)}$ has no tuples $\gamma$ such that $\gamma_{i^{\prime}}\in B$ for all $i^{\prime}\in\operatorname{Perm}(i)$ (otherwise we could replace $\delta_{i}^{\ell+1}$ by $\delta_{i}^{\ell+1}\cap\sigma$). By Lemma 50, $\sigma$ is a perfect linear congruence. Let $i=(\mathbf{A}_{k},\alpha)$. Then there exists $\zeta\leqslant\mathbf{A}_{k}\times\mathbf{A}_{k}\times\mathbf{Z}_{p}$ such that $\operatorname{pr}_{1,2}\zeta=\sigma^{*}$ and $(a_{1},a_{2},b)\in\zeta$ implies that $(a_{1},a_{2})\in\sigma\Leftrightarrow(b=0)$. Let us define a conjunctive formula $\Theta$ with variables $\{x_{j}\mid j\in I\}\cup\{z_{i^{\prime}}\mid i^{\prime}\in\operatorname{Perm}(i)\}$. For every $j\in I$ and every $j^{\prime}\in\operatorname{Perm}(j)$ we add the constraint $\delta_{j}^{\ell+1}(x_{j},x_{j^{\prime}})$, for every $i^{\prime}\in\operatorname{Perm}(i)$ we add the constraint $\zeta(x_{i},x_{i^{\prime}},z_{i^{\prime}})$, finally add $R_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ once with the corresponding variables. Since $\delta_{i}^{\ell+1}$ was chosen minimal and $\sigma\not\supseteq\delta_{i}^{\ell+1}$, $\Theta^{(\ell+1)}$ is not satisfiable if $z_{i^{\prime}}=0$ for every $i^{\prime}\in\operatorname{Perm}(i)$. Nevertheless, it is satisfied for some $z_{i^{\prime}}$. Also, since $\delta_{i}^{\ell}\subseteq\sigma$, the formula $\Theta^{(\ell)}$ has a solution with $z_{i^{\prime}}=0$ for every $i^{\prime}\in\operatorname{Perm}(i)$. By condition (5) $D^{(\ell+1)}\leqslant_{T}D^{(\ell)}$ for some $T\in\{\mathcal{BA},\mathcal{C}\}$. Thus, the projections $L_{1}$ and $L_{0}$ of the solution sets of $\Theta^{(\ell)}$ and $\Theta^{(\ell+1)}$, respectively, onto the $z$-variables are different (in the point $(0,0,\dots,0)$), which by Lemma 19 implies that $L_{0}<_{T}L_{1}\leqslant\mathbf{Z}_{p}^{|\operatorname{Perm}(i)|}$. This contradicts Lemma 29. ∎
4.4 Proof of Theorem 47 (Fixing an operation)
Lemma 51.
Suppose $\mathbf{A}=(A;w^{\mathbf{A}})$, where $w^{\mathbf{A}}$ is an $n$-ary idempotent operation, $0_{\mathbf{A}}$ is a perfect linear congruence witnessed by $\zeta\leqslant\mathbf{A}\times\mathbf{A}\times\mathbf{Z}_{p}$. Then
  1. (1)
    for every $(a,b)\in 0_{\mathbf{A}}^{*}$ there exists a unique $c$ such that $(a,b,c)\in\zeta$,
  2. (2)
    $n-1$ is divisible by $p$,
  3. (3)
    $w^{\mathbf{A}}(a,\dots,a,b,a,\dots,a)=b$ for every $(a,b)\in 0_{\mathbf{A}}^{*}$ and any position of $b$,
  4. (4)
    for every $a\in A$ and $c\in\mathbf{Z}_{p}$ there exists at most one $b\in A$ such that $(a,b,c\in\zeta)$,
  5. (5)
    for every $b\in A$ and $c\in\mathbf{Z}_{p}$ there exists at most one $a\in A$ such that $(a,b,c\in\zeta)$,
  6. (6)
    if $(a,b,d),(b,c,e)\in\zeta$ then $(a,c,d+e)\in\zeta$.
Proof.
Let us prove (1). Assuming the opposite we take $d_{1}$ and $d_{2}$ such that $(a,b,d_{1}),(a,b,d_{2})\in\zeta$. Applying $w$ to these tuples and using idempotency of $\mathbf{A}$ we derive that $(a,b,k_{1}\cdot d_{1}+k_{2}\cdot d_{2})\in\zeta$ for any $k_{1}$ and $k_{2}$ such that $k_{1}+k_{2}-1$ is divisible by $n-1$. Since $d_{1}\neq d_{2}$ we can choose $k_{1}$ and $k_{2}$ such that $k_{1}\cdot d_{1}+k_{2}\cdot d_{2}=0(mod\;\;p)$, which implies $(a,b,0)\in\zeta$ and contradicts the definition of $\zeta$.
Let us prove (2). Applying $w$ to $n$ tuples $(a,b,d)$ for some $d\neq 0$, we get a tuple $(a,b,n\cdot d)$. By (1) we obtain $n\cdot d=d(mod\;\;\;p)$ and $n-1$ is divisible by $p$.
Let us prove (3) using (2). Applying $w$ to $n-1$ tuples $(a,b,d)$ and one tuple $(b,b,0)$ we get a tuple $(w^{\mathbf{A}}(a,\dots,a,b,a,\dots,a),b,0)$. Hence $w^{\mathbf{A}}(a,\dots,a,b,a,\dots,a)=b$.
Let us prove (4). Assume that $(a,b_{1},d),(a,b_{2},d)\in\zeta$. By properties (2) and (3), applying $w$ to $n-2$ tuples $(a,b_{2},d)$, one tuple $(a,b_{1},d)$ and one tuple $(b_{2},b_{2},0)$ we obtain $(b_{2},b_{1},0)\in\zeta$, which means $b_{1}=b_{2}$.
Property (5) is proved in the same way as (4).
To prove (6) apply $w$ to the tuples $(a,b,d),(b,c,e)$ and $(n-2)$ tuples $(b,b,0)$. By (2) and (3) we have $(a,c,d+e)\in\zeta$, which completes the proof. ∎
Using claim (1) in Lemma 51 we can define a binary operation $\zeta:A\times A\to\mathbf{Z}_{p}$ by $\zeta(x_{1},x_{2})=z\leftrightarrow(x_{1},x_{2},z)\in\zeta$. Then property (6) means $\zeta(a,c)=\zeta(a,b)+\zeta(b,c)$ for any $(a,b),(b,c)\in 0_{\mathbf{A}}^{*}$.
Theorem 52.
Suppose $\mathbf{A}\in\mathcal{V}_{n}$, $0_{\mathbf{A}}$ is a perfect linear congruence. Then there exists an algebra $\mathbf{C}\in(\mathbf{A}/0_{\mathbf{A}}^{*}\boxtimes\mathbf{Z}_{p})\cap\mathcal{V}_{n}$ such that there exists an injective homomorphism $h\colon\mathbf{A}\to\mathbf{C}$.
Proof.
Suppose $0_{\mathbf{A}}$ is witnessed by $\zeta\leqslant\mathbf{A}\times\mathbf{A}\times\mathbf{Z}_{p}$. Then we can use all the properties of $\zeta$ proved in Lemma 51.
Let us define an algebra $\mathbf{C}=(A/0_{\mathbf{A}}^{*}\times\mathbf{Z}_{p};w^{\mathbf{C}})$. Let $\phi:A/0_{\mathbf{A}}^{*}\to A$ be an injection such that $\phi(\theta)\in\theta$ for every equivalence class $\theta$. Thus, we just chose a representative from every equivalence class of $0_{\mathbf{A}}^{*}$. Put
$(w^{\mathbf{C}})^{(1)}(x_{1},\dots,x_{n})=w^{\mathbf{A}/0_{\mathbf{A}}^{*}}(x_{1}^{(1)},\dots,x_{n}^{(1)}),$
$(w^{\mathbf{C}})^{(2)}(x_{1},\dots,x_{n})=\zeta\left(w^{\mathbf{A}}(\phi(x_{1}^{(1)}),\dots,\phi(x_{n}^{(1)})),\phi(w^{\mathbf{A}/0_{\mathbf{A}}^{*}}(x_{1}^{(1)},\dots,x_{n}^{(1)}))\right)+x_{1}^{(2)}+\dots+x_{n}^{(2)}.$
Let us define an injective homomorphism $h\colon\mathbf{A}\to\mathbf{C}$ by $h(a)=(a/0_{\mathbf{A}}^{*},\zeta(a,\phi(a/0_{\mathbf{A}}^{*})))$. By Lemma 51(5) $h$ is injective. Let us prove that $h$ is a homomorphism. It follows immediately from the definition of $h$ and $w^{\mathbf{C}}$ that
$(h(w^{\mathbf{A}}(x_{1},\dots,x_{n})))^{(1)}=w^{\mathbf{A}/0_{\mathbf{A}}^{*}}(x_{1}/0_{\mathbf{A}}^{*},\dots,x_{n}/0_{\mathbf{A}}^{*})=(w^{\mathbf{C}}(h(x_{1}),\dots,h(x_{n})))^{(1)}.$
Applying $w^{\mathbf{A}}$ to the tuples $(x_{i},\phi(x_{i}/0_{\mathbf{A}}^{*}),h^{(2)}(x_{i}))\in\zeta$ for $i=1,2,\dots,n$ we obtain
$\zeta(w^{\mathbf{A}}(x_{1},\dots,x_{n}),w^{\mathbf{A}}(\phi(x_{1}/0_{\mathbf{A}}^{*}),\dots,\phi(x_{n}/0_{\mathbf{A}}^{*})))=h^{(2)}(x_{1})+\dots+h^{(2)}(x_{n}).$
Using Lemma 51(6) we derive
$)\displaystyle h^{(2)}(w^{\mathbf{A}}(x_{1},\dots,x_{n})$)=) =
$\zeta(w^{\mathbf{A}}(x_{1},\dots,$ $x_{n}),\phi(w^{\mathbf{A}}(x_{1},\dots,x_{n})/0_{\mathbf{A}}^{*}))=$
$\zeta(w^{\mathbf{A}}$ $(x_{1},\dots,x_{n}),w^{\mathbf{A}}(\phi(x_{1}/0_{\mathbf{A}}^{*}),\dots,\phi(x_{n}/0_{\mathbf{A}}^{*})))+$
$\;\;\;\;\;\;\;\zeta(w^{\mathbf{A}}(\phi(x_{1}/0_{\mathbf{A}}^{*}),\dots,\phi(x_{n}/0_{\mathbf{A}}^{*})),\phi(w^{\mathbf{A}}(x_{1},\dots,x_{n})/0_{\mathbf{A}}^{*}))=$
$h^{(2)}($ $x_{1})+\dots+h^{(2)}(x_{n})+$
$\zeta(w^{\mathbf{A}}(\phi(h^{(1)}(x_{1})),\dots,\phi(h^{(1)}(x_{n}))),\phi(w^{\mathbf{A}/0_{\mathbf{A}}^{*}}(h^{(1)}(x_{1}),\dots,h^{(1)}(x_{1}))))=$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(w^{\mathbf{C}})^{(2)}(h(x_{1}),\dots,h(x_{n}))$
Hence, $h(w^{\mathbf{A}}(x_{1},\dots,x_{n}))=w^{\mathbf{C}}(h(x_{1}),\dots,h(x_{n}))$ and $h$ is a homomorphism.
It follows from the definition that $\mathbf{C}\in\mathbf{A}/0_{\mathbf{A}}^{*}\boxtimes\mathbf{Z}_{p}$. It remains to show that $w^{\mathbf{C}}\in\mathcal{V}_{n}$. Since $w^{\mathbf{A}}$ is a WNU and the addition $x_{1}^{(2)}+\dots+x_{n}^{(2)}$ in the definition of $(w^{\mathbf{C}})^{(2)}$ is symmetric, $w^{\mathbf{C}}$ is also a WNU.
Since $h$ is an isomorphism from $A$ to $h(A)$, $w^{\mathbf{C}}$ is a special WNU on $h(A)$. Notice that $h(\phi(x/0_{\mathbf{A}}^{*}))=(x/0_{\mathbf{A}}^{*},0)$, which means that $(b,0)\in h(A)$ for any $b\in A/0_{\mathbf{A}}^{*}$. For $x\in C$ put $x^{\prime}=(x^{(1)},0)$.
Let us show that $w^{\mathbf{C}}$ is idempotent. Since $w^{\mathbf{C}}$ is idempotent on $h(A)$, we have $w^{\mathbf{C}}(x^{\prime},x^{\prime},\dots,x^{\prime})=x^{\prime}$ and $(w^{\mathbf{C}})^{(2)}(x^{\prime},x^{\prime},\dots,x^{\prime})=0$. Hence, $(w^{\mathbf{C}})^{(2)}(x,x,\dots,x)=(w^{\mathbf{C}})^{(2)}(x^{\prime},x^{\prime},\dots,x^{\prime})+n\cdot x^{(2)}=x^{(2)}$ and $w^{\mathbf{C}}(x,x,\dots,x)=x$.
Let us show that $w^{\mathbf{C}}$ is special and $w^{\mathbf{C}}(x,x,\dots,x,w^{\mathbf{C}}(x,x,\dots,x,y))=w^{\mathbf{C}}(x,x,\dots,x,y)$. For $(w^{\mathbf{C}})^{(1)}$ it follows from the fact that $w^{\mathbf{A}/0_{\mathbf{A}}^{*}}$ is special. Since $w^{\mathbf{C}}$ is special on $f(A)$, we have
$w^{\mathbf{C}}(x^{\prime},\dots,x^{\prime},w(x^{\prime},\dots,x^{\prime},y^{\prime}))=w^{\mathbf{C}}(x^{\prime},\dots,x^{\prime},y^{\prime}).$
Looking at the second components in the above equality we obtain $(w^{\mathbf{C}})^{(2)}(x^{\prime},x^{\prime},\dots,x^{\prime},y^{\prime})=0$. Therefore, using the definition of $w^{\mathbf{C}}$ we derive
$(w^{\mathbf{C}})^{(2)}(x,x,\dots,x,w^{\mathbf{C}}(x,x,\dots,$ $x,y))=$
$2\cdot(w^{\mathbf{C}})^{(2)}(x^{\prime},x^{\prime},\dots,x^{\prime},y^{\prime})$ $+2\cdot(n-1)\cdot x^{(2)}+y^{(2)}=$
$\;\;(w^{\mathbf{C}})^{(2)}(x^{\prime},x^{\prime},\dots,x^{\prime},y^{\prime})+(n-1)\cdot x^{(2)}+y^{(2)}=$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(w^{\mathbf{C}})^{(2)}(x,x,\dots,x,y)$
For a set $B$ and prime $p$ by $\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$ we denote the set of all operations $f$ on $B\times Z_{p}$ such that $f^{(1)}(x_{1},\dots,x_{n})$ depends only on $x_{1}^{(1)},\dots,x_{n}^{(1)}$ and $f^{(2)}(x_{1},\dots,x_{n})=f^{(2,1)}(x_{1}^{(1)},\dots,x_{n}^{(1)})+a_{1}x_{1}^{(2)}+\dots+a_{n}x_{n}^{(2)}$ for some $f^{(2,1)}:B^{n}\to Z_{p}$ and $a_{1},\dots,a_{n}\in Z_{p}$. The subfunction $a_{1}x_{1}^{(2)}+\dots+a_{n}x_{n}^{(2)}$ we sometimes denote by $f^{(2,2)}$. Sometimes we write $f^{(1)}(x_{1}^{(1)},\dots,x_{n}^{(1)})$ instead of $f^{(1)}(x_{1},\dots,x_{n})$ to point out that $f^{(1)}$ depends only on first components.
For an operation $f$ of arity $n$ we define operations $t_{\ell}^{f}$ for $\ell\in\{n,n-1,\dots,1,0\}$ as follows. Put $t_{n}^{f}(x_{1},\dots,x_{n},y_{1},\dots,y_{n})=f(x_{1},\dots,x_{n})$. For $\ell\in\{n-1,\dots,1,0\}$ put
$t_{\ell}^{f}(x_{1},\dots,x_{\ell},y_{1},\dots,y_{n})=f($ $t_{\ell+1}(x_{1},\dots,x_{\ell},y_{1},y_{1},\dots,y_{n}),$
$t_{\ell+1}(x_{1},\dots,x_{\ell},y_{2},y_{1},\dots,y_{n}),\dots,t_{\ell+1}(x_{1},\dots,x_{\ell},y_{n},y_{1},\dots,y_{n}))$
Lemma 53.
Suppose $f$ is a $k$-WNU. Then $t_{0}^{f}$ is a $k$-WNU.
Proof.
Suppose tuples $(a_{1},\dots,a_{n}),(b_{1},\dots,b_{n})\in\{c,d\}^{n}$ contain exactly $k$ elements $d$. We need to show that $t_{0}^{f}(a_{1},\dots,a_{n})=t_{0}^{f}(b_{1},\dots,b_{n})$. We prove by induction on $\ell$ starting with $\ell=n$ to $\ell=0$ that $t_{\ell}^{f}(x_{1},\dots,x_{\ell},a_{1},\dots,a_{n})=t_{e\\ }^{f}(x_{1},\dots,x_{\ell},b_{1},\dots,b_{n})$ for all $x_{1},\dots,x_{\ell}$. For $\ell=n$ it is obvious. From the inductive assumption on $t_{\ell+1}^{f}$ we conclude that
$a_{i}=b_{j}\Longrightarrow t_{\ell+1}(x_{1},\dots,x_{\ell},a_{i},a_{1},\dots,a_{n})=t_{\ell+1}(x_{1},\dots,x_{\ell},b_{j},b_{1},\dots,b_{n}).$
Then the inductive step follows from the definition of $t_{\ell}^{f}$ and the fact that $f$ is a $k$-WNU. ∎
Note, that if we write the term defining $t_{0}^{f}(y_{1},\dots,y_{n})$ then for every $(i_{1},\dots,i_{n})\in[n]^{n}$ there exists exactly one internal formula $f(y_{i_{1}},\dots,y_{i_{n}})$. For two tuples $(i_{1},i_{2},\dots,i_{n}),(j_{1},j_{2},\dots,j_{n})\in[n]^{n}$ by $t_{0,(i_{1},i_{2},\dots,i_{n})}^{f,(j_{1},j_{2},\dots,j_{n})}$ we denote the operation defined by the same term as $t_{0}^{f}$ but with $f(y_{i_{1}},\dots,y_{i_{n}})$ replaced by $f(y_{j_{1}},\dots,y_{j_{n}})$.
For a tuple $\alpha$ and an element $b$ by $N_{b}(\alpha)$ we denote the number of elements in $\alpha$ that are equal to $b$. By $T_{a,b}^{n,k}$ we denote the set of all tuples $\gamma\in\{a,b\}^{n}$ such that $N_{b}(\gamma)=k$ and $\gamma(1)=a$.
For two tuples $\alpha,\beta\in T_{a,b}^{n,k}$ we define a tuple $\xi(\alpha,\beta)=(i_{1},\dots,i_{n})\in[n]^{n}$ as follows. Let $j_{1}<\dots<j_{k}$ and $s_{1}<\dots<s_{k}$ be the lists of positions of $b$ in $\alpha$ and $\beta$ respectively. Put $i_{\ell}=1$ if $\beta(\ell)=a$, and $i_{\ell}=j_{m}$ if $\ell=s_{m}$. For example, $\xi((a,a,b,a,b,a,b,b,a),(a,b,a,b,b,a,a,a,b))=(1,3,1,5,7,1,1,1,8)$. Notice that for any $\alpha,\beta,\gamma\in T_{a,b}^{n,k}$ the tuple $\xi(\alpha,\beta)$ is a permutation of the tuple $\xi(\alpha,\gamma)$.
Lemma 54.
Suppose $f$ is a special idempotent WNU operation of arity $n$, $f\in\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$. Then $f^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$ and $p$ divides $n-1$.
Proof.
By the definition of $\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$ we have $f^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=a_{1}x_{1}+\dots+a_{n}x_{n}$. Since $f$ is a WNU, we have $a_{1}=a_{2}=\dots=a_{n}$. Since $f$ is idempotent, $n\cdot a_{1}=1\mod p$. Since $f$ is special we have $f(y,y,\dots,y,f(y,y,\dots,y,x))=f(y,y,\dots,y,x)$, which implies $a_{1}\cdot a_{1}=a_{1}$ and $a_{1}=1$. ∎
The next lemma follows immediately from the definition of $t_{0}^{f}$ and $t_{0,(i_{1},i_{2},\dots,i_{n})}^{f,(j_{1},j_{2},\dots,j_{n})}$.
Lemma 55.
Suppose $f\in\operatorname{Pol}(\sigma_{B\times\mathbb{Z}})$, $f^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$, where $p$ divides $n-1$, and $f^{(1)}(x_{i_{1}},\dots,x_{i_{n}})=f^{(1)}(x_{j_{1}},\dots,x_{j_{n}})$. Then $(t_{0,(i_{1},i_{2},\dots,i_{n})}^{f,(j_{1},j_{2},\dots,j_{n})})^{(1)}=(t_{0}^{f})^{(1)}$ and $(t_{0,(i_{1},i_{2},\dots,i_{n})}^{f,(j_{1},j_{2},\dots,j_{n})})^{(2)}(x_{1},\dots,x_{n})=(t_{0}^{f})^{(2)}(x_{1},\dots,x_{n})-f^{(2)}(x_{i_{1}},\dots,x_{i_{n}})+f^{(2)}(x_{j_{1}},\dots,x_{j_{n}})$.
Lemma 56.
Suppose $f\in\operatorname{Pol}(\sigma_{B\times\mathbb{Z}})$, $f^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$, where $p$ divides $n-1$. Then $(t_{0}^{f})^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$.
Proof.
$(t_{0}^{f})^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=$
$\sum\limits_{i_{1},\dots,i_{n}\in[n]}$ $(x_{i_{1}}^{(2)}+\dots+x_{i_{n}}^{(2)})=n^{n-1}\cdot(x_{i_{1}}^{(2)}+\dots+x_{i_{n}}^{(2)})=x_{i_{1}}^{(2)}+\dots+x_{i_{n}}^{(2)}.$
Lemma 57.
Suppose $f\in\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$, $f^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$, $m\in\mathbb{N}$. Then there exists $g\in\operatorname{Clo}(f)$ such that $g^{(2,2)}(x_{1},\dots,x_{2m+1})=x_{1}^{(2)}-x_{2}^{(2)}+x_{3}^{(2)}-\dots-x_{2m}^{(2)}+x_{2m+1}^{(2)}$.
Proof.
Consider a term defining $x_{1}-x_{2}+x_{3}-\dots+x_{2m+1}$ from $x_{1}+x_{2}+\dots+x_{n}$ in $\mathbf{Z}_{p}$. The same term defines the required $g$ from $f$. ∎
To simplify notations we use operations coming from Lemma 57 as follows. Whenever we write $\bigoplus\limits_{i=1}^{m}(a_{i}\cdot(x_{i}\ominus x_{i}^{\prime}))\oplus x_{m+1}$ for $a_{1},\dots,a_{m}\in\mathbf{Z}_{p}$, we mean
$g(\underbrace{x_{1},x_{1}^{\prime},\dots,x_{1},x_{1}^{\prime}}_{2a_{1}},\underbrace{x_{2},x_{2}^{\prime},\dots,x_{2},x_{2}^{\prime}}_{2a_{2}},\dots,\underbrace{x_{m},x_{m}^{\prime},\dots,x_{m},x_{m}^{\prime}}_{2a_{m}},x_{m+1}),$
where $g$ is a $(2\cdot\sum_{i=1}^{m}a_{i}+1)$-ary operation coming from Lemma 57. Notice that we use this notation if the only important part of the obtained operation $f$ is $f^{(2,2)}$.
We say that an operation $f$ is symmetric on a tuple $(a_{1},\dots,a_{n})$ if $f(a_{1},\dots,a_{n})=f(a_{\sigma(1)},\dots,a_{\sigma(n)})$ for any permutation $\sigma\colon[n]\to[n]$. We say that an operation $f$ is weakly symmetric on a tuple $(a_{1},\dots,a_{n})$ if $f(a_{1},\dots,a_{n})=f(a_{\sigma(1)},\dots,a_{\sigma(n)})$ for any permutation $\sigma\colon[n]\to[n]$ such that $\sigma(1)=1$.
Suppose $P\subseteq\{(c,d)\mid c,d\in B,c\neq d\}$ and $a,b\in B$. We say that an $n$-ary operation $f\in\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$ is $(P,a,b,k)$-symmetric if
  1. (1)
    $p$ divides $n-1$,
  2. (2)
    $f^{(1)}$ is XY-symmetric,
  3. (3)
    $f^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$,
  4. (4)
    $f^{(2,1)}$ is weakly symmetric on all tuples $\alpha\in\{c,d\}^{n}$ such that $\alpha(1)=c$ and $(c,d)\in P$,
  5. (5)
    $f^{(2,1)}$ is weakly symmetric on all tuples $\alpha\in\{a,b\}^{n}$ such that $\alpha(1)=a$ and $N_{b}(\alpha)\leqslant k$.
An operation is called $P$-symmetric if it satisfies only items (1)-(4).
Lemma 58.
Suppose $P\subseteq\{(c,d)\mid c,d\in B,c\neq d\}$, $10\leqslant k<n-$ 1, $a,b\in B$, and $f\in\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$ is a $(P,a,b,k)$-symmetric operation. Then there exists a $(P,a,b,k+1)$-symmetric operation $g\in\operatorname{Clo}(f)$ of arity $n$.
Proof.
The function $f$ satisfies all the properties we require for $g$ except for the property (5) for $N_{b}(\alpha)=k+1$. If $f$ is also weekly symmetric on such tuples then we just take $g=f$. Otherwise, consider two tuples $\alpha,\beta\in\{a,b\}^{n}$ such that $\alpha(1)=\beta(1)=a$, $N_{b}(\alpha)=N_{b}(\beta)=k+1$, and $f^{(2,1)}(\alpha)\neq f^{(2,1)}(\beta)$.
Define a new operation
$,\displaystyle g(y_{1},$$\dots,y_{n}):=$
$\bigoplus\limits_{\gamma\in T_{a,b}^{n,k+1}}\frac{f^{(2,1)}(\alpha)-f^{(2,1)}(\gamma)}{f^{(2,1)}(\beta)-f^{(2,1)}(\alpha)}\cdot\left(t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(y_{1},\dots,y_{n})\ominus t_{0}^{f}(y_{1},\dots,y_{n})\right)\oplus f(y_{1},\dots,y_{n})$
Let us show that $g$ satisfies the required properties (2)-(5). Since $f^{(1)}$ is XY-symmetric, by Lemmas 53 and 55, $g^{(1)}$ is also XY-symmetric. Property (3) follows from Lemma 56. Let us prove property (4). Consider two tuples $\delta,\delta^{\prime}\in\{c,d\}^{n}$ such that $\delta(1)=\delta^{\prime}(1)=c$, $N_{d}(\delta)=N_{d}(\delta^{\prime})$, and $(c,d)\in P$. We need to prove that $g^{(2,1)}(\delta)=g^{(2,1)}(\delta^{\prime})$. Since $(c,d)\in P$, by Lemma 55 for every $\gamma\in T_{a,b}^{n,k+1}$ we have $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta)=t_{0}^{f}(\delta)$ and $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta^{\prime})=t_{0}^{f}(\delta^{\prime})$. Hence $g^{(2,1)}(\delta)-g^{(2,1)}(\delta^{\prime})=f^{(2,1)}(\delta)-f^{(2,1)}(\delta^{\prime})=0$.
Let us prove property (5). Consider two tuples $\delta,\delta^{\prime}\in T_{a,b}^{n,k^{\prime}}$, where $k^{\prime}\leqslant k+1$. We need to prove that $g^{(2,1)}(\delta)=g^{(2,1)}(\delta^{\prime})$. if $k^{\prime}\leqslant k$ then it follows from the property (5) for $f$ that $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta)=t_{0}^{f}(\delta)$ and $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta^{\prime})=t_{0}^{f}(\delta^{\prime})$, and therefore $g^{(2,1)}(\delta)-g^{(2,1)}(\delta^{\prime})=f^{(2,1)}(\delta)-f^{(2,1)}(\delta^{\prime})=0$.
Assume that $k^{\prime}=k+1$. Then $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta)=t_{0}^{f}(\delta)$ whenever $\gamma\neq\delta$. Therefore, in the definition of $g$ we only care about elements of the $\bigoplus$ corresponding to $\gamma=\delta$. Hence, by Lemma 55 we have
$g^{(2,1)}(\delta)-g^{(2,1)}(\delta^{\prime})=$
$\frac{f^{(2,1)}(\alpha)-f^{(2,1)}(\delta)}{f^{(2,1)}(\beta)-f^{(2,1)}(\alpha)}\cdot((t_{0,\xi(\delta,\alpha)}^{f,\xi(\delta,\beta)})^{(2,1)}(\delta)-(t_{0}^{f})^{(2,1)}(\delta))-$
$\frac{f^{(2,1)}(\alpha)-f^{(2,1)}(\delta^{\prime})}{f^{(2,1)}(\beta)-f^{(2,1)}(\alpha)}\cdot((t_{0,\xi(\delta^{\prime},\alpha)}^{f,\xi(\delta^{\prime},\beta)})^{(2,1)}(\delta^{\prime})-(t_{0}^{f})^{(2,1)}(\delta^{\prime}))+f^{(2,1)}(\delta)-f^{(2,1)}(\delta^{\prime})=$
$\frac{f^{(2,1)}(\alpha)-f^{(2,1)}(\delta)}{f^{(2,1)}(\beta)-f^{(2,1)}(\alpha)}\cdot(f^{(2,1)}(\beta)-f^{(2,1)}(\alpha))-$
$\frac{f^{(2,1)}(\alpha)-f^{(2,1)}(\delta^{\prime})}{f^{(2,1)}(\beta)-f^{(2,1)}(\alpha)}\cdot(f^{(2,1)}(\beta)-f^{(2,1)}(\alpha))+f^{(2,1)}(\delta)-f^{(2,1)}(\delta^{\prime})=$
$f^{(2,1)}(\alpha)-f^{(2,1)}(\delta)-f^{(2,1)}(\alpha)+f^{(2,1)}(\delta^{\prime})+f^{(2,1)}(\delta)-f^{(2,1)}(\delta^{\prime})=0$
Corollary 59.
Suppose $g\in\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$, $g^{(1)}$ is an $n$-ary XY-symmetric operation, $p$ divides $n-1$, $g^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$. Then $\operatorname{Clo}(g)$ contains an $n$-ary operation $h$ such that $h^{(1)}$ is XY-symmetric, $h^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$, and $h$ is weakly symmetric on all tuples $\alpha$ having two different elements.
Proof.
Notice that $g$ is $(\varnothing,a,b,0)$-symmetric for any $a,b\in B$. Using Lemma 58 we can increase the set $P$ of tuples on which $g^{(2,1)}$ is weakly symmetric until we get a required operation.
Formally, we prove as follows. Consider operations $f\in\operatorname{Clo}(g)$ that is $P$-symmetric for an inclusion maximal set $P$. If $P$ contains all pairs then we found the required operation. Otherwise, choose $(a,b)\notin P$. Choose a maximal $k$ such that there exists a $(P,a,b,k)$-symmeric operation $f\in\operatorname{Clo}(g)$.
Then applying Lemma 58 we can always increase $k$ if $k<n-1$, which contradicts the maximality of $k$. Notice that if $k=n-2$ then Lemma 58 guarantees that the pair $(a,b)$ can be included into $P$, which contradicts our assumption about the maximality of $P$. ∎
Theorem 60.
Suppose $\mathbf{A}\in(\mathbf{B}\boxtimes\mathbb{Z}_{p})\cap\mathcal{V}_{n}$, $w^{\mathbf{B}}$ is XY-symmetric. Then there exists a term $t$ of arity $n$ such that $t^{\mathbf{A}}$ is XY-symmetric.
Proof.
By the definition $w^{\mathbf{A}}\in\mathcal{C}_{B\boxtimes\mathbf{Z}_{p}}$. By Lemma 54, $(w^{\mathbf{A}})^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}$ and $p$ divides $n-1$. By Corollary 59, there exists an $n$-ary term $\tau$ such that $\tau^{\mathbf{B}}$ is XY-symmetric and $\tau^{\mathbf{A}}$ is weakly symmetric on all tuples containing two different elements. By $t$ define the term
$w(\tau(x_{1},\dots,x_{n}),\tau(x_{2},\dots,x_{n},x_{1}),\tau(x_{3},\dots,x_{n},x_{1},x_{2}),\dots,\tau(x_{n},x_{1},\dots,x_{n-1})).$
Let us show that $t$ is XY-symmetric.
Since $\tau^{\mathbf{A}}$ is weakly symmetric on any $\gamma\in\{a,b\}^{n}$, $\tau^{\mathbf{A}}(\gamma)$ depends only on $\gamma(1)$ and $N_{b}(\gamma)$. Notice that $(t^{\mathbf{A}})^{(1)}=t^{\mathbf{B}}$ is XY-symmetric and
$(t^{\mathbf{A}})^{(2,2)}(x_{1}^{(2)},\dots,x_{n}^{(2)})=n\cdot(x_{1}^{(2)}+\dots+x_{n}^{(2)})=x_{1}^{(2)}+\dots+x_{n}^{(2)}.$
Suppose $\alpha\in\{a,b\}^{n}$ for some $a,b\in B$. Then
$(t^{\mathbf{A}})^{(2,1)}(\alpha)=N_{a}(\alpha)\cdot(\tau^{\mathbf{A}})^{(2,1)}(a,\dots,a,\underbrace{b,\dots,b}_{N_{b}(\alpha)})+N_{b}(\alpha)\cdot(\tau^{\mathbf{A}})^{(2,1)}$ $(\underbrace{b,\dots,b}_{N_{b}(\alpha)},a,\dots,a)+$
$(w^{\mathbf{A}})^{(2,1)}(\tau^{\mathbf{B}}(\alpha),\dots,\tau^{\mathbf{B}}(\alpha)).$
Hence $(t^{\mathbf{A}})^{(2,1)}(\alpha)$ depends only on $N_{b}(\alpha)$, which means that $(t^{\mathbf{A}})^{(2,1)}$ is symmetric on $\alpha$. Therefore, $t^{\mathbf{A}}$ is XY-symmetric. ∎
Theorem 47.
Suppose $\mathbf{A},\mathbf{B}\in\mathcal{V}_{n}$, $0_{\mathbf{A}}$ is a perfect linear congruence, $\mathbf{A}/{0_{\mathbf{A}}}^{*}\times\mathbf{B}$ has an XY-symmetric term operation of arity $n$. Then $\mathbf{A}\times\mathbf{B}$ has an XY-symmetric term operation.
Proof.
By Theorem 52, there exists an algebra $\mathbf{C}\in\mathbf{(}(\mathbf{A}/0_{\mathbf{A}}^{*})\boxtimes\mathbf{Z}_{p})\cap\mathcal{V}_{n}$ such that there exists an injective homomorphism $h\colon\mathbf{A}\to\mathbf{C}$. Let $\delta$ be the canonical congruence on $\mathbf{C}$ such that $\mathbf{C}/\delta\cong\mathbf{A}/0_{\mathbf{A}}^{*}$. Put $\mathbf{D}=\mathbf{C}\times\mathbf{B}$ and extend the congruence $\delta$ on $\mathbf{D}$. Then $\mathbf{D}\in((\mathbf{D}/\delta)\boxtimes\mathbf{Z}_{p})\cap\mathcal{V}_{n}$. By Theorem 60 there exists a term $\tau$ such that $\tau^{\mathbf{D}}$ is XY-symmetric. Hence, both $\tau^{\mathbf{C}}$ and $\tau^{\mathbf{B}}$ are XY-symmetric. Since $\mathbf{A}$ is isomorphic to a subalgebra of $\mathbf{C}$, $\tau^{\mathbf{A}}$ is also XY-symmetric. ∎
5 Proof of the properties of strong subuniverses
In the section we prove all the statements formulated in Section 2 and this is the most technical part of the paper. We start with a few additional notations, then we formulate necessary properties of binary absorbing and central subuniverses that are mostly taken from [34]. In Subsection 5.3 we show that intersection of strong subalgebras behaves well, which is one of the main properties of strong subalgebras and definitely the most difficult to prove. In the next subsection we show the properties of a bridge connecting linear and PC congruences. For instance, there we prove that Linear and PC congruences can never be connected by a bridge and bridges for the PC congruences are trivial. In the next subsections we show that we should intersect strong subuniverses of the same type to obtain an empty set, and prove that factorization of strong subalgebras modulo a congruence behaves well. Finally, in Subsection 5.7 we prove most of the statements formulated in Section 2.3.
5.1 Additional definitions
In this section we call a relation symmetric if any permutation of its variables gives the same relation. For a relation $R\leqslant\mathbf{A}_{1}\times\dots\mathbf{A}_{n}$ by $\operatorname{LeftLinked}(R)$ we denote the minimal equivalence relation on $\operatorname{pr}_{1}(R)$ such that $(a_{1},a_{2},\dots,a_{n}),(b_{1},a_{2},\dots,a_{n})\in R$ implies $(a_{1},b_{1})\in\operatorname{LeftLinked}(R)$. Similarly, $\operatorname{RightLinked}(R)$ is the minimal equivalence relation on $\operatorname{pr}_{n}(R)$ such that $(a_{1},\dots,a_{n-1},a_{n}),(a_{1},\dots,a_{n-1},b_{n})\in R$ implies $(a_{n},b_{n})\in\operatorname{RightLinked}(R)$.
A relation $R\lneq A\times B$ is called central if there exists $b\in B$ such that $A\times\{b\}\subseteq R$.
5.2 Subuniverses of types $\mathcal{BA},\mathcal{C},\mathcal{S}$
Here we formulate some properties of strong subuniverse that we will use later.
Lemma 61 ([34], Lemma 6.25).
Suppose $B<_{T_{1}}A$, $C<_{T_{2}}A$, $B\cap C=\varnothing$, $T_{1},T_{2}\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$. Then $T_{1}=T_{2}\in\{\mathcal{BA},\mathcal{C}\}$.
Lemma 62 ([3], Lemma 2.9, [34], Lemma 6.1, Theorem 6.9).
Suppose $R\leqslant A^{n}$ is defined by a pp-formula $\Phi$ containing a relation $S$ and $\Phi^{\prime}$ is obtained by $\Phi$ by replacement of each appearance of $S$ by $S^{\prime}<_{T}S$, where $T\in\{\mathcal{BA},\mathcal{C}\}$. Then $\Phi^{\prime}$ defines a relation $R^{\prime}$ such that $R^{\prime}\leqslant_{T}R$.
The above lemma implies an easier claim.
Lemma 63.
Suppose $B\leqslant_{T}\mathbf{A}$ and $C\leqslant\mathbf{A}$, where $T\in\{\mathcal{BA},\mathcal{C}\}$. Then $B\cap C\leqslant_{T}C$.
Lemma 64.
Suppose $C\lll^{A}B\lll A$ and $D<_{\mathcal{BA},\mathcal{C}}B$. Then $C\cap D\neq\varnothing$.
Proof.
Assume the converse. Consider a minimal $C^{\prime\prime}$ such that $C\lll^{A}C^{\prime}<_{T(\sigma)}^{A}C^{\prime\prime}\lll^{A}B$ and $C^{\prime\prime}\cap D\neq\varnothing$. By Lemma 63 $C^{\prime\prime}\cap D<_{\mathcal{BA},\mathcal{C}}C^{\prime\prime}$, which by Lemma 61 implies that $T=\mathcal{D}$. By Lemma 65 we have $(C^{\prime\prime}\cap D)/\sigma<_{\mathcal{BA}}C^{\prime\prime}/\sigma$, which contradicts the definition of a dividing subuniverse. ∎
Lemma 65.
Suppose $B\leqslant_{T}A$ and $\sigma$ is a congruence on $A$, where $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$. Then $B/\sigma\leqslant_{T}A/\sigma$.
Proof.
For $T=\mathcal{BA}$ it is straightforward, for $T=\mathcal{C}$ see Lemma 6.8 in [34], for $T=\mathcal{S}$ it is just a combination of the results for $\mathcal{BA}$ and $\mathcal{C}$. ∎
Lemma 66 ([3], Proposition 2.14, [34], Lemma 3.2).
Suppose $B\leqslant\mathbf{A}$, $n\geqslant 2$. Then $B$ is an absorbing subuniverse with an operation of arity $n$ if and only if there does not exist $S\leqslant A^{n}$ such that $S\cap B^{n}=\varnothing$ and $S\cap(B^{i-1}\times A\times B^{n-i})\neq\varnothing$ for every $i\in\{1,2,\dots,n\}$.
Lemma 67.
Suppose $B<_{T}A^{n}$, where $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$. Then there exists $C<_{T}A$.
Proof.
For $T\in\{\mathcal{BA},\mathcal{C}\}$ see Lemma 6.24 in [34]. For $T=\mathcal{S}$ just repeat the same proof word to word replacing $\mathcal{BA}$ by $\mathcal{S}$. ∎
Lemma 68 ([4], Proposition 2.15 (i)).
Suppose $R\leqslant_{sd}A_{1}\times A_{2}$, $B_{1}$ and $B_{2}$ are absorbing subuniverses on $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$, respectively, $(R\cap(B_{1}\times B_{2}))\leqslant_{sd}B_{1}\times B_{2}$, $R$ is linked. Then $(R\cap(B_{1}\times B_{2}))$ is linked.
Lemma 69 ([34], Theorem 6.15).
Suppose $\mathbf{R}\leqslant_{sd}\mathbf{A}\times\mathbf{B}$, $C=\{c\in A\mid\forall b\in B\colon(c,b)\in R\}$. Then one of the following conditions holds:
  1. 1.
    $C$ is a central subuniverse of $\mathbf{A}$;
  2. 2.
    $\mathbf{B}$ has a nontrivial binary absorbing subuniverse.
Lemma 70 ([11], Theorem 3.11.1).
Suppose $R\lneq_{sd}\mathbf{A}\times\mathbf{B}$, $R$ is linked. Then there exists a BA or central subuniverse on $\mathbf{A}$ or $\mathbf{B}$.
Lemma 71 ([33], Lemma 7.2).
Suppose $0_{\mathbf{A}}\subseteq\sigma\leqslant{\mathbf{A}}^{2}$ and $\omega<_{BA}\sigma$. Then $\omega\cap 0_{\mathbf{A}}\neq\varnothing$.
The following Lemma can be derived from Lemma 3.11.2 and 3.11.3 from [11], but we will give a separate proof.
Lemma 72.
Suppose $R\leqslant_{sd}A\times B$, $A$ is BA and center free, $\operatorname{LeftLinked}(R)=A^{2}$, and $C=\{c\in B\mid A\times\{c\}\subseteq R\}$. Then $C\neq\varnothing$ and $C\leqslant_{\mathcal{BA},\mathcal{C}}B$.
Proof.
First, let us show that $C\neq\varnothing$. For every $n\geqslant 2$ put $W_{n}=\{(a_{1},\dots,a_{n})\mid\exists b\forall i\colon R(a_{i},b)\}$. If $W_{|A|}=A^{|A|}$, then $C\neq\varnothing$. Otherwise, choose the minimal $n\geqslant 2$ such that $W_{n}\neq A^{n}$. Since $\operatorname{LeftLinked}(R)=A^{2}$, $\operatorname{LeftLinked}(W_{n})=A^{2}$. Looking at $W_{n}$ as at binary relation $W_{n}\leqslant A\times A^{n-1}$ and using Lemmas 70 and 67 we derive the existence of BA or central subuniverse on $A$, which contradicts our assumptions.
Thus, $C\neq\varnothing$. By Lemma 69, $C$ is a central subuniverse of $B$. It remains to show that it is also a BA subuniverse. Assume the converse, then by Lemma 66 there exists a relation $S\leqslant B\times B$ such that $S\cap(C\times C)=\varnothing$, $S\cap(B\times C)\neq\varnothing$, and $S\cap(C\times B)\neq\varnothing$.
Put $W=\{(a_{1},\dots,a_{|A|})\mid\exists(b,c)\in S\colon c\in C\wedge\forall i(a_{i},b)\}$. By Lemma 62, $W<_{\mathcal{C}}A^{|A|}$, which by Lemma 67 implies the existence of a central subuniverse on $A$ and contradicts our assumptions. ∎
Lemma 73.
Suppose $f\colon\mathbf{A}\to\mathbf{A}^{\prime}$ is a surjective homomorphism and $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{D}\}$. Then $C^{\prime}<_{T(\sigma)}^{A}B^{\prime}\Rightarrow f^{-1}(C^{\prime})<_{T(f^{-1}(\sigma))}^{A}f^{-1}(B^{\prime})$.
Proof.
For $T\in\{\mathcal{BA},\mathcal{C}\}$ it follows from the properties of a homomorphism (see Section 3.15 in [11]). For $T=\mathcal{S}$ there exists $D^{\prime}\leqslant C$ such that $D^{\prime}<_{\mathcal{BA},\mathcal{C}}B^{\prime}$. Then $f^{-1}(D^{\prime})<_{BA,\mathcal{C}}f^{-1}(B^{\prime})$, hence $f^{-1}(C^{\prime})<_{S}f^{-1}(B^{\prime})$.
Suppose $T\in\{\mathcal{PC},\mathcal{L}\}$. Let $\delta=f^{-1}(\sigma)$, that is $\delta=\{(a,b)\mid(f(a),f(b))\in\sigma\}$. Then $\mathbf{A}^{\prime}/\sigma=\mathbf{A}/\delta$, $\mathbf{B}^{\prime}/\sigma\cong\mathbf{B}/\delta$ and $\mathbf{C}^{\prime}/\sigma\cong\mathbf{C}/\delta$, which implies the required properties of a PC/linear subuniverse. ∎
Corollary 74.
Suppose $R\leqslant_{sd}A_{1}\times\dots\times A_{n}$, $C_{1}<_{T(\sigma)}^{A_{1}}B_{1}\leqslant A_{1}$ where $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{D}\}$. Then $R\cap(C_{1}\times A_{2}\times\dots\times A_{n})<_{T(\sigma)}^{R}R\cap(B_{1}\times A_{2}\times\dots\times A_{n})$.
Proof.
It is sufficient to consider a homomorphism $f_{1}\colon\mathbf{R}\to\mathbf{A}_{1}$ sending every tuple to its first coordinate and apply Lemma 73 to $f_{1}^{-1}$. ∎
5.3 Intersection property
In this subsection we prove a fundamental property of our subuniverses. Precisely, we will show (Lemma 78) that if $C\lll A$ and $\delta$ is a dividing congruence for $B\lll A$, then $(B\cap C)/\delta$ is either empty, or of size 1, or equal to $B/\delta$.
Lemma 75.
Suppose $R\leqslant\mathbf{A}^{n}$ is symmetric, $\mathbf{A}$ is BA and center free, $\operatorname{pr}_{1,2}(R)=A^{2}$, and $(a_{1},\dots,a_{n})\in R\Rightarrow(a_{1},a_{1},a_{2}\dots,a_{n-1})\in R$. Then $R=A^{n}$.
Proof.
Let us prove the claim by induction on $n$. For $n=2$ it follows from the condition $\operatorname{pr}_{1,2}(R)=A^{2}$. Suppose $n>2$. Since $\operatorname{pr}_{1,n}(R)=A^{2}$ for any $a,c\in A$ we have $(a,b_{2},\dots,b_{n-1},c)\in R$ for some $b_{2},\dots,b_{n-1}\in A$. By the conditions of the lemma we have $(a,\dots,a,c),(a,\dots,a,a)\in R$. Since $a$ and $c$ can be chosen arbitrary, we have $\operatorname{RightLinked}(R)=A^{2}$. By the inductive assumption $\operatorname{pr}_{1,2,\dots,n-1}(R)=A^{n-1}$. Then Lemma 70 implies the existence of BA or central subuniverse on $A$ or $A^{n-1}$, which together with Lemma 67 contradicts the fact that $\mathbf{A}$ is BA and center free. ∎
Lemma 76.
Suppose $\sigma$ is a dividing congruence for $B\leqslant A$, $R\leqslant(\mathbf{A}/\sigma)^{n}$, $R$ is reflexive, symmetric, and $(a_{1},\dots,a_{n})\in R\Rightarrow(a_{1},a_{1},a_{2}\dots,a_{n-1})\in R$. Then either $(B/\sigma)^{n}\subseteq R$, or $R=\{(a/\sigma,a/\sigma,\dots,a/\sigma)\mid a\in A\}$.
Proof.
Consider two cases:
Case 1. $\operatorname{pr}_{1,2}(R)$ is the equality relation. Since $R$ is symmetric, we derive that $R=\{(a/\sigma,a/\sigma,\dots,a/\sigma)\mid a\in A\}$.
Case 2. $\operatorname{pr}_{1,2}(R)$ is not the equality relation. Since $\sigma$ is irreducible and $R$ is reflexive, we derive that $\operatorname{pr}_{1,2}(R)\supseteq\sigma^{*}/\sigma\supseteq(B/\sigma)^{2}$. It remains to apply Lemma 75 to $R\cap(B/\sigma)^{n}$ to show that $(B/\sigma)^{n}\subseteq R$. ∎
Lemma 77.
Suppose $B_{k}<_{T_{k}(\sigma_{k})}^{A}B_{k-1}<_{T_{k-1}(\sigma_{k-1})}^{A}\dots<_{T_{2}(\sigma_{2})}^{A}B_{1}<_{T_{1}(\sigma_{1})}^{A}B_{0}=A$, $\delta$ is a congruence on $A$, $T_{1},T_{2},\dots,T_{k}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{D}\}$, $m\in\{1,2,\dots,k\}$, and $T_{m}=\mathcal{D}$. Then $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}$ is either $B_{m-1}/\sigma_{m}$, or $B_{m}/\sigma_{m}$. Additionally, if $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m}/\sigma_{m}$ then $\sigma_{m}\supseteq(\delta\cap\sigma_{1}\cap\dots\cap\sigma_{m-1})$.
Proof.
We prove the statement by induction on $k$. Consider two cases:
Case 1. $k=m$. Define $|A|$-ary relation $S_{0}\leqslant(A/\sigma_{k})^{|A|}$ by
$S_{0}=\{(a_{1}/\sigma_{k},\dots,a_{|A|}/\sigma_{k})\mid\forall i,j\colon(a_{i},a_{j})\in\delta\cap\bigcap_{\ell=1}^{k-1}\sigma_{\ell}\}.$
By Lemma 76 we have one of the two subcases:
Subcase 1A. $S_{0}=\{(b/\sigma_{k},\dots,b/\sigma_{k})\mid b\in A\}.$ Then $\sigma_{k}\supseteq(\delta\cap\sigma_{1}\cap\dots\cap\sigma_{k-1})$, which is the additional condition we needed to prove. Therefore, $\delta\cap B_{k-1}^{2}\subseteq\sigma_{k}\cap B_{k-1}^{2}$, $(B_{k}\circ\delta)\cap B_{k-1}\subseteq(B_{k}\circ\sigma_{k})\cap B_{k-1}=B_{k}$, and $((B_{k}\circ\delta)\cap B_{k-1})/\sigma_{k}=B_{k}/\sigma_{k}$.
Subcase 1B. $(B_{k-1}/\sigma_{k})^{|A|}\subseteq S_{0}.$ For $n\in\{0,1,\dots,k\}$ define $S_{n}$ by
$S_{n}=\{(a_{1}/\sigma_{k},\dots,a_{|A|}/\sigma_{k})\mid\forall i,j\colon(a_{i},a_{j})\in\delta\cap\bigcap_{\ell=1}^{k-1}\sigma_{\ell},\forall i\colon a_{i}\in B_{n}\}.$
Notice that for $n=0$ we get a definition of $S_{0}$. Assume that $(B_{k-1}/\sigma_{k})^{|A|}\subseteq S_{k-1}$. Then there exists an equivalence class $E$ of $\delta$ such that $(E\cap B_{k-1})/\sigma_{k}=B_{k-1}/\sigma_{k-1}$. Hence, $((B_{k}\circ\delta)\cap B_{k-1})/\sigma_{k}=B_{k-1}/\sigma_{k}$, which completes this case. Thus, we assume that $(B_{k-1}/\sigma_{k})^{|A|}\not\subseteq S_{k-1}$. Consider the minimal $n$ such that $(B_{k-1}/\sigma_{k})^{|A|}\not\subseteq S_{n}.$ Consider three subsubcases.
Subsubcase 1B1. $T_{n}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65, we obtain that $S_{n}\cap(B_{k-1}/\sigma_{k})^{|A|}<_{T_{n}}(B_{k-1}/\sigma_{k})^{|A|}$, which by Lemma 67 implies the existence of BA or central subuniverse on $B_{k-1}/\sigma_{k}$ and contradicts the definition of a dividing congruence.
Subsubcase 1B2. $T_{n}=\mathcal{S}$. Choose $G\leqslant B_{n}$ such that $G<_{\mathcal{BA},\mathcal{C}}B_{n-1}$. By Lemma 64 $G\cap B_{k}\neq\varnothing$, hence even if we replace $B_{n}$ by $G$ in the definition of $S_{n}$ the intersection $S_{n}\cap(B_{k-1}/\sigma_{k})^{|A|}$ will be nonempty. Then Lemmas 62 and 65 imply the existence of both BA and central subuniverse on $B_{k-1}/\sigma_{k}$ and contradicts the definition of a dividing congruence.
Subsubcase 1B3. $T_{n}=\mathcal{D}$. Define a new relation $R\leqslant(A/\sigma_{k})^{|A|}\times A/\sigma_{n}$ by
$R=\{(a_{1}/\sigma_{k},\dots,a_{|A|}/\sigma_{k},b/\sigma_{n})\mid\forall i,j\colon(a_{i},a_{j})\in\delta\cap\bigcap_{\ell=1}^{k-1}\sigma_{\ell},\forall i\colon a_{i}\in B_{n-1},(a_{i},b)\in\sigma_{n}\}.$
Put $R^{\prime}=R\cap((B_{k-1}/\sigma_{k})^{|A|}\times B_{n-1}/\sigma_{n})$. By the choice of $n$ we have $\operatorname{pr}_{1,2,\dots,|A|}(R^{\prime})=(B_{k-1}/\sigma_{k})^{|A|}$ and $(B_{k-1}/\sigma_{k})^{|A|}\times B_{n}/\sigma_{n}\not\subseteq R^{\prime}$. Notice that $\operatorname{pr}_{|A|+1}(R^{\prime})=((B_{k-1}\circ\sigma_{k})\cap B_{n-1})/\sigma_{n}$. By the inductive assumption either $\operatorname{pr}_{|A|+1}(R^{\prime})=B_{n-1}/\sigma_{n}$, or $\operatorname{pr}_{|A|+1}(R^{\prime})=B_{n}/\sigma_{n}$. The second case contradicts the above conditions, therefore $\operatorname{pr}_{|A|+1}(R^{\prime})=B_{n-1}/\sigma_{n}$.
Since $\operatorname{pr}_{1,2,\dots,|A|}(R^{\prime})=(B_{k-1}/\sigma_{k})^{|A|}$, there exists $d\in B_{n-1}/\sigma_{n}$ such that $(B_{k-1}/\sigma_{k})^{|A|}\times\{d\}\subseteq R^{\prime}$. Then $R^{\prime}$ can be viewed as a binary relation with a center containing $d$. Then Lemma 69 implies the existence of a BA subuniverse on $B_{k-1}/\sigma_{k}$ or a center on $B_{n-1}/\sigma_{n}$, which contradicts the definition of a dividing subuniverse.
Case 2. $k>m$. By the inductive assumption $((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}$ is either $B_{m-1}/\sigma_{m}$, or $B_{m}/\sigma_{m}$. In the second case we also have $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m}/\sigma_{m}$, which completes this case. Thus, we assume that $((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m-1}/\sigma_{m}$. Let us consider three cases depending on the type of the reduction $T_{k}$.
Subcase 2A. $T_{k}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65 we obtain $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}\leqslant_{T_{k}}((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}$, which by the definition of a dividing congruence implies $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}=((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m-1}/\sigma_{m}$ and completes this case.
Subcase 2B. $T_{k}=\mathcal{S}$. Choose $G\leqslant B_{k}$ such that $G<_{\mathcal{BA},\mathcal{C}}B_{k-1}$. Then $((G\circ\delta)\cap B_{m-1})/\sigma_{m}\leqslant_{\mathcal{BA},\mathcal{C}}((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}$, which by the definition of a dividing congruence implies $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}=((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m-1}/\sigma_{m}$ and completes this case.
Subcase 2C. $T_{k}=\mathcal{D}$. Define a binary relation $R\leqslant B_{k-1}/\sigma_{k}\times B_{m-1}/\sigma_{m}$ by
$R=\{(a/\sigma_{k},b/\sigma_{m})\mid a\in B_{k-1},b\in B_{m-1},(a,b)\in\delta\}.$
By the above assumption $R$ is subdirect. Since $k>m$, we have $B_{k-1}/\sigma_{k}\times B_{m}/\sigma_{m}\subseteq R$, hence the right-hand side of $R$ has a center. By the definition of a dividing congruence both $B_{k-1}/\sigma_{k}$ and $B_{m-1}/\sigma_{m}$ are BA and center-free. Then Lemma 69 implies that $R=B_{k-1}/\sigma_{k}\times B_{m-1}/\sigma_{m}$ and $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m-1}/\sigma_{m}$. ∎
Lemma 78.
Suppose $B\lll A$, $C\lll A$, $B\cap C\neq\varnothing$. Then
  1. (d)
    if $\delta$ is a dividing congruence for $B\leqslant A$, then $|(B\cap C)/\delta|=1$ or $(B\cap C)/\delta=B/\delta$. Moreover, if $|(B\cap C)/\delta|=1$ then $\delta\supseteq\delta_{1}\cap\dots\cap\delta_{s}$, where $\delta_{1},\dots,\delta_{s}$ are the dividing congruences from the definition of $B\lll A$ and $C\lll A$.
  2. (s)
    if $G<_{\mathcal{BA},\mathcal{C}}B$, then $G\cap C\neq\varnothing$.
Proof.
Suppose
$B=B_{k}<_{T_{k}(\sigma_{k})}^{A}B_{k-1}<_{T_{k-1}(\sigma_{k-1})}^{A}\dots<_{T_{2}(\sigma_{2})}^{A}B_{1}<_{T_{1}(\sigma_{1})}^{A}B_{0}=A$
$C=C_{\ell}<_{\mathcal{T}_{\ell}(\omega_{\ell})}^{A}C_{\ell-1}<_{\mathcal{T}_{\ell-1}(\omega_{\ell-1})}^{A}\dots<_{\mathcal{T}_{2}(\omega_{2})}^{A}C_{1}<_{\mathcal{T}_{1}(\omega_{1})}^{A}C_{0}=A$
where $k$ and $\ell$ were chosen minimal. Put $\sigma=\bigcap_{i=1}^{k}\sigma_{i}$ and $\omega=\bigcap_{i=1}^{\ell}\omega_{i}$. We prove the lemma by induction on $k+\ell$.
Let us prove (s) first. By the inductive assumption $G\cap C_{\ell-1}\neq\varnothing$. Let us consider the type $\mathcal{T}_{\ell}$ of the subuniverse $C_{\ell}$.
$\mathcal{T}_{\ell}=\mathcal{D}$. By the inductive assumption $(B\cap C_{\ell-1})/\omega_{\ell}$ is either of size 1, or equal to $C_{\ell-1}/\omega_{\ell}$. In the first case we have $G\cap C_{\ell}=G\cap B\cap C_{\ell}=G\cap B\cap C_{\ell-1}=G\cap C_{\ell-1}\neq\varnothing$. In the second case $(G\cap C_{\ell-1})/\omega_{\ell}<_{\mathcal{BA},\mathcal{C}}C_{\ell-1}/\omega_{\ell}$, which contradicts the definition of a divisible congruence $\omega_{\ell}$.
$\mathcal{T}_{\ell}\in\{\mathcal{BA},\mathcal{C}\}$. By Lemma 63 $G\cap C_{\ell-1}<_{\mathcal{BA},\mathcal{C}}B_{k}\cap C_{\ell-1}$, $B_{k}\cap C_{\ell}<_{\mathcal{T}_{\ell}}B_{k}\cap C_{\ell-1}$, and by Lemma 61 $G\cap C_{\ell}\neq\varnothing$.
$\mathcal{T}_{\ell}=\mathcal{S}$. Let $G^{\prime}\leqslant C_{\ell}$ and $G^{\prime}<_{\mathcal{BA},\mathcal{C}}C_{\ell-1}$. By the inductive assumption $B_{k}\cap G^{\prime}\neq\varnothing$. Then by Lemma 63 $G\cap C_{\ell-1}<_{\mathcal{BA},\mathcal{C}}B_{k}\cap C_{\ell-1}$, $B_{k}\cap G^{\prime}<_{\mathcal{BA},\mathcal{C}}B_{k}\cap C_{\ell-1}$, and by Lemma 61 $G\cap G^{\prime}\neq\varnothing$. Hence $G\cap C_{\ell}\neq\varnothing$.
Let us prove (d). Define an $|A|$-ary relation $S\leqslant(A/\delta)^{|A|}$ by
$S=\{(a_{1}/\delta,\dots,a_{|A|}/\delta)\mid\forall i,j\colon(a_{i},a_{j})\in\sigma\cap\omega\}.$
By Lemma 76 we have one of the two cases.
Case 1. $S=\{(b/\delta,\dots,b/\delta)\mid b\in A\}.$ Therefore, $\sigma\cap\omega\subseteq\delta$, which is the additional property we need. Notice that $(a,b)\in\sigma\cap\omega$ for any $a,b\in B\cap C$, hence $a/\delta=b/\delta$ for any $a,b\in B\cap C$, and $|(B\cap C)/\delta|=1$, which completes this case.
Case 2. $(B/\delta)^{|A|}\subseteq S.$ For $m\in\{0,1,\dots,k\}$ and $n\in\{0,1,\dots,\ell\}$ put
$S_{m,n}=\{(a_{1}/\delta,\dots,a_{|A|}/\delta)\mid\forall i,j\colon(a_{i},a_{j})\in\sigma\cap\omega,\forall i\colon a_{i}\in B_{m},a_{i}\in C_{n}\}.$
Notice that $S_{0,0}=S$. Consider two subcases.
Subcase 2A. $(B/\delta)^{|A|}\not\subseteq S_{k,0}$. Choose a minimal $m$ such that $(B/\delta)^{|A|}\not\subseteq S_{m,0}$. Consider 3 subsubcases depending on the type $T_{m}$.
Subsubcase 2A1. $T_{m}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65, we obtain $S_{m,0}\cap(B/\delta)^{|A|}<_{T_{m}}(B/\delta)^{|A|}$, which by Lemma 67 implies the existence of BA or central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2A2. $T_{m}=\mathcal{S}$. Choose $G\leqslant B_{m}$ such that $G<_{\mathcal{BA},\mathcal{C}}B_{m-1}$. By Lemma 64 $G\cap B\neq\varnothing$, hence even if we replace $B_{m}$ by $G$ in the definition of $S_{m,0}$ the intersection $S_{m,0}\cap(B/\delta)^{|A|}$ will be nonempty. Then Lemmas 62 and 65 imply the existence of both BA and central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2A3. $T_{m}=\mathcal{D}$. Define a new relation $R\leqslant(A/\delta)^{|A|}\times A/\sigma_{m}$ by
$R=\{(a_{1}/\delta,\dots,a_{|A|}/\delta,b/\sigma_{m})\mid\forall i,j\colon(a_{i},a_{j})\in\sigma\cap\omega,\forall i\colon a_{i}\in B_{m-1},(a_{i},b)\in\sigma_{m}\}.$
Put $R^{\prime}=R\cap((B/\delta)^{|A|}\times B_{m-1}/\sigma_{m})$. By the choice of $m$ we have $\operatorname{pr}_{1,2,\dots,|A|}(R^{\prime})=(B/\delta)^{|A|}$ and $(B/\delta)^{|A|}\times B_{m}/\sigma_{m}\not\subseteq R^{\prime}$. Also, $\operatorname{pr}_{|A|+1}(R^{\prime})=((B\circ\delta)\cap B_{m-1})/\sigma_{m}$. Lemma 77 implies that $((B\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m-1}/\sigma_{m}$ Since $\operatorname{pr}_{1,2,\dots,|A|}(R^{\prime})=(B/\delta)^{|A|}$, there exists $d\in B_{m-1}/\sigma_{m}$ such that $(B/\delta)^{|A|}\times\{d\}\subseteq R^{\prime}$. Then $R^{\prime}$ can be viewed as a binary relation with a center containing $d$. Then Lemma 69 implies the existence of a BA subuniverse on $B/\delta$ or a center on $B_{m-1}/\sigma_{m}$, which contradicts the definition of a dividing subuniverse.
Subcase 2B. $(B/\delta)^{|A|}\subseteq S_{k,0}$ and $(B/\delta)^{|A|}\not\subseteq S_{k,\ell}$. Choose a minimal $n$ such that $(B/\delta)^{|A|}\not\subseteq S_{k,n}$. Consider 3 subsubcases depending on the type $\mathcal{T}_{n}$.
Subsubcase 2B1. $\mathcal{T}_{n}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65, we obtain $S_{k,n}\cap(B/\delta)^{|A|}<_{\mathcal{T}_{n}}(B/\delta)^{|A|}$, which by Lemma 67 implies the existence of BA or central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2B2. $\mathcal{T}_{n}=\mathcal{S}$. Choose $G\leqslant C_{n}$ such that $G<_{\mathcal{BA},\mathcal{C}}C_{n-1}$. By the inductive assumption $G\cap B_{k}\neq\varnothing$. Then even if we replace $C_{n}$ by $G$ in the definition of $S_{m,n}$ we get a nonempty intersection with $(B/\delta)^{|A|}$. By Lemmas 62 and 65 we obtain both BA and central subuniverse on $(B/\delta)^{|A|}$, which by Lemma 67 implies the existence of BA and central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2B3. $\mathcal{T}_{n}=\mathcal{D}$. Define a new relation $R\leqslant(A/\delta)^{|A|}\times A/\omega_{n}$ by
$R=\{(a_{1}/\delta,\dots,a_{|A|}/\delta,b/\omega_{n})\mid\forall i,j\colon(a_{i},a_{j})\in\sigma\cap\omega,\forall i\colon a_{i}\in B_{k},a_{i}\in C_{n-1},(a_{i},b)\in\omega_{n}\}.$
Put $R^{\prime}=R\cap((B/\delta)^{|A|}\times C_{n-1}/\omega_{n})$. By the choice of $n$ we have $\operatorname{pr}_{1,2,\dots,|A|}(R^{\prime})=(B/\delta)^{|A|}$ and $(B/\delta)^{|A|}\times C_{n}/\omega_{n}\not\subseteq R^{\prime}$. Also, $\operatorname{pr}_{|A|+1}(R^{\prime})=(B_{k}\cap C_{n-1})/\omega_{n}$. By the inductive assumption either $(B_{k}\cap C_{n-1})/\omega_{n}=C_{n-1}/\omega_{n}$, or $(B_{k}\cap C_{n-1})/\omega_{n}|=C_{n}/\omega_{n}$. In the second case we have $S_{k,n-1}\cap(B/\delta)^{|A|}=S_{k,n}\cap(B/\delta)^{|A|}$, which contradicts the choice of $n$. Thus, we assume that $\operatorname{pr}_{|A|+1}(R^{\prime})=C_{n-1}/\omega_{n}$. Since $\operatorname{pr}_{1,2,\dots,|A|}(R^{\prime})=(B/\delta)^{|A|}$, there exists $d\in C_{n-1}/\omega_{n}$ such that $(B/\delta)^{|A|}\times\{d\}\subseteq R^{\prime}$. Then $R^{\prime}$ can be viewed as a binary relation with a center containing $d$. Then Lemma 69 implies the existence of a BA subuniverse on $B/\delta$ or a center on $C_{n-1}/\omega_{n}$, which contradicts the definition of a dividing congruence.
Subcase 2C. $(B/\delta)^{|A|}\subseteq S_{k,\ell}$. Hence, $(B\cap C)/\delta=B/\delta$, which completes the proof. ∎
Corollary 79.
Suppose $R\leqslant_{sd}\mathbf{A}_{1}\times\mathbf{A}_{2}$, $B_{1}\lll A_{1}$, $B_{2}\lll A_{2}$, $\sigma$ is a dividing congruence for $B_{1}\lll A_{1}$. Then $\operatorname{pr}_{1}(R\cap(B_{1}\times B_{2}))/\sigma$ is either empty, or of size 1, or equal to $B_{1}/\sigma$.
Proof.
Consider homomorphisms $f_{1}\colon\mathbf{R}\to\mathbf{A}_{1}$ and $f_{2}\colon\mathbf{R}\to\mathbf{A}_{2}$ that maps each tuple to the first and the second coordinate respectively. By Lemma 73 $f_{1}^{-1}(B_{1})\lll\mathbf{R}$ and $f_{2}^{-1}(B_{2})\lll\mathbf{R}$. Put $\delta=f^{-1}(\sigma)$. By Lemma 78 $(f_{1}^{-1}(B_{1})\cap f_{2}^{-1}(B_{2}))/\delta$ is either empty, or of size 1, or equal to $R/\delta$. Translating this into the language of $\sigma$ and $R\cap(B_{1}\times B_{2})$ we obtained the required statement. ∎
Corollary 80.
Suppose $B_{1}\lll A$, $B_{2}\lll A$, $C_{1}<_{\mathcal{D}(\sigma_{1})}^{A}B_{1}$, $C_{1}^{\prime}<_{\mathcal{D}(\sigma_{1})}^{A}B_{1}$, $C_{2}<_{\mathcal{D}(\sigma_{2})}^{A}B_{2}$, $C_{2}^{\prime}<_{\mathcal{D}(\sigma_{2})}^{A}B_{2}$, $C_{1}^{\prime}\cap C_{2}\neq\varnothing$, $C_{1}\cap C_{2}^{\prime}\neq\varnothing$, and $C_{1}^{\prime}\cap C_{2}^{\prime}\neq\varnothing$. Then $C_{1}\cap C_{2}\neq\varnothing$.
Proof.
By Lemma 78 $(B_{1}\cap B_{2})/\sigma_{1}$ is either of size 1, or equal to $B_{1}/\sigma_{1}$. In the first case we have $C_{1}\cap B_{2}=B_{1}\cap B_{2}$ and $C_{1}\cap C_{2}=B_{1}\cap C_{2}\neq\varnothing$. Similarly, if $(B_{1}\cap B_{2})/\sigma_{2}$ is of size 1, then $C_{1}\cap C_{2}=C_{1}\cap B_{2}\neq\varnothing$.
Thus, we assume that $(B_{1}\cap B_{2})/\sigma_{1}=B_{1}/\sigma_{1}$ and $(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$. Let $S=\{(a/\sigma_{1},a/\sigma_{2})\mid a\in B_{1}\cap B_{2}\}$. Then $S\leqslant_{sd}(B_{1}/\sigma_{1})\times(B_{2}/\sigma_{2})$.
Applying Lemma 78 (d) to $C<_{\mathcal{D}(\sigma_{1})}^{A}B_{1}$, $B_{2}\lll A$, and the congruence $\sigma_{2}$ we derive one of the two cases:
Case 1. There exists $a\in B_{1}/\sigma_{1}$ such that $(a,b)\in S$ for every $b\in B_{2}/\sigma_{2}$. If $S$ is full, we immediately obtain $C_{1}\cap C_{2}\neq\varnothing$. Otherwise, Lemma 69 implies the existence of a BA or center on $B_{1}/\sigma_{1}$ or $B_{2}/\sigma_{2}$, which contradicts our assumptions.
Case 2. For every $a\in B_{1}/\sigma_{1}$ there exists a unique $b\in B_{2}/\sigma_{2}$ such that $(a,b)\in S$. Choosing $a\in B_{1}/\sigma_{1}$ corresponding to $C_{1}^{\prime}$ we derive that $C_{2}=C_{2}^{\prime}$. Hence, $C_{1}\cap C_{2}=C_{1}\cap C_{2}^{\prime}\neq\varnothing$. ∎
5.4 Properties of PC or Linear congruences
To prove the following lemma we will need several standard algebraic definitions. Two algebras $\mathbf{A}$ and $\mathbf{B}$ are called polynomially equivalent if $A=B$ the clone generated by $\mathbf{A}$ and all constants operations coincides with the clone generated by $\mathbf{B}$ and all constant operations. An algebra is called affine if it is polynomially equivalent to an $\mathbf{R}$-module. An algebra is called Abelian if all operations $t\in\operatorname{Clo}({\mathbf{A}})$ (of arbitrary arity $n+1$) satisfy the following condition
$\forall x,y,u_{1},\dots,u_{n},v_{1},\dots,v_{n}\in A\colon$
$t(x,u_{1},\dots,u_{n})=$ $t(x,v_{1},\dots,v_{n})\Rightarrow t(y,u_{1},\dots,u_{n})=t(y,v_{1},\dots,v_{n}).$
An equivalent definition is given in the following lemma.
Lemma 81 ([20]).
An algebra $\mathbf{A}$ is Abelian if and only if there exists a congruence $\delta$ on $\mathbf{A}^{2}$ such that $\{(a,a)\mid a\in A\}$ is a block of $\delta$.
Lemma 82.
An algebra $\mathbf{A}$ is Abelian if and only if there exists a bridge $\delta$ from $0_{\mathbf{A}}$ to $0_{\mathbf{A}}$ such that $\widetilde{\delta}=\operatorname{pr}_{1,2}(\delta)=\operatorname{pr}_{3,4}(\delta)=A^{2}$.
Proof.
A congruence on $\mathbf{A}^{2}$ from Lemma 81 is a required bridge.
To obtain a congruence from a bridge, compose the bridge with itself sufficient number of times to obtain a reflexive symmetric transitive relation on $\mathbf{A}^{2}$. ∎
Lemma 83 ([20]).
Suppose a finite algebra $\mathbf{A}$ has a WNU term operation. Then $\mathbf{A}$ is Abelian if and only if it is affine.
Lemma 84.
Suppose $\sigma$ is a congruence on $\mathbf{A}$, $\delta$ is a bridge from $\sigma$ to $\sigma$ such that $\operatorname{pr}_{1,2}(\delta)=\widetilde{\delta}=A^{2}$ and $\delta(x_{1},x_{2},x_{3},x_{4})=\delta(x_{3},x_{4},x_{1},x_{2})$. Then there exists an abelian group $(G;+,-)$ such that $(A/\sigma;\delta/\sigma)\cong(G;x_{1}-x_{2}=x_{3}-x_{4})$.
Proof.
By Lemmas 82 and 83, $\mathbf{A}/\sigma$ is affine and, therefore, polynomially equivalent to an $\mathbf{R}$-module $\mathbf{G}=(G,+,0,-,(r)_{r\in\mathbf{R}})$. Composing the bridge $\delta$ with itself we get a bridge $\delta_{0}\supseteq\delta$ which is an equivalence relation on $\mathbf{A}^{2}$. Then $\delta_{0}/\sigma$ is a congruence on the $\mathbf{R}$-module $\mathbf{G}^{2}$. Congruences in $\mathbf{R}$-modules come from submodules ($u$ and $v$ are congruent if $u-v$ is in a submodule). Since the diagonal of $(A/\sigma)^{2}$ is the block of the congruence $\delta_{0}/\sigma$, the corresponding submodule is $\{(x,y):x=y\}$. Hence $\delta_{0}/\sigma=\{((x_{1},y_{1}),(x_{2},y_{2})):x_{1}-x_{2}=y_{1}-y_{2}\}$. It remains to show that $\delta=\delta_{0}$. Since $\delta/\sigma$ is preserved by the Maltsev operation $x-y+z$ and $\widetilde{\delta}=\operatorname{pr}_{1,2}(\delta)=A^{2}$, we derive that $\operatorname{pr}_{1,2,3}(\delta)=A^{3}$. Since the last element of $\delta/\sigma$ is uniquely determined by the first three, we derive $\delta=\delta_{0}$. ∎
Lemma 85.
Suppose $\delta$ is a bridge from $0_{A}$ to $0_{A}$, $\operatorname{pr}_{1,2}(\delta)\subseteq\widetilde{\delta}$, $\operatorname{pr}_{1,2}(\delta)$ is linked. Then $A$ is BA and center free.
Proof.
We prove by induction on the size of $A$. First, we build a symmetric bridge $\sigma$ from $\delta$ by
$\sigma(x_{1},x_{2},x_{3},x_{4})=\exists x_{5}\exists x_{6}\;\delta(x_{1},x_{2},x_{5},x_{5})\wedge\delta(x_{3},x_{4},x_{5},x_{6}).$
It follows from the definition of the bridge that $\sigma$ satisfies all the properties satisfied by $\delta$ but additionally we have $\sigma(x_{1},x_{2},x_{3},x_{4})=\sigma(x_{3},x_{4},x_{1},x_{2})$. Put $\omega=\operatorname{pr}_{1,2}(\sigma)$.
Assume that there exists $B<_{T}A$, where $T\in\{\mathcal{BA},\mathcal{C}\}$. Since $\operatorname{pr}_{1,2}(\sigma)$ is linked either $((A\setminus B)\times B)\cap\omega\neq\varnothing$, or $(B\times(A\setminus B)\cap\omega\neq\varnothing$. Without loss of generality let it be the first case. Then choose $(a,b)\in((A\setminus B)\times B)\cap\omega$. Consider two cases:
Case 1. There exists a subalgebra $D\lneq A$ such that $a,b\in D$. Let $E$ be the maximal subuniverse of $D$ such that $\omega\cap E^{2}$ is linked and $a,b\in E$. Then we apply the inductive assumption to $\sigma\cap E^{4}$ and derive that $E\cap B$ cannot be a proper BA or central subuniverse. By Lemma 63 $E\cap B<_{T}E$, which gives a contradiction.
Case 2. There does not exist a subalgebra $D\lneq A$ such that $a,b\in D$. Hence $\{a\}\circ\omega=A$. Put $\xi(x_{1},x_{2})=\sigma(a,x_{1},x_{2},b)$. Since $(a,b,a,b)$ and $(a,a,b,b)$ are from $\sigma$, both $\operatorname{pr}_{1}(\xi)$ and $\operatorname{pr}_{2}(\xi)$ contains $a$ and $b$. Be the definition of case 2 we derive $\operatorname{pr}_{1}(\xi)=\operatorname{pr}_{2}(\xi)=A$. Put $C=B\circ\xi$. By Lemma 62 $C\leqslant_{T}A$. Notice that $b\notin C$ as otherwise we would have a tuple $(a,c,b,b)\in\sigma$ for some $c\in B$, which is not possible because $a\notin B$. Thus, we have $B<_{T}A$, $C<_{T}A$, $b\in B$, $a\in C$, $b\notin C$. If $T=BA$ then applying the binary absorbing operation to $(a,b,a,b)$ and $(a,a,b,b)$ we get a tuple $(a,b_{1},b_{2},b)$, where $b_{1},b_{2}\in B$. This implies that $C\cap B\neq\varnothing$. If $T=C$ then by Corollary 22 we have
($(\{a\}\times B\times C\times B)\cap\sigma\neq\varnothing$
($(\{a\}\times C\times C\times C)\cap\sigma\neq\varnothing\;\;\;\;\;\;\;\Rightarrow\;\;\;\;(\{a\}\times C\times C\times B)\cap\sigma\neq\varnothing$
($(\{a\}\times C\times B\times B)\cap\sigma\neq\varnothing$
Consider $C^{\prime}=\operatorname{pr}_{2}(\sigma\cap(\{a\}\times C\times C\times B))$. Using Lemma 62 and the fact that $\operatorname{pr}_{1}(\xi)=A$, we derive $C^{\prime}\leqslant_{C}C$. Moreover, by the definition of a bridge $a\notin C^{\prime}$ or $C\cap B\neq\varnothing$. Thus, in all the remaining cases we either have $C\cap B\neq\varnothing$, or $C^{\prime}<_{C}C$. In the first case put $C^{\prime}=B\cap C$.
To complete the proof we consider the case when $\varnothing\neq C^{\prime}<_{T}C<_{T}A$. Notice that $|C|>1$. Applying the inductive assumption to $\sigma\cap C^{4}$ and using the fact that $\{a\}\circ\omega=A$, we get a contradiction with $C^{\prime}<_{T}C$. ∎
Lemma 86.
Suppose $\sigma$ is an irreducible congruence on $A$, $\delta$ is a bridge from $\sigma$ to $\sigma$ such that $\widetilde{\delta}\supsetneq\sigma$. Then
  1. 1.
    $\sigma^{*}$ is a congruence;
  2. 2.
    $B/\sigma$ is BA and center free for each block $B$ of $\sigma^{*}$.
  3. 3.
    if $\delta(x_{1},x_{2},x_{3},x_{4})=\delta(x_{3},x_{4},x_{1},x_{2})$, then there exists a prime $p$ such that for every block $B$ of $\sigma^{*}$ we have $(B/\sigma;(\delta\cap B^{4})/\sigma)\cong(\mathbb{Z}_{p}^{n_{B}};x_{1}-x_{2}=x_{3}-x_{4})$, where $n_{B}\geqslant 0$.
Proof.
Since the relations $\sigma$, $\sigma^{*}$, $\delta$ are stable under $\sigma$ we can factorize them by $\sigma$ and consider $0_{A/\sigma}$, $\sigma^{*}/\sigma$, and $\delta/\sigma$ instead. To avoid new notations we assume that $\sigma=0_{A}$.
Consider some block $B$ of $\operatorname{LeftLinked}(\sigma^{*})$ that is not a block of $\sigma$ (equivalently, of size greater than one). Put $\delta^{\prime}=\delta\cap B^{4}\cap(\sigma^{*}\times\sigma^{*})$. Since $\widetilde{\delta}\supseteq\sigma^{*}$, $\delta^{\prime}$ satisfies all the conditions of Lemma 85, which implies that $B$ is BA and center free. Applying Lemma 70 to $\sigma^{*}\cap B^{2}$ we derive that $B^{2}\subseteq\sigma^{*}$. Therefore, $\operatorname{LeftLinked}(\sigma^{*})=\sigma^{*}$ and $\sigma^{*}$ is a congruence.
To prove the rest consider a block $B$ of $\sigma^{*}$ of size at least 2 and apply Lemma 84 to the bridge $\delta^{\prime}=\delta\cap B^{4}$. Then, $(B;\delta)\cong(G_{B};x_{1}-x_{2}=x_{3}-x_{4})$ for some Abelian group $(G_{B};+,-)$. It remains to show there exists a prime $p$ such that each Abelian group $G_{p}$ is isomorphic to $(\mathbf{Z}_{p}^{n};+,-)$.
First, we simplify the bridge $\delta$ and consider $\omega=\delta\cap(\sigma^{*}\times\sigma^{*})$. Notice that $\omega$ satisfies the same properties.
Assume that $G_{B}$ has elements of coprime orders $p_{1}$ and $p_{2}$ or $G_{B}$ has a element of order $p_{1}^{k}$, where $k\geqslant 2$. Composing the relation $x_{1}-x_{2}=x_{3}-x_{4}$ we can define a relation $(k+1)\cdot x_{1}=k\cdot x_{2}+x_{3}$ for any $k\in\mathbb{N}$. In fact
$((k+1)\cdot x_{1}=k\cdot x_{2}+x_{3})=\exists x_{4}\;(k\cdot x_{1}=(k-1)\cdot x_{2}+x_{4})\wedge(x_{1}-x_{2}=x_{3}-x_{4}).$
Hence $p_{1}\cdot x_{1}=p_{1}\cdot x_{2}$ is also pp-definable from $x_{1}-x_{2}=x_{3}-x_{4}$. Let this pp-definition define the binary relation $S$ if we replace $x_{1}-x_{2}=x_{3}-x_{4}$ by $\omega$. It follows from the definition that $S$ is a congruence on $A$ and $S\supsetneq\sigma$ and $\sigma^{*}\not\subseteq S$, which contradicts the irreducibility of $\sigma$. Hence the order of any element of $G_{B}$ is a prime number. Similarly, if elements of different groups $G_{B_{1}}$ and $G_{B_{2}}$ have different orders, we we can define a similar relation “$p_{1}\cdot x_{1}=p_{1}\cdot x_{2}$” and again get a contradiction with the irreducibility of $\sigma$. ∎
Lemma 7.
Suppose $\sigma$ is an irreducible congruence on $\mathbf{A}$. Then the following conditions are equivalent:
  1. 1.
    $\sigma$ is a linear congruence;
  2. 2.
    there exists a bridge $\delta$ from $\sigma$ to $\sigma$ such that $\widetilde{\delta}\supsetneq\sigma$.
Proof.
$1\Rightarrow 2$. By property 3 of linear congruence we have a bridge $\delta$ such that $\widetilde{\delta}\supseteq\sigma$.
$2\Rightarrow 1$. We derive another bridge $\delta^{\prime}(x_{1},x_{2},x_{3},x_{4})=\exists x_{5}\exists x_{6}\;\delta(x_{1},x_{2},x_{5},x_{6})\wedge\delta(x_{3},x_{4},x_{5},x_{6})$. and apply Lemma 86 to it. ∎
Lemma 87.
Suppose $\sigma$ is a congruence on $\mathbf{A}$, $\delta$ is a reflexive bridge from $\sigma$ to $\sigma$ satisfying
  1. 1.
    $\delta(x_{1},x_{2},x_{3},x_{4})=\delta(x_{3},x_{4},x_{1},x_{2})$;
  2. 2.
    $(a,b,a,b),(b,a,b,a)\in\delta$ for every $(a,b)\in\operatorname{pr}_{1,2}(\delta)$;
  3. 3.
    $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\delta))=A^{2}$.
Then there exists $a,b\in A$ such that $a\neq b$ and $(a,a,b,b)\in\delta$.
Proof.
As before, we assume that $\sigma$ is the equality relation as otherwise we can factorize all the relations by $\sigma$. We prove the lemma by induction on the size of $A$. Consider two cases:
Case 1. There exists $B<_{T}A$ such that $|B|>1$ and $T\in\{\mathcal{BA},\mathcal{C}\}$. Put $\delta^{\prime}=\delta\cap B^{4}$. By Lemma 68 $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\delta^{\prime}))=B^{2}$. Assume that $\delta^{\prime}$ is not a bridge, then $\operatorname{pr}_{1,2}(\delta^{\prime})$ is the equality relation. Then $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\delta^{\prime}))=B^{2}$ implies the existence of $(a,a,b,b)\in\delta^{\prime}$ such that $a\neq b$, which is what we need. If $\delta^{\prime}$ is a bridge then the inductive assumption implies the existence of the corresponding $(a,a,b,b)\in\delta^{\prime}$.
Case 2. Otherwise, there exists $\{a\}<_{T}A$, where $T\in\{\mathcal{BA},\mathcal{C}\}$. Choose $(a,b,c,d)\in\delta$ such that $c\neq a$, which exists because $\operatorname{pr}_{1,3}(\delta)$ is linked. If $a=b$ then $c=d$ and we found the required pair $(a,c)$. If $\{a\}\circ\operatorname{pr}_{1,3}(\delta)\neq A$ then $\{a\}\circ\operatorname{pr}_{1,3}(\delta)<_{T}A$ and $a,c\in\{a\}\circ\operatorname{pr}_{1,3}$, which is Case 1. Otherwise, consider a tuple $(a,e,b,f)\in\delta$. By the assumption on $\delta$ we have $(b,a,b,a)\in\delta$. Let $g$ be a ternary absorbing operation on $A$ witnessing that $\{a\}$ absorbs $A$. Applying this operation to the tuples $(a,e,b,f),(b,a,b,a),(a,a,a,a)$ we obtain $(a,a,g(b,b,a),a)\in\delta$, hence $g(b,b,a)=a$. It remains to apply $g$ to $(b,a,b,a),(b,a,b,a),(a,e,b,f)$ and obtain $(a,a,b,a)\in\delta$ which contradicts the definition of a bridge.
Case 3. There does not exist a BA or central subuniverse on $A$. Consider $\operatorname{pr}_{1,2,3}(\delta)$ as a binary relation and put $C=\{(a,b)\mid\{a\}\times\{b\}\times A\subseteq\operatorname{pr}_{1,2,3}(\delta)\}$. By Lemma 72 $C\leqslant_{BA}\operatorname{pr}_{1,2}(\delta)$. Since $(a,a)\in\operatorname{pr}_{1,2}(\delta)$ for any $a\in A$, Lemma 71 implies that $(a,a)\in C$ for some $a\in A$. Then $(a,a,b,b)\in\delta$ for any $b\neq a$, which completes this case. ∎
Lemma 9.
Suppose $\sigma$ is a PC congruence on $A$. Then any reflexive bridge $\delta$ from $\sigma$ to $\sigma$ such that $\operatorname{pr}_{1,2}(\delta)=\operatorname{pr}_{3,4}(\delta)=\sigma^{*}$ can be represented as $\delta(x_{1},x_{2},x_{3},x_{4})=\sigma(x_{1},x_{3})\wedge\sigma(x_{2},x_{4})$ or $\delta(x_{1},x_{2},x_{3},x_{4})=\sigma(x_{1},x_{4})\wedge\sigma(x_{2},x_{3})$.
Proof.
Define a new bridge $\xi$ by
$\xi(x_{1},x_{2},x_{5},x_{6})=\exists x_{3}\exists x_{4}\;\;\delta(x_{1},x_{2},x_{3},x_{4})\wedge\delta(x_{5},x_{6},x_{3},x_{4}).$
Consider $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\xi))$ and $\operatorname{RightLinked}(\operatorname{pr}_{1,2,4}(\xi))$. Since $\sigma$ is irreducible, we have one of the three cases:
Case 1. $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\xi))=\operatorname{RightLinked}(\operatorname{pr}_{1,2,4}(\xi))=\sigma$. Hence, for any $(a,b,c,d)\in\xi$ the elements $c/\sigma$ and $d/\sigma$ are uniquely determined by $a/\sigma$ and $b/\sigma$. Since $\xi$ is symmetric, we have $(a,b,a,b)\in\xi$. Therefore, $(a,c)\in\sigma$ and $(b,d)\in\sigma$. Since $\operatorname{pr}_{1,2}(\delta)=\operatorname{pr}_{3,4}(\delta)$, for any $(a,b,c,d)\in\delta$ the elements $c/\sigma$ and $d/\sigma$ are also uniquely determined by $a/\sigma$ and $b/\sigma$.
Define two new relations $\zeta_{1}$ and $\zeta_{2}$ and check whether one of them is a bridge showing that $\sigma$ is linear. Put
$\zeta_{1}(x_{1},x_{2},x_{3},x_{4})=$ $\exists y\exists y^{\prime}\exists z\exists z^{\prime}\exists t_{1}\exists t_{2}\exists t_{3}\exists t_{4}\;\;$
$\delta(x_{1},y,z,t_{1})\wedge\delta(x_{2},y,z^{\prime},t_{2})\wedge\delta(x_{3},y^{\prime},z,t_{3})\wedge\delta(x_{4},y^{\prime},z^{\prime},t_{4})$
$\zeta_{2}(x_{1},x_{2},x_{3},x_{4})=$ $\exists y\exists z\;\;\delta(x_{1},y,z,x_{3})\wedge\delta(x_{2},y,z,x_{4})$
Choose some $(a,b,c,d)\in\delta$ such that $(a,b)\in\sigma^{*}\setminus\sigma$. Since $(a,b,c,d),(b,b,b,b)\in\delta$, we have $(a,b,a,b)\in\zeta_{1}$ and $\operatorname{pr}_{1,2}(\zeta_{1})\supseteq\sigma^{*}$. Consider several subcases:
Subcase 1A. $\zeta_{1}$ is a bridge. Since $\widetilde{\zeta}_{1}$ must be equal to $\sigma$ and $(a,a,c,c)\in\zeta_{1}$ for any $(a,b,c,d)\in\delta$, we have $\operatorname{pr}_{1,3}(\delta)=\sigma$.
Subcase 1B. $\zeta_{1}$ is not a bridge. Then there exists $(a,a,b,c)\in\zeta_{1}$ such that $(b,c)\notin\sigma$. Let the evaluation of the variables in the definition of $\zeta_{1}$ be $y=d$, $y^{\prime}=d^{\prime}$, $z=e$, $z^{\prime}=e^{\prime}$, and $t_{i}=f_{i}$ for $i=1,2,3,4$. Since the first two coordinates of $\delta$ uniquely (up to $\sigma$) determine the last two, we have $(e,e^{\prime})\in\sigma$. Hence $(b,d^{\prime},e,f_{3}),(c,d^{\prime},e,f_{4})\in\delta$. Hence $\operatorname{pr}_{1,2}(\zeta_{2})\supsetneq\sigma$ and using the fact that the first two and the last two coordinates of $\delta$ uniquely determine each other, we derive that $\zeta_{2}$ is a bridge. Since $\widetilde{\zeta}_{2}$ must be equal to $\sigma$, we have $\operatorname{pr}_{1,4}(\delta)=\sigma$.
Thus, we derived that either $\operatorname{pr}_{1,4}(\delta)=\sigma$ or $\operatorname{pr}_{1,3}(\delta)=\sigma$. Repeating the same argument but switching $x_{1}$ and $x_{2}$ we derive that $\operatorname{pr}_{2,4}(\delta)=\sigma$ or $\operatorname{pr}_{2,3}(\delta)=\sigma$. This completes this case.
Case 2. $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\xi))\supseteq\sigma^{*}$. Choose a block $B$ of $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\xi))$ that is not a block of $\sigma$. Let us check that $a,b,d\in B$ for any $c\in B$ and $(a,b,c,d)\in\xi$. Since $(a,b,a,b)\in\xi$, we have $a\in B$. Since $\operatorname{RightLinked}(\operatorname{pr}_{1,2,3}(\xi))\supseteq\sigma^{*}$ and $(a,b),(c,d)\in\sigma^{*}$, we have $b,d\in B$. It remains to apply Lemma 87 to $\xi\cap B^{4}$.
Case 3. $\operatorname{RightLinked}(\operatorname{pr}_{1,2,4}(\xi))\supseteq\sigma^{*}$. This case can be considered in the same way as Case 2. ∎
Lemma 8.
Suppose $\sigma_{1}$ is a linear congruence, $\sigma_{2}$ is an irreducible congruence, $\delta$ is a bridge from $\sigma_{1}$ to $\sigma_{2}$. Then $\sigma_{2}$ is a also linear congruence.
Proof.
Without loss of generality we can assume that the relation $\widetilde{\delta}$ is rectangular as otherwise we can compose it with itself many times to obtain rectangularity. To simplify we replace the bridge $\delta$ by $\delta\cap(\sigma_{1}^{*}\times\sigma_{2}^{*})$. Let $\sigma_{1}$ and $\sigma_{2}$ be congruences on algebras $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$, respectively.
Assume that $\operatorname{RightLinked}(\widetilde{\delta})\supsetneq\sigma_{2}$. Then composing $\delta$ with itself we derive a bridge from $\sigma_{2}$ to $\sigma_{2}$ witnessing that $\sigma_{2}$ is linear.
Thus, assume that $\operatorname{RightLinked}(\widetilde{\delta})=\sigma_{2}$. Notice that $\operatorname{LeftLinked}(\widetilde{\delta})\supseteq\sigma_{1}^{*}$ as otherwise $\operatorname{LeftLinked}(\widetilde{\delta})=\sigma_{1}$, $A_{1}/\sigma_{1}\cong A_{2}/\sigma_{2}$, and $\sigma_{2}$ is also linear.
Put $\delta^{\prime}(x_{1},x_{2},x_{3},x_{4})=\delta(x_{1},x_{2},x_{3},x_{4})\wedge\widetilde{\delta}(x_{1},x_{3})$ and consider two cases:
Case 1. $\delta^{\prime}$ is a bridge, then we obtained a new bridge with the property $\operatorname{pr}_{1,3}(\delta^{\prime})=\widetilde{\delta}^{\prime}$.
Case 2. $\delta^{\prime}$ is not a bridge. Hence, there does not exist $(a,b,c,d)\in\delta$ such that $(a,b)\notin\sigma_{1}$ and $(a,c)\in\widetilde{\delta}$. Put $\delta^{\prime\prime}(x_{1},x_{2},x_{3},x_{4})=\exists z\;\widetilde{\delta}(x_{1},x_{3})\wedge\delta(x_{2},z,x_{3},x_{4})$. Let us show that $\delta^{\prime\prime}$ is a bridge. If $(x_{1},x_{2})\in\sigma_{1}$ then $(x_{2},x_{3})\in\widetilde{\delta}$ and by the assumption we have $(x_{2},z)\in\sigma_{1}$ and $(x_{3},x_{4})\in\sigma_{2}$. If $(x_{3},x_{4})\in\sigma_{2}$ then $(x_{2},z)\in\sigma_{1}$ and $(x_{2},x_{3})\in\widetilde{\delta}$. Hence $(x_{1},x_{2})\in\sigma_{1}$. As $\delta^{\prime}\neq\delta$, there exists $(a,b,c,d)\in\delta$ with $(a,c)\notin\widetilde{\delta}$. Choosing $e\in A_{1}$ such that $(e,c)\in\widetilde{\delta}$ we derive that $(e,a,c,d)\in\delta^{\prime\prime}$ and $(e,a)\notin\sigma_{1}$. Hence, $\delta^{\prime\prime}$ is a bridge.
Thus, in both cases we build a bridge $\omega$ from $\sigma_{1}$ to $\sigma_{2}$ such that $\operatorname{pr}_{1,3}(\omega)=\widetilde{\omega}=\widetilde{\delta}$.
Assume that $\sigma_{2}$ is not linear. Define a bridge $\xi_{1}$ by
$\xi_{1}(x_{1},x_{2},x_{3},x_{4})=\exists x_{5}\exists x_{6}\;\;\omega(x_{5},x_{6},x_{1},x_{2})\wedge\omega(x_{5},x_{6},x_{3},x_{4}).$
By Lemma 9 $\xi_{1}(x_{1},x_{2},x_{3},x_{4})=\sigma_{2}(x_{1},x_{3})\wedge\sigma_{2}(x_{2},x_{4})$. Similarly, we define a bridge $\xi_{2}$ by
$\xi_{2}(x_{1},x_{2},x_{3},x_{4})=\exists x_{5},x_{6}\;\;\omega(x_{5},x_{6},x_{1},x_{2})\wedge\omega(x_{6},x_{5},x_{3},x_{4}).$
Using the facts that $(x_{5},x_{1}),(x_{6},x_{3})\in\widetilde{\omega}$, $(x_{5},x_{6})\in\sigma_{1}^{*}$, and $\operatorname{LeftLinked}(\widetilde{\omega})\supseteq\sigma_{1}^{*}$ we derive that $(x_{1},x_{3})\in\sigma_{2}$. Then Lemma 9 implies that $\xi_{2}(x_{1},x_{2},x_{3},x_{4})=\sigma_{2}(x_{1},x_{3})\wedge\sigma_{2}(x_{2},x_{4})$. Hence $(b,a,c,d)\in\omega$ whenever $(a,b,c,d)\in\omega$.
Define a new relation $\zeta$ by
$\zeta(x_{1},x_{2},x_{3},x_{4})=\exists y_{1}\exists y_{2}\exists y_{3}\;\;\omega(y_{1},y_{2},x_{1},x_{2})\wedge\omega(y_{1},y_{3},x_{1},x_{3})\wedge\omega(y_{2},y_{3},x_{1},x_{4}).$
If $x_{1}=x_{2}$ then $y_{1}=y_{2}$ and by the property of $\xi_{1}$ we have $x_{3}=x_{4}$. Consider some $(c,d)\in\sigma_{2}^{*}$. Then for some $(a,b)\in\sigma_{1}^{*}$ we have $(a,b,c,d)\in\omega$. Since $(a,b)\in\sigma_{1}^{*}\subseteq\operatorname{LeftLinked}(\widetilde{\omega})$ and $(a,c)\in\widetilde{\omega}$, we have $(a,a,c,c),(b,b,c,c)\in\omega$. Sending $(x_{1},x_{2},x_{3},x_{4})$ to $(c,d,c,d)$ and $(y_{1},y_{2},y_{3})$ to $(a,b,a)$ we witness that $(c,d,c,d)\in\zeta$. Sending $(x_{1},x_{2},x_{3},x_{4})$ to $(c,c,d,d)$ and $(y_{1},y_{2},y_{3})$ to $(a,a,b)$ we witness that $(c,c,d,d)\in\zeta$. Hence $\zeta(x_{1},x_{2},x_{3},x_{4})\wedge\zeta(x_{3},x_{4},x_{1},x_{2})$ defines a bridge witnessing that $\sigma_{2}$ is linear. ∎
Lemma 10.
Suppose $\delta$ is a bridge from a PC congruence $\sigma_{1}$ on $\mathbf{A}_{1}$ to an irreducible congruence $\sigma_{2}$ on $\mathbf{A}_{2}$, $\operatorname{pr}_{1,2}(\delta)=\sigma_{1}^{*}$, and $\operatorname{pr}_{3,4}(\delta)=\sigma_{2}^{*}$. Then
  1. 1.
    $\sigma_{2}$ is a PC congruence;
  2. 2.
    $\mathbf{A}_{1}/\sigma_{1}\cong\mathbf{A}_{2}/\sigma_{2}$;
  3. 3.
    $\{(a/\sigma_{1},b/\sigma_{2})\mid(a,b)\in\widetilde{\delta}\}$ is bijective;
  4. 4.
    $\delta(x_{1},x_{2},x_{3},x_{4})=\widetilde{\delta}(x_{1},x_{3})\wedge\widetilde{\delta}(x_{2},x_{4})$ or $\delta(x_{1},x_{2},x_{3},x_{4})=\widetilde{\delta}(x_{1},x_{4})\wedge\widetilde{\delta}(x_{2},x_{3})$.
Proof.
By Lemma 8 $\sigma_{2}$ must be also be PC congruence. Then composing the bridge $\delta$ with itself we must get a trivial bridge. Hence $\widetilde{\delta}$ is bijective and gives an isomorphism $\mathbf{A}_{1}/\sigma_{1}\cong\mathbf{A}_{2}/\sigma_{2}$. Define a new bridge $\xi$ by
$\xi(x_{1},x_{2},x_{5},x_{6})=\exists x_{3},x_{4}\;\delta(x_{1},x_{2},x_{3},x_{4})\wedge\widetilde{\delta}(x_{5},x_{3})\wedge\widetilde{\delta}(x_{6},x_{4}).$
Then $\xi$ satisfies Lemma 9, which immediately implies the required condition 4. ∎
5.5 Types interaction
Lemma 88 ([33], Lemma 8.19).
Suppose $\omega$, $\sigma_{1}$, and $\sigma_{2}$ are congruences on $\mathbf{A}$, $\omega\cap\sigma_{1}=\omega\cap\sigma_{2}$, and $\omega\setminus\sigma_{1}\neq\varnothing$. Then there exists a bridge $\delta$ from $\sigma_{1}$ to $\sigma_{2}$ such that $\widetilde{\delta}=\sigma_{1}\circ\sigma_{2}$.
Proof.
Let us define a bridge $\delta$ by
$\delta(x_{1},x_{2},y_{1},y_{2})=\exists z_{1}\exists z_{2}\;\sigma_{1}(x_{1},z_{1})\wedge\sigma_{1}(x_{2},z_{2})\wedge\omega(z_{1},z_{2})\wedge\sigma_{2}(z_{1},y_{1})\wedge\sigma_{2}(z_{2},y_{2}).$
As it follows from the definition, the first two variables of $\delta$ are stable under $\sigma_{1}$ and the last two variables are stable under $\sigma_{2}$.
Let us show that for any $(a_{1},a_{2},a_{3},a_{4})\in\delta$ we have $(a_{1},a_{2})\in\sigma_{1}\Leftrightarrow(a_{3},a_{4})\in\sigma_{2}$. In fact, if $(x_{1},x_{2})\in\sigma_{1}$, then $(z_{1},z_{2})\in\sigma_{1}$. Since $\omega\cap\sigma_{1}=\omega\cap\sigma_{2}$, we have $(z_{1},z_{2})\in\sigma_{2}$. Therefore, $(y_{1},y_{2})\in\sigma_{2}$.
Choose $(a,b)\in\omega\setminus\sigma_{1}$. Then $(a,b,a,b)\in\delta$ (put $z_{1}=a$, $z_{2}=b$), which gives the last necessary property of the bridge.
It follows immediately from the definition of $\delta$ that $\widetilde{\delta}=\sigma_{1}\circ\sigma_{2}$. ∎
Lemma 89.
Suppose $C_{1}<_{T_{1}(\sigma_{1})}^{A}B_{1}\lll A$, $C_{2}<_{T_{2}(\sigma_{2})}^{A}B_{2}\lll A$, $T_{1},T_{2}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{D}\}$, $C_{1}\cap B_{2}\neq\varnothing$, $B_{1}\cap C_{2}\neq\varnothing$, $C_{1}\cap C_{2}=\varnothing$. Then
  1. 1.
    $T_{1}=T_{2}\in\{\mathcal{BA},\mathcal{C},\mathcal{D}\}$;
  2. 2.
    if $T_{1}=T_{2}=\mathcal{D}$, then there is a bridge $\delta$ from $\sigma_{1}$ to $\sigma_{2}$ such that $\widetilde{\delta}=\sigma_{1}\circ\sigma_{2}$.
Proof.
Assume that $T_{1}=\mathcal{S}$. Choose $S_{1}\leqslant C_{1}$ such that $S_{1}<_{\mathcal{BA},\mathcal{C}}B_{1}$. By Lemma 78(s) $S_{1}\cap C_{2}\neq\varnothing$ and therefore $C_{1}\cap C_{2}\neq\varnothing$, which gives a contradiction.
Similarly, we prove that the case $T_{2}=\mathcal{S}$ cannot happen.
Assume that $T_{1},T_{2}\in\{\mathcal{BA},\mathcal{C}\}$. Then by Lemma 63 $C_{1}\cap B_{2}<_{T_{1}}B_{1}\cap B_{2}$, $B_{1}\cap C_{2}<_{T_{2}}B_{1}\cap B_{2}$ and the claim follows from Lemma 61.
Assume that $T_{1}\in\{\mathcal{BA},\mathcal{C}\}$ and $T_{2}=\mathcal{D}$. By Lemma 78(d), $(B_{1}\cap B_{2})/\sigma_{2}=C_{2}/\sigma_{2}$ or $(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$. First case would imply that $B_{1}\cap B_{2}=B_{1}\cap C_{2}$, which contradicts $C_{1}\cap C_{2}=\varnothing$ and $C_{1}\cap B_{2}\neq\varnothing$. Thus, we assume that $(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$. Combining Lemmas 63 and 65 we obtain $(C_{1}\cap B_{2})/\sigma_{2}<_{T_{1}}(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$, which contradicts the definition of a dividing congruence.
It remains to consider the case when $T_{1}=T_{2}=\mathcal{D}$. Let $\omega_{1},\dots,\omega_{s}$ be all the dividing congruences coming from $B_{1}\lll^{A}A$ and $B_{2}\lll^{A}A$. Put $\omega=\cap_{i=1}^{s}\omega_{i}$. By Lemma 78(d) we have $\sigma_{1}\supseteq\sigma_{2}\cap\omega$ and $\sigma_{2}\supseteq\sigma_{1}\cap\omega$. By choosing $c_{1}\in C_{1}\cap B_{2}$ and $c_{2}\in B_{1}\cap C_{2}$, we obtain $(c_{1},c_{2})\in\omega\setminus\sigma_{1}$. Thus, $\sigma_{1}\cap\omega=\sigma_{2}\cap\omega$ and $\omega\setminus\sigma_{1}\neq\varnothing$, and Lemma 88 implies the existence of a bridge $\delta$ from $\sigma_{1}$ to $\sigma_{2}$ such that $\widetilde{\delta}=\sigma_{1}\circ\sigma_{2}$. ∎
5.6 Factorization of strong subalgebras
In this section first we prove several technical lemmas, then we show we can factorize subalgebras of type $T$ by a congruence keeping the type of the subalgebra (see Lemma 93).
Lemma 90.
Suppose $R\leqslant_{sd}A_{1}\times A_{2}$, $R\cap(B_{1}\times B_{2})\neq\varnothing$, $B_{1}\lll A_{1}$, $B_{2}\lll A_{2}$, $\sigma$ is a congruence on $A_{1}$, $B_{1}/\sigma$ is BA and center free, there exists $c\in A_{2}$ such that $(E\times\{c\})\cap R\neq\varnothing$ for every $E\in B_{1}/\sigma$. Then there exists $c\in B_{2}$ such that $(E\times\{c\})\cap R\neq\varnothing$ for every $E\in B_{1}/\sigma$.
Proof.
Consider a minimal $B_{2}^{\prime}$ such that $B_{2}\lll^{A_{2}}B_{2}^{\prime\prime}<_{T(\delta)}^{A_{2}}B_{2}^{\prime}\lll A_{2}$ and $c$ can be chosen from $B_{2}^{\prime}$. Define
$S^{\prime}=\{(a_{1}/\sigma,\dots,a_{|A_{1}|}/\sigma)\mid\exists c\in B_{2}^{\prime}\;\forall i\colon a_{i}\in B_{1},(a_{i},c)\in R\},$
$S^{\prime\prime}=\{(a_{1}/\sigma,\dots,a_{|A_{1}|}/\sigma)\mid\exists c\in B_{2}^{\prime\prime}\;\forall i\colon a_{i}\in B_{1},(a_{i},c)\in R\}.$
If $T\in\{\mathcal{BA},\mathcal{C}\}$ then by Lemmas 63 and 65 $S^{\prime\prime}<_{T}S^{\prime}=(B_{1}/\sigma)^{|A_{1}|}$ and by Lemma 67 there exists a BA or central subuniverse on $B_{1}/\sigma$, which contradicts our assumptions.
If $T=\mathcal{S}$ then choose $D_{2}\leqslant B_{2}^{\prime\prime}$ such that $D_{2}<_{\mathcal{BA},\mathcal{C}}B_{2}^{\prime}$. Combining Corollary 74 and Lemma 78(s) we obtain that $R\cap(B_{1}\times D_{2})\neq\varnothing.$ Hence, Lemmas 63 and 65 imply the existence of both BA and central subuniverse on $S^{\prime}=(B_{1}/\sigma)^{|A_{1}|}$ and by Lemma 67 there exists a BA and central subuniverse on $B_{1}/\sigma$, which contradicts our assumptions.
Suppose $T=\mathcal{D}$. Define a relation $S\leqslant(B_{1}/\sigma)^{|A_{1}|}\times B_{2}^{\prime}/\delta$ by
$S=\{(a_{1}/\sigma,\dots,a_{|A_{1}|}/\sigma,b/\delta)\mid\exists c\in B_{2}^{\prime}\;\forall i\colon a_{i}\in B_{1},(a_{i},c)\in R,(c,b)\in\delta\}.$
By the choice of $B_{2}^{\prime}$ there exists $d\in B_{2}^{\prime}/\delta$ such that $(B_{1}/\sigma)^{|A_{1}|}\times\{d\}\subseteq S$ but $(B_{1}/\sigma)^{|A_{1}|}\times B_{2}^{\prime\prime}/\delta\neq S$. Since $\operatorname{pr}_{|A_{1}|+1}(S)/\delta=\operatorname{pr}_{2}(R\cap(B_{1}\times B_{2}^{\prime}))/\delta$, Corollary 79 implies that $\operatorname{pr}_{|A_{1}|+1}(S)/\delta=B_{2}^{\prime}/\delta$. Hence, $S$ is a central relation. Combining Lemmas 69 and 67 we get a contradiction with the assumption that $B_{1}/\sigma$ and $B_{2}^{\prime}/\delta$ are BA and center free. ∎
Lemma 91.
Suppose $R\leqslant_{sd}A_{1}\times A_{2}$, $R\cap(B_{1}\times B_{2})\neq\varnothing$, $B_{1}\lll A_{1}$, $B_{2}\lll A_{2}$, $\sigma$ is a congruence on $A_{1}$, $B_{1}/\sigma$ is BA and center free, $(\operatorname{LeftLinked}(R)\cap B_{1}^{2})/\sigma=(B_{1}/\sigma)^{2}$. Then $(\operatorname{LeftLinked}(R\cap(B_{1}\times B_{2})))/\sigma=(B_{1}/\sigma)^{2}$.
Proof.
By $R_{n}$ we denote the binary relation defined by $R_{n}=\underbrace{R\circ R^{-1}\circ R\circ R^{-1}\circ\dots\circ R\circ R^{-1}}_{2n}.$ For sufficiently large $n$ we have $(R_{n}\cap B_{1}^{2})/\sigma=(B_{1}/\sigma)^{2}$.
Consider two cases:
Case 1. $(R_{1}\cap B_{1}^{2})/\sigma=(B_{1}/\sigma)^{2}$. Applying Lemma 72 to $S=\{(a/\sigma,b)\mid a\in B_{1},(a,b)\in R\}$ we derive the existence of $c\in A_{2}$ such that for every $E\in B_{1}/\sigma$ we have $(E\times\{c\})\cap R\neq\varnothing$. By Lemma 90 $c$ can be chosen from $B_{2}$ and $(\operatorname{LeftLinked}(R\cap(B_{1}\times B_{2})))/\sigma=(B_{1}/\sigma)^{2}$.
Case 2. $(R_{1}\cap B_{1}^{2})/\sigma\neq(B_{1}/\sigma)^{2}$. Consider the maximal $n=2^{k}$ such that $(R_{n}\cap B_{1}^{2})/\sigma\neq(B_{1}/\sigma)^{2}$. Hence, $(R_{2n}\cap B_{1}^{2})/\sigma=(B_{1}/\sigma)^{2}$. Applying Lemma 72 to $S=\{(a/\sigma,b)\mid a\in B_{1},(a,b)\in R_{n}\}$ we derive the existence of $c\in A_{1}$ such that for every $E\in B_{1}/\sigma$ we have $(E\times\{c\})\cap R_{n}\neq\varnothing$. By Lemma 90, $c$ can be chosen from $B_{1}$. Hence, the relation $(R_{n}\cap B_{1}^{2})/\sigma$ is central, which by Lemma 69 implies that $(R_{n}\cap B_{1}^{2})/\sigma=(B_{1}/\sigma)^{2}$ and contradicts our assumption. ∎
Lemma 92.
Suppose, $\sigma$ is a dividing congruence for $B\lll A$, $\delta$ is a congruence on $A$, and $(\delta\cap B^{2})/\sigma\neq B^{2}/\sigma$. Then
  1. (1)
    $(\delta\cap B^{2})\subseteq(\sigma\cap B^{2})$;
  2. (2)
    $\sigma\supseteq\delta\cap\omega$;
  3. (3)
    $(\delta\vee(\sigma\cap\omega))\cap B^{2}=\sigma\cap B^{2}$;
  4. (4)
    $(\delta\vee(\sigma\cap\omega))\cap\omega=\sigma\cap\omega$;
where $\omega$ is the intersection of all the dividing congruences coming from $B\lll A$.
Proof.
Let us prove (2) first. Consider two equivalence classes $C_{1}$ and $C_{2}$ of $\sigma$ such that $C_{1}\cap B\neq\varnothing$, $C_{2}\cap B\neq\varnothing$, and $((C_{1}\cap B)\times(C_{2}\cap B))\cap\delta=\varnothing$. Then $C_{1}\cap B<_{\mathcal{D}(\sigma)}^{A}B\lll A$ and by Lemma 77 for $m=k$ and $B_{k}=C_{1}\cap B$ we obtain $(((C_{1}\cap B)\circ\delta)\cap B)/\sigma=\{C_{1}\}$ and $\sigma\supseteq\delta\cap\omega$.
(1) follows immediately from (2).
Let us prove (4). Consider the binary relation $R=\delta\circ(\sigma\cap\omega)$. Notice that $\operatorname{LeftLinked}(R)=\delta\vee(\sigma\cap\omega)$. Put $\delta^{\prime}=\operatorname{LeftLinked}(R)$. Consider two cases:
Case 1. $(\delta^{\prime}\cap B^{2})/\sigma=(B/\sigma)^{2}$. By Lemma 91 $\operatorname{LeftLinked}(R\cap B^{2})/\sigma=(B/\sigma)^{2}$. Then there exist $a_{1},a_{2},c\in B$ and $b_{1},b_{2}\in A$ such that $(a_{1},b_{1}),(a_{2},b_{2})\in\delta$, $(b_{1},c),(b_{2},c)\in\sigma\cap\omega$, $(a_{1},a_{2})\notin\sigma$. Since $a_{1},a_{2},c\in B$, we have $(a_{1},c),(a_{2},c)\in\omega$. Therefore, $(a_{1},b_{1}),(a_{2},b_{2})\in\omega$, hence by (2) we have $(a_{1},b_{1}),(a_{2},b_{2})\in\sigma$, which contradicts the assumption $(a_{1},a_{2})\notin\sigma$.
Case 2. $(\delta^{\prime}\cap B^{2})/\sigma\neq(B/\sigma)^{2}$. Applying (2) to $\delta^{\prime}$ we obtain $\sigma\supseteq\delta^{\prime}\cap\omega$. Since $\delta^{\prime}\supseteq(\sigma\cap\omega)$, we obtain $\delta^{\prime}\cap\omega=\sigma\cap\omega$.
Condition (3) immediately follows from (4). ∎
Lemma 93.
Suppose $B\lll A$, $\sigma$ is a congruence on $A$ such that $|B/\sigma|>1$, $B/\sigma$ is BA and center free. Then there exists a dividing congruence $\delta$ for $B\lll A$ such that $\delta\supseteq\sigma$.
Proof.
Let $\delta\supseteq\sigma$ be a maximal congruence such that $|B/\delta|>1$. It follows from Lemma 73 that $B/\delta$ is BA and center free. Let us show that $\delta$ is irreducible. Assume that it is not true and $\delta$ can be represented as an intersection of binary relations $S_{1},\dots,S_{k}\supsetneq\delta$. Then for some $i$ we have $B^{2}\not\subseteq S_{i}$. Notice that $\operatorname{LeftLinked}(S_{i})$ is a congruence that is larger than $\delta$, hence by the choice of $\delta$ we have $B^{2}\subseteq\operatorname{LeftLinked}(S_{i})$.
By Lemma 91 $\operatorname{LeftLinked}(S_{i}\cap(B\times B))/\sigma=(B/\sigma)^{2}$. Hence, $(S_{i}\cap(B\times B))/\sigma$ is a linked relation, which by Lemma 70 implies the existence of a BA or central subuniverse on $B/\delta$ and contradicts our assumption. Thus, the congruence $\delta$ is irreducible.
Lemma 94.
Suppose $\delta$ is a congruence on $A$, $C<_{T(\sigma)}^{A}B\lll A$, where $T\in\{\mathcal{PC},\mathcal{L}\}$. Then $C/\delta=B/\delta$, or $C/\delta<_{\mathcal{S}}B/\delta$, or $C/\delta<_{T}^{A/\delta}B/\delta$. Moreover, if $C/\delta<_{T}^{A/\delta}B/\delta$ then $\delta\cap B^{2}\subseteq\sigma\cap B^{2}$.
Proof.
Let $R\leqslant A/\sigma\times A/\delta$ be defined by $R=\{(a/\sigma,a/\delta)\mid a\in B\}$. Consider two cases:
Case 1. $\operatorname{LeftLinked}(R)=(B/\sigma)^{2}$. By Lemma 72 there exists $U\subseteq A/\delta$ such that $(B/\sigma)\times U\subseteq R$. Notice that $C/\delta\supseteq U$ hence $C/\delta=B/\delta$ or $C/\delta<_{\mathcal{S}}B/\delta$.
Case 2. $\operatorname{LeftLinked}(R)\neq(B/\sigma)^{2}$. Let $\omega$ be the intersection of all the dividing congruences coming from $B\lll A$. Since $(\delta\cap B^{2})/\sigma\neq B^{2}/\sigma$, by Lemma 92(4) $\sigma\cap\omega=\delta^{\prime}\cap\omega$, where $\delta^{\prime}=\delta\vee(\sigma\cap\omega)$. Since $B/\delta^{\prime}\cong B/\sigma$, $B/\delta^{\prime}$ is BA and center free. By Lemma 93 applied to $\delta^{\prime}$ there exists $\delta^{\prime\prime}\supseteq\delta^{\prime}$ that is a dividing congruence for $B\lll A$. By Lemma 92(2) $\delta^{\prime\prime}\cap\omega\subseteq\sigma$, which together with $\delta^{\prime\prime}\cap\omega\supseteq\delta^{\prime}\cap\omega=\sigma\cap\omega$ implies $\delta^{\prime\prime}\cap\omega=\sigma\cap\omega$. Hence $C/\delta<_{\mathcal{T}(\delta^{\prime\prime}/\delta)}^{A/\delta}B/\delta$, where $\mathcal{T}\in\{\mathcal{PC},\mathcal{L}\}$.
To show that $T=\mathcal{T}$, we apply Lemma 88 to $\delta^{\prime\prime}\cap\omega=\sigma\cap\omega$ and obtain a bridge from $\delta^{\prime\prime}$ to $\sigma$. This, by Lemma 8, implies that congruences $\delta^{\prime\prime}$ and $\sigma$ must be of the same type and therefore $T=\mathcal{T}$. ∎
5.7 Proof of the remaining statements from Section 2.3
Lemma 13.
Suppose $B\lll A$, $|B|>1$. Then there exists $C<_{T}^{A}B$, where $T\in\{\mathcal{BA},\mathcal{C},\mathcal{L},\mathcal{PC}\}$.
Proof.
If $B$ has a nontrivial BA or central subuniverse we take this subuniverse as $C$. Otherwise, we apply Lemma 93, where $\sigma$ is the equality relation. Then there exists an irreducible congruence $\delta$ for $B$. It remains to choose any block $D$ of $B/\delta$ and put $C=D\cap B$. ∎
Lemma 20.
Suppose $B\lll A$, $D\lll A$. Then
  1. (i)
    $B\cap D\dot{\lll}A$;
  2. (t)
    $C<_{T(\sigma)}^{A}B\Rightarrow C\cap D\dot{\leqslant}_{T(\sigma)}^{A}B\cap D$.
Proof.
Let us prove (t) first. If $T\in\{\mathcal{BA},\mathcal{C}\}$ then it follows from Lemma 63. If $T=\mathcal{S}$ then consider $E\leqslant C$ such that $E<_{\mathcal{BA},\mathcal{C}}B$. If $C\cap D=\varnothing$, then the lemma holds. Otherwise, Lemma 78(s) implies that $E\cap D\neq\varnothing$. By Lemma 63. $E\cap D<_{\mathcal{BA},\mathcal{C}}B\cap D$. Hence $C\cap D<_{\mathcal{S}}B\cap D$. If $T\in\{\mathcal{PC},\mathcal{L}\}$ then by Lemma 78(d) $(B\cap D)/\sigma$ is either empty, or of size 1, or equal to $B/\sigma$. In the first case $C\cap D=\varnothing$ and we are done. In the last case we have $C\cap D<_{T(\sigma)}^{A}B\cap D$. In the second case we either have $C\cap D=B\cap D$, or $C\cap D=\varnothing$, which is what we need.
Let us prove (i). Consider the sequence $B=B_{k}<_{T_{k}}B_{k-1}<_{T_{k-1}}\dots<_{T_{2}}B_{1}<_{T_{1}}<A$ and apply (t) to each $B_{i}<_{T_{i}}B_{i-1}$. Then we have $B\cap D\dot{\lll}^{A}D$, which together with $D\lll A$ implies $B\cap D\dot{\lll}A$. ∎
Lemma 14.
Suppose $f\colon\mathbf{A}\to\mathbf{A}^{\prime}$ is a surjective homomorphism, then
  1. (f)
    $C\lll^{A}B\Rightarrow f(C)\lll f(B)$;
  2. (b)
    $C^{\prime}\lll^{A^{\prime}}B^{\prime}\Rightarrow f^{-1}(C^{\prime})\lll f^{-1}(B)$;
  3. (ft)
    $C<_{T(\sigma)}^{A}B\lll A\Longrightarrow(f(C)=f(B)\text{ or }f(C)<_{S}f(B)\text{ or }f(C)<_{T}^{A^{\prime}}f(B))$;
  4. (bt)
    $C^{\prime}<_{T(\sigma)}^{A^{\prime}}B^{\prime}\Rightarrow f^{-1}(C^{\prime})<_{T(f^{-1}(\sigma))}^{A^{\prime}}f^{-1}(B^{\prime})$;
  5. (fs)
    $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$ and $C<_{T}B$ $\Longrightarrow f(C)\leqslant_{T}f(B)$;
  6. (fm)
    $C\leqslant_{\mathcal{M}T}^{A}B\lll A$ and $f(B)$ is S-free $\Longrightarrow f(C)\leqslant_{\mathcal{M}T}^{A^{\prime}}f(B)$;
  7. (bm)
    $C^{\prime}\leqslant_{\mathcal{M}T}^{A^{\prime}}B^{\prime}\lll A^{\prime}\Longrightarrow f^{-1}(C)\leqslant_{\mathcal{M}T}^{A}f^{-1}(B)$.
Proof.
Let $\delta$ be the congruence defined by the homomorphism $f$, that is $f(\mathbf{A})\cong\mathbf{A}/\delta$. (bt). It follows from Lemma 73.
(b). It is sufficient to apply (bt) several times.
(ft). It follows from Lemma 94.
(f). It is sufficient to apply (ft) several times.
(fs). It follows from Lemma 65.
(fm). Let $C=C_{1}\cap\dots\cap C_{t}$, where $C_{i}<_{T(\sigma_{i})}^{A}B$. We prove by induction on $t$. If $t=1$ then it follows from Lemma 94. Assume that $C_{i}/\delta=B/\delta$ for some $i$. To simplify notations assume that $i=t$. Then consider $D_{j}=C_{j}\cap C_{t}$ for $j=1,2,\dots,t-1$. By Lemma 20 $D_{j}\leqslant_{T(\sigma_{i})}^{A}C_{t}$ for every $j$. By the inductive assumption $C/\delta=(D_{1}\cap\dots\cap D_{t-1})/\delta\leqslant_{\mathcal{M}T}^{A}C_{t}/\delta=B/\delta$. Hence $C/\delta\leqslant_{\mathcal{M}T}^{A/\delta}B/\delta$. Thus, it remains to consider the case when $C_{i}/\delta\neq B/\delta$ for every $i$. By the additional condition to Lemma 94 we have $\delta\cap B^{2}\subseteq\sigma_{i}\cap B^{2}$. By Lemma 94 we have $C_{i}/\delta<_{T}^{A/\delta}B/\delta$.
Let us show that $(C_{1}\cap\dots\cap C_{t})/\delta=C_{1}/\delta\cap\dots\cap C_{t}/\delta$. The inclusion $\subseteq$ is obvious. Let us prove $\supseteq$. Suppose $E\in C_{1}/\delta\cap\dots\cap C_{t}/\delta$. Since $(C_{i}\circ\delta)\cap B=C_{i}$ for every $i$, we have $E\cap B\subseteq C_{1}\cap\dots\cap C_{t}$. Hence $E\in(C_{1}\cap\dots\cap C_{t})/\delta$.
Thus, we showed that $(C_{1}\cap\dots\cap C_{t})/\delta=C_{1}/\delta\cap\dots\cap C_{t}/\delta<_{\mathcal{M}T}^{A/\delta}B/\delta$.
(bm). It is sufficient to apply (bt) and consider the intersection of the corresponding PC and linear subuniverses. ∎
Corollary 16.
Suppose $\delta$ is a congruence on $\mathbf{A}$, $\mathbf{B},\mathbf{C}\leqslant\mathbf{A}$. Then
  1. (f)
    $C/\delta\lll^{A/\delta}B/\delta\Longleftrightarrow C\circ\delta\lll^{A}B\circ\delta$;
  2. (t)
    $C/\delta<_{T}^{A/\delta}B/\delta\Longleftrightarrow C\circ\delta<_{T}^{A}B\circ\delta$.
Proof.
(t,$\Rightarrow$). It follows from Lemma 73.
(t,$\Leftarrow$). if $T\in\{\mathcal{BA},\mathcal{C},\mathcal{S}\}$ then it follows from Lemma 65. If $T\in\{\mathcal{PC},\mathcal{L},\mathcal{D}\}$ then it follows immediately from the definition.
(f). To prove (f) it is sufficient to apply (t) several times. ∎
Corollary 17.
Suppose $\delta$ is a congruence on $\mathbf{A}$. Then
  1. (f)
    $C\lll^{A}B\Rightarrow C\circ\delta\lll^{A}B\circ\delta$;
  2. (t)
    $C<_{T(\sigma)}^{A}B\lll^{A}A\Longrightarrow(C\circ\delta=B\circ\delta\text{ or}C\circ\delta<_{\mathcal{S}}^{A}B\circ\delta\text{ or }C\circ\delta<_{T}^{A}B\circ\delta)$;
  3. (e)
    $\delta\subseteq\sigma$ and $C<_{T(\sigma)}^{A}B\lll^{A}A\Longrightarrow C\circ\delta<_{T}^{A}B\circ\delta$.
Proof.
(f). Corollary 15(f) implies $C/\delta\lll^{A/\delta}B/\delta$, then Corollary 16(f) implies $C\circ\delta\lll^{A}B\circ\delta$.
(t). Again the proof is just a combination of Corollary 15(t) and Corollary 16(t).
(e). We use (t) and notice that $C\circ\delta<_{\mathcal{S}}^{A}B\circ\delta$ implies that $B/\sigma$ has BA and central subuniverse, which contradicts the definition of a dividing congruence. ∎
Corollary 18.
Suppose $R\leqslant_{sd}A_{1}\times\dots\times A_{n}$, $B_{i}\lll A_{i}$ for $i\in[n]$. Then
  1. (r)
    $R\cap(B_{1}\times\dots\times B_{n}))\dot{\lll}R$;
  2. (r1)
    $\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))\dot{\lll}A_{1}$;
  3. (b)
    $\forall i\colon C_{i}\lll^{A_{i}}B_{i}\Longrightarrow(R\cap(C_{1}\times\dots\times C_{n}))\dot{\lll}^{R}(R\cap(B_{1}\times\dots\times B_{n}))$;
  4. (b1)
    $\forall i\colon C_{i}\lll^{A_{i}}B_{i}\Longrightarrow\operatorname{pr}_{1}(R\cap(C_{1}\times\dots\times C_{n}))\dot{\lll}^{A_{1}}\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))$;
  5. (m)
    $\forall i\colon C_{i}\leqslant_{\mathcal{M}T}^{A_{i}}B_{i}\Longrightarrow R\cap(C_{1}\times\dots\times C_{n})\dot{\leqslant}_{\mathcal{M}T}^{R}R\cap(B_{1}\times\dots\times B_{n})$;
  6. (m1)
    $\forall i\colon C_{i}\leqslant_{\mathcal{M}T}^{A_{i}}B_{i}$, $\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))$ is S-free $\Longrightarrow$
    $\operatorname{pr}_{1}(R\cap(C_{1}\times\dots\times C_{n}))\dot{\leqslant}_{\mathcal{M}T}^{A_{1}}\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))$.
Proof.
(r). Let $f_{i}:\mathbf{R}\to\mathbf{A}_{i}$ be the homomorphism sending each tuple to its $i$-th coordinate. By Lemma 14(b) $f_{i}^{-1}(B_{i})\lll R$ for every $i$. Then by Lemma 20(i) $R\cap(B_{1}\times\dots\times B_{n})=\bigcap_{i=1}^{n}f_{i}^{-1}(B_{i})\dot{\lll}R$.
(r1). Additionally to (r) we apply $f_{1}$ to the intersection and use Lemma 14(f).
(b). By (r) we have $R^{\prime}:=R\cap(B_{1}\times\dots\times B_{n}))\dot{\lll}R$ and $R^{\prime\prime}:=R\cap(C_{1}\times\dots\times C_{n}))\dot{\lll}R$. By Lemma 20(i) $R^{\prime\prime}=R^{\prime}\cap R^{\prime\prime}\dot{\lll}^{R}R^{\prime}$.
(b1). Again, additionally to (b) we apply $f_{1}$ and use Lemma 14(f).
(m). By the definition of type $\mathcal{M}T$ each $C_{i}$ can be represented as $C_{i,1}\cap\dots\cap C_{i,n_{i}}$ where $C_{i,j}\leqslant_{T}^{A_{i}}B_{i}$ for all $i$ and $j$. We will use notations $R^{\prime}$ and $R^{\prime\prime}$ from (b). By Lemma 14(b) $f_{i}^{-1}(C_{i,j})\leqslant_{T}^{R}f_{i}^{-1}(B_{i})$. By Lemma 20(t) and property (r), that we already proved, we have
$f_{i}^{-1}(C_{i,j})\cap R^{\prime}\leqslant_{T}^{R}f_{i}^{-1}(B_{i})\cap R^{\prime}=R^{\prime}.$
Hence $R^{\prime\prime}=R^{\prime}\cap\bigcap_{i=1}^{t}\bigcap_{j=1}^{n_{i}}f_{i}^{-1}(C_{i,j})\dot{\leqslant}_{\mathcal{M}T}^{R}R^{\prime}$.
(m1). It is sufficient to apply Corollary 15(m) to (m). ∎
Theorem 21.
Suppose
  1. 1.
    $C_{i}<_{T_{i}(\sigma_{i})}^{A}B_{i}\lll A$, where $T_{i}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{PC}\}$ for $i=1,2,\dots,n$, $n\geqslant 2$;
  2. 2.
    $\bigcap\limits_{i\in[n]}C_{i}=\varnothing$;
  3. 3.
    $B_{j}\cap\bigcap\limits_{i\in[n]\setminus\{j\}}C_{i}\neq\varnothing$ for every $j\in[n]$.
Then one of the following conditions hold:
  1. (ba)
    $T_{1}=\dots=T_{n}=\mathcal{BA}$;
  2. (l)
    $T_{1}=\dots=T_{n}=\mathcal{L}$ and for every $k,\ell\in[n]$ there exists a bridge $\delta$ from $\sigma_{k}$ and $\sigma_{\ell}$ such that $\widetilde{\delta}=\sigma_{k}\circ\sigma_{\ell}$;
  3. (c)
    $n=2$ and $T_{1}=T_{2}=\mathcal{C}$;
  4. (pc)
    $n=2$, $T_{1}=T_{2}=\mathcal{PC}$, and $\sigma_{1}=\sigma_{2}$.
Proof.
For $n=2$ it follows from Lemmas 89 and 8. Put $B_{2}^{\prime}=B_{2}\cap C_{3}\cap\dots\cap C_{n}$ and $C_{2}^{\prime}=C_{2}\cap B_{2}^{\prime}$. By Lemma 20 $C_{2}^{\prime}\leqslant_{T_{2}(\sigma_{2})}B_{2}^{\prime}\lll A$. Also we have $C_{1}\cap C_{2}^{\prime}=\varnothing$, $C_{1}\cap B_{2}^{\prime}\neq\varnothing$, and $B_{1}\cap C_{2}^{\prime}\neq\varnothing$. Then Lemmas 89 and 8 imply that $T_{1}=T_{2}$, and if $T_{1}\in\{\mathcal{PC},\mathcal{L}\}$ then there is a bridge $\delta$ from $\sigma_{1}$ to $\sigma_{2}$ such that $\widetilde{\delta}=\sigma_{1}\circ\sigma_{2}$. If $T_{1}=\mathcal{PC}$ then $\sigma_{1}\circ\sigma_{2}=\sigma_{1}=\sigma_{2}$, as otherwise composing the bridge $\delta$ with itself we would get a nontrivial bridge from $\sigma_{i}$ to $\sigma_{i}$. Additionally, this implies that $n$ cannot be greater than 2 for $T_{1}=\mathcal{PC}$, as in this case the intersection $C_{1}\cap C_{2}$ must be empty.
Thus, we proved the required conditions for $T_{1}(\sigma_{1})$ and $T_{2}(\sigma_{2})$. Similarly, we can prove this for any $T_{i}(\sigma_{i})$ and $T_{j}(\sigma_{j})$. ∎
Corollary 22.
Suppose
  1. 1.
    $R\leqslant_{sd}A_{1}\times\dots\times A_{n}$;
  2. 2.
    $C_{i}<_{T_{i}(\sigma_{i})}^{A_{i}}B_{i}\lll A_{i}$, where $T_{i}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{PC}\}$ for $i=1,2,\dots,n$, $n\geqslant 2$;
  3. 3.
    $R\cap(C_{1}\times\dots\times C_{n})=\varnothing$;
  4. 4.
    $R\cap(C_{1}\times\dots\times C_{j-1}\times B_{j}\times C_{j+1}\times\dots\times C_{n})\neq\varnothing$ for every $j\in[n]$.
Then one of the following conditions hold:
  1. (ba)
    $T_{1}=\dots=T_{n}=\mathcal{BA}$;
  2. (l)
    $T_{1}=\dots=T_{n}=\mathcal{L}$ and for every $k,\ell\in[n]$ there exists a bridge $\delta$ from $\sigma_{k}$ and $\sigma_{\ell}$ such that $\widetilde{\delta}=\sigma_{k}\circ\operatorname{pr}_{k,\ell}(R)\circ\sigma_{\ell}$;
  3. (c)
    $n=2$ and $T_{1}=T_{2}=\mathcal{C}$;
  4. (pc)
    $n=2$, $T_{1}=T_{2}=\mathcal{PC}$, $A_{1}/\sigma_{1}\cong A_{2}/\sigma_{2}$, and the relation $\{(a/\sigma_{1},b/\sigma_{2})\mid(a,b)\in R\}$ is bijective.
Proof.
Again, let $f_{i}:\mathbf{R}\to\mathbf{A}_{i}$ be the homomorphism sending each tuple to its $i$-th coordinate. Denote $C_{i}^{\prime}=f_{i}^{-1}(C_{i})$ and $B_{i}^{\prime}=f_{i}^{-1}(B_{i})$. By Lemma 14(b) and (bt) we have $C_{i}^{\prime}\leqslant{T_{i}(\sigma_{i}^{\prime})}^{R}B_{i}^{\prime}\lll R$, where $\sigma_{i}^{\prime}=f_{i}^{-1}(\sigma_{i})$. Since $B_{i}^{\prime}$ and $C_{i}^{\prime}$ satisfy conditions of Theorem 21 we obtain most of the properties and the only nontrivial one is the fact that $\widetilde{\delta}=\operatorname{pr}_{k,\ell}(R)$ for any $k$ and $\ell$. Notice that $\widetilde{\delta}$ for the bridge coming from Theorem 21 is equal to $\sigma_{k}^{\prime}\circ\sigma_{\ell}^{\prime}$. Translating congruences $\sigma_{k}^{\prime}$ and $\sigma_{\ell}^{\prime}$ to $\sigma_{k}$ and $\sigma_{\ell}$, we derive that $(a,b)\in\widetilde{\delta}$ if and only if there exists a tuple $(a_{1},\dots,a_{n})\in R$ such that $(a_{k},a)\in\sigma_{k}$ and $(a_{\ell},b)\in\sigma_{\ell}$. This implies $\widetilde{\delta}=\sigma_{k}\circ\operatorname{pr}_{k,\ell}(R)\circ\sigma_{\ell}$. The additional condition for $T_{1}=\mathcal{PC}$ follows from the fact that $\sigma_{1}^{\prime}=\sigma_{2}^{\prime}$. ∎
Lemma 23.
Suppose $C\leqslant_{\mathcal{M}T}^{A}B$. Then $C<_{T}^{A}\dots<_{T}^{A}B$ and $C\lll^{A}B$.
Proof.
Suppose $C=C_{1}\cap\dots\cap C_{n}$ where $C_{i}<_{T}^{A}B$. Put $D_{j}=\cap_{i=1}^{j}C_{i}$. By Lemma 20(t) we have $D_{j+1}\leqslant_{T}^{A}D_{j}$. Since $D_{n}=C$ and $D_{1}=C_{1}$ we obtain the required property. ∎
Lemma 11.
Suppose $\sigma$ is a linear congruence on $\mathbf{A}\in\mathcal{V}_{n}$ such that $\sigma^{*}=A^{2}$. Then $\mathbf{A}/\sigma\cong\mathbf{Z}_{p}$ for some prime $p$.
Proof.
Since $\sigma$ is linear, applying Lemma 27 a nontrivial bridge from $\sigma$ to $\sigma$ we derive that $\sigma$ is a perfect linear congruence. Hence we have $\zeta\leqslant\mathbf{A}\times\mathbf{A}\times\mathbf{Z}_{p}$ with $\operatorname{pr}_{1,2}(\zeta)=A^{2}$. Choose some element $a\in A$ and put $\xi(x,z)=\zeta(x,a,z)$. Then $\xi$ is a bijective relation giving an isomorphism $\mathbf{A}/\sigma\cong\mathbf{Z}_{p}$. ∎
Lemma 12.
Suppose $\sigma$ is a PC congruence on $\mathbf{A}$ and $\sigma^{*}=A^{2}$. Then $\mathbf{A}/\sigma$ is a PC algebra.
Proof.
To show that $\mathbf{A}/\sigma$ is a PC algebra it is sufficient to show that any reflexive $R\leqslant(\mathbf{A}/\sigma)^{m}$ can be represented as a conjunction of equality relations. Assume the converse and consider a relation $R\leqslant(\mathbf{A}/\sigma)^{m}$ of the minimal arity that is not like this. Then projection of $R$ onto any subset of coordinates gives a full relation. Choose some tuple $(a_{1},\dots,a_{m})\notin R$. Notice that $\{a_{i}\}<_{\mathcal{PC}}^{A/\sigma}A/\sigma$ for every $i$. Applying Corollary 22 we derive that $m$ must be equal to 2 and using the fact that $R$ is reflexive, we derive that $R$ is the equality relation, which contradicts our assumptions. ∎
Lemma 95.
Suppose $R\leqslant_{sd}\mathbf{A}_{1}\times\mathbf{A}_{2}$, $C_{i}\leqslant_{\mathcal{D}(\sigma_{i})}^{A_{i}}B_{i}\lll A_{1}$ for $i=1,2$, $S$ is the rectangular closure of $R$, $R\cap(B_{1}\times C_{2})\neq\varnothing$, $R\cap(C_{1}\times B_{2})\neq\varnothing$, and $S\cap(C_{1}\times C_{2})\neq\varnothing$. Then $R\cap(C_{1}\times C_{2})\neq\varnothing$.
Proof.
Let $\delta_{i}$ be the intersection of all irreducible congruences coming from $C_{i}\lll A_{i}$ for $i=1,2$. Let $C_{i}^{\prime}=C_{i}\circ\delta_{i}$, $B_{i}^{\prime}=B_{i}\circ\delta_{i}$, By Corollary 17(e) $C_{i}^{\prime}<_{\mathcal{D}(\sigma_{i})}^{A_{i}}B_{i}^{\prime}\lll A_{i}$ for $i=1,2$.
Assume that $R\cap(C_{1}^{\prime}\times C_{2}^{\prime})\neq\varnothing$. Notice that $C_{i}=C_{i}^{\prime}\cap B_{i}$. If $R\cap(C_{1}\times C_{2})\neq\varnothing$, then we are done. Otherwise, without loss of generality (we can switch 1 and 2 if it is not true) there are $B_{1}\lll^{A_{1}}F_{1}<_{T(\xi)}^{A_{1}}E_{1}\lll A_{1}$ and $B_{2}\lll^{A_{2}}E_{2}\lll A_{2}$ such that $R\cap((C_{1}^{\prime}\cap F_{1})\times(C_{2}^{\prime}\cap E_{2}))=\varnothing$ and $R\cap((C_{1}^{\prime}\cap E_{1})\times(C_{2}^{\prime}\cap E_{2}))\neq\varnothing$. Since $B_{1}^{\prime}\cap F_{1}\supseteq B_{1}$, we have $R\cap((B_{1}^{\prime}\cap F_{1})\times(C_{2}^{\prime}\cap E_{2}))\neq\varnothing$. and by Theorem 21 $T=\mathcal{D}$. Since $\xi\supseteq\delta_{1}$, $C_{1}^{\prime}\cap F_{1}=C_{1}^{\prime}\cap E_{1}$, which contradicts our assumptions.
Thus, we assume that $R\cap(C_{1}^{\prime}\times C_{2}^{\prime})=\varnothing$. Let $R^{\prime}=\delta_{1}\circ R\circ\delta_{2}$. Notice that $R^{\prime}\cap(C_{1}^{\prime}\times C_{2}^{\prime})=\varnothing$.
Consider two cases:
Case 1. $(\operatorname{LeftLinked}(R^{\prime})\cap(B_{1}^{\prime})^{2})/\delta_{1}=(B_{1}^{\prime})^{2}/\delta_{1}$. By Lemma 91 $\operatorname{LeftLinked}(R^{\prime}\cap(B_{1}^{\prime}\times C_{2}^{\prime}))/\delta_{1}=(B_{1}^{\prime})^{2}/\delta_{1}$, which implies that $R^{\prime}\cap(C_{1}^{\prime}\times C_{2}^{\prime})\neq\varnothing$.
Case 2. $(\operatorname{LeftLinked}(R^{\prime})\cap(B_{1}^{\prime})^{2})/\delta_{1}\neq(B_{1}^{\prime})^{2}/\delta_{1}$. As $R^{\prime}\cap(B_{1}^{\prime}\times C_{2}^{\prime})\neq\varnothing$, there should be $C_{1}^{\prime\prime}<_{\mathcal{D}(\sigma_{1})}^{A_{1}}B_{1}^{\prime}$ such that $R^{\prime}\cap(C_{1}^{\prime\prime}\times C_{2}^{\prime})\neq\varnothing$. Notice that $C_{1}^{\prime\prime}\neq C_{1}^{\prime}$. Since $S\cap(C_{1}\times C_{2})\neq\varnothing$, $C_{1}^{\prime}\times C_{1}^{\prime\prime}\subseteq\operatorname{LeftLinked}(R^{\prime})$, which contradicts condition 1 of Lemma 92. ∎
Lemma 96.
Suppose $R\leqslant_{sd}\mathbf{A}_{1}\times\mathbf{A}_{2}$, $C_{1}\leqslant_{\mathcal{\mathcal{D}(\sigma)}}^{A_{1}}B_{1}\lll A_{1}$, $B_{2}\lll A_{2}$, $S$ is a rectangular closure of $R$, $R\cap(B_{1}\times B_{2})\neq\varnothing$, $S\cap(C_{1}\times B_{2})\neq\varnothing$. Then $R\cap(C_{1}\times B_{2})\neq\varnothing$.
Proof.
Assume the converse. Consider $C_{2}^{\prime}$ and $B_{2}^{\prime}$ such that $B_{2}\lll^{A_{2}}C_{2}^{\prime}<_{T}^{A_{2}}B_{2}^{\prime}\lll A_{2}$, $R\cap(C_{1}\times C_{2}^{\prime})=\varnothing$, and $R\cap(C_{1}\times B_{2}^{\prime})\neq\varnothing$. By Theorem 21 $T=\mathcal{D}$. Then by Lemma 95 $R\cap(C_{1}\times C_{2}^{\prime})\neq\varnothing$, which contradicts our assumptions. ∎
Lemma 24.
Suppose $R\leqslant_{sd}\mathbf{A}_{1}\times\mathbf{A}_{2}$, $C_{i}\leqslant_{\mathcal{\mathcal{MD}}}^{A_{i}}B_{i}\lll A_{i}$ for $i\in\{1,2\}$, $S$ is a rectangular closure of $R$, $R\cap(B_{1}\times B_{2})\neq\varnothing$, $S\cap(C_{1}\times C_{2})\neq\varnothing$. Then $R\cap(C_{1}\times C_{2})\neq\varnothing$.
Proof.
To prove the lemma it is sufficient to combine Lemma 23 and Lemma 96. ∎
Lemma 97.
$C_{1}<_{\mathcal{M}T}^{A}B_{1}\lll A$, $T\in\{\mathcal{PC},\mathcal{L},\mathcal{D}\}$, $B_{2}\lll A$, $C_{1}\cap B_{2}=\varnothing$, $B_{1}\cap B_{2}\neq\varnothing$. Then $(C_{1}\circ(\omega_{1}\cap\dots\cap\omega_{s}))\cap B_{2}=\varnothing$, where $\omega_{1},\dots,\omega_{s}$ are all congruences of type $T$ on $\mathbf{A}$ such that $\omega_{i}^{*}\supseteq B_{1}^{2}$.
Proof.
We prove the claim by induction on the size of $B_{1}$ starting with $B_{1}=A$. Thus, the inductive assumption is that the lemma holds for all greater $B_{1}$.
Let $C_{1}=C_{1}^{1}\cap\dots\cap C_{1}^{t}$, where $C_{1}^{i}<_{\mathcal{D}(\sigma_{i})}^{A}B_{1}$ for every $i\in[t]$. Notice that $\sigma_{i}\in\{\omega_{1},\dots,\omega_{s}\}$. By the definition of $<_{\mathcal{D}}^{A}$ we have $C_{1}^{i}=(C_{1}^{i}\circ\sigma_{i})\cap B_{1}$. By Corollary 17(e) and (f) we have $C_{1}^{i}\circ\sigma_{i}<_{\mathcal{D}(\sigma_{i})}^{A}B_{1}\circ\sigma_{i}\lll A$. Applying Theorem 21 to $(C_{1}^{1}\circ\sigma_{1})\cap\dots\cap(C_{1}^{t}\circ\sigma_{t})\cap B_{1}\cap B_{2}=\varnothing$ we obtain one of the two cases:
Case 1. $(C_{1}^{1}\circ\sigma_{1})\cap\dots\cap(C_{1}^{t}\circ\sigma_{t})\cap B_{2}=\varnothing$. Since
$(C_{1}^{1}\circ\sigma_{1})\cap\dots\cap(C_{1}^{t}\circ\sigma_{t})\supseteq(C_{1}^{1}\cap\dots\cap C_{1}^{t})\circ(\sigma_{1}\cap\dots\cap\sigma_{t})\supseteq C_{1}\circ(\omega_{1}\cap\dots\cap\omega_{s}),$
we derive that the required intersection is empty and complete this case.
Case 2. There exist $B_{1}\lll^{A}B_{1}^{\prime}<_{\mathcal{D}}^{A}B_{1}^{\prime\prime}\lll^{A}A$ such that $\bigcap_{i\in[t]}(C_{1}^{i}\circ\sigma_{i})\cap B_{1}^{\prime}\cap B_{2}=\varnothing$ and $\bigcap_{i\in[t]}(C_{1}^{i}\circ\sigma_{i})\cap B_{1}^{\prime\prime}\cap B_{2}\neq\varnothing$. By Lemma 20(i) $\bigcap_{i\in[t]}(C_{1}^{i}\circ\sigma_{i})\cap B_{2}\lll A$. Applying the inductive assumption to $B_{1}^{\prime\prime}$ we derive that $(B_{1}^{\prime}\circ\bigcap_{i\in[s]}\omega_{s})\cap\bigcap_{i\in[t]}(C_{1}^{i}\circ\sigma_{i})\cap B_{2}=\varnothing$. Since
$(B_{1}^{\prime}\circ\bigcap_{i\in[s]}\omega_{s})\cap\bigcap_{i\in[t]}(C_{1}^{i}\circ\sigma_{i})\supseteq(C_{1}\circ\bigcap_{i\in[s]}\omega_{s})\cap\bigcap_{i\in[t]}(C_{1}^{i}\circ\sigma_{i})\supseteq C_{1}\circ\bigcap_{i\in[s]}\omega_{s}$
we obtain the required condition. ∎
Lemma 25.
Suppose $C_{1}<_{\mathcal{M}T}^{A}B_{1}\lll A$, $B_{2}\lll A$, $C_{1}\cap B_{2}=\varnothing$, $B_{1}\cap B_{2}\neq\varnothing$, $\sigma$ is a maximal congruence on $\mathbf{A}$ such that $(C_{1}\circ\sigma)\cap B_{2}=\varnothing$. Then $\sigma=\omega_{1}\cap\dots\cap\omega_{s}$, where $\omega_{1},\dots,\omega_{s}$ are congruences of type $T$ on $\mathbf{A}$ such that $\omega_{i}^{*}\supseteq B_{1}^{2}$.
Proof.
By Corollary 15(m) consider two cases:
Case 1. There exists $E<_{S}B_{1}/\sigma$. By Theorem 21 $E\cap B_{2}/\sigma\neq\varnothing$ and $E\cap C_{1}/\sigma\neq\varnothing$, hence $C_{1}/\sigma\cap B_{2}/\sigma\neq\varnothing$, which contradicts our assumptions.
Case 2. $C_{1}/\sigma\leqslant_{\mathcal{M}T}B_{1}/\sigma$. Since $C_{1}/\sigma\cap B_{2}/\sigma=\varnothing$ and $B_{1}/\sigma\cap B_{2}/\sigma\neq\varnothing$, we have $C_{1}/\sigma<_{\mathcal{M}T}B_{1}/\sigma$. By Lemma 97 $(C_{1}/\sigma\circ(\delta_{1}\cap\dots\cap\delta_{r}))\cap B_{2}/\sigma=\varnothing$, where $\delta_{1},\dots,\delta_{r}$ are all the congruences of type $T$ such that $\delta_{i}^{*}\supseteq B_{1}^{2}$. Extend each congruence $\delta_{i}$ to $\mathbf{A}$ so that $\mathbf{A}/\omega_{i}\cong(\mathbf{A}/\sigma)/\delta_{i}$. Then $\omega_{i}^{*}\supseteq B_{1}^{2}$ and $(C_{1}\circ(\omega_{1}\cap\dots\cap\omega_{r}))\cap B_{2}=\varnothing$. Since $\omega_{1}\cap\dots\cap\omega_{r}\supseteq\sigma$ and $\sigma$ is a maximal congruence satisfying this condition, we obtain $\sigma=\omega_{1}\cap\dots\cap\omega_{r}$, which completes the proof. ∎